### v1

# Revisiting Semi-Supervised Learning in the Era of Foundation Models

Abstract
Semi-supervised learning (SSL) leverages abundant unlabeled data alongside limited labeled data to enhance learning. As vision foundation models (VFMs) increasingly serve as the backbone of vision applications, it remains unclear how SSL interacts with these pre-trained models. To address this gap, we develop new SSL benchmark datasets where frozen VFMs underperform and systematically evaluate representative SSL methods. We make a surprising observation: parameter-efficient fine-tuning (PEFT) using only labeled data often matches SSL performance, even without leveraging unlabeled data. This motivates us to revisit self-training, a conceptually simple SSL baseline, where we use the supervised PEFT model to pseudo-label unlabeled data for further training. To overcome the notorious issue of noisy pseudo-labels, we propose ensembling multiple PEFT approaches and VFM backbones to produce more robust pseudo-labels. Empirical results validate the effectiveness of this simple yet powerful approach, providing actionable insights into SSL with VFMs and paving the way for more scalable and practical semi-supervised learning in the era of foundation models. [^1] Our code is available at https://github.com/OSU-MLB/SSL-Foundation-Models.
semi-supervised learning, self-training, vision foundation models, foundation models, benchmark, parameter-efficient fine-tuning
1 Introduction
The quality of machine learning models is often closely tied to the amount of labeled data available, but annotation can be costly or labor-intensive. As a result, semi-supervised learning (SSL), which leverages abundant unlabeled data alongside limited labeled data, has emerged as a promising paradigm for developing ML models without the need for extensive labeling (Zhu, 2008; Chapelle et al., 2006; Van Engelen & Hoos, 2020; Yang et al., 2022). Over the past few decades, numerous SSL algorithms have been developed to advance this field. Exemplar methods from the deep learning era include Mean Teacher (Tarvainen & Valpola, 2018), MixMatch (Berthelot et al., 2019), and FixMatch (Sohn et al., 2020), which dynamically impose objective functions on unlabeled data based on distillation and consistency, thereby enhancing learning performance. It is worth noting that many of these methods were originally designed to train neural networks from scratch.
Meanwhile, pre-training on external labeled or unlabeled data has become the de facto standard across many machine learning application domains (Gui et al., 2024). For example, in computer vision, many recent algorithms are built upon vision foundation models (VFMs) such as CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2024). These models were pre-trained on massive datasets—–millions, if not billions, of data points. As a result, they have demonstrated remarkable generalizability across a wide range of tasks, often requiring only slight fine-tuning or, in some cases, functioning effectively as a frozen backbone (Zhang et al., 2024).
Figure 1: Illustration of our proposed SSL pipeline. To effectively leverage abundant unlabeled data alongside scarce labeled data in the era of VFMs, our approach follows four phases: (1) Supervised Parameter-Efficient Fine-Tuning, where we harness labeled data by fine-tuning pre-trained VFMs using various PEFT algorithms; (2) Pseudo-Label Generation, where we exploit fine-tuned VFMs’ generalization ability to generate pseudo-labels for unlabeled data; (3) Pseudo-Label Ensemble, where we enhance robustness by aggregating pseudo-labels from multiple fine-tuned VFMs; and (4) Self-Training, where we consolidate all knowledge into one model.
Given the promising advancements in both fields, we explore their interaction in this paper. Specifically, we seek to address the following questions: Are existing SSL algorithms still effective when using VFMs as the backbone? What adjustments, if any, are needed to improve their performance? Finally, can we leverage the power of VFMs to develop more effective, yet simpler, SSL algorithms?
To this end, we introduce new SSL benchmark datasets based on the Visual Task Adaptation Benchmark (VTAB) (Zhai et al., 2020), a diverse suite of classification tasks designed to evaluate visual representations. Our focus is on tasks where frozen VFMs underperform, indicating the need for further fine-tuning, and where SSL could offer a beneficial solution. We then systematically evaluate three recent representative SSL methods—FixMatch (Sohn et al., 2020), FlexMatch (Zhang et al., 2022), SoftMatch (Chen et al., 2023) and FineSSL (Gan & Wei, 2024). Hyperparameters are carefully selected using techniques proposed for unsupervised domain adaptation (Ericsson et al., 2023), ensuring that the data leakage issue discussed in (Oliver et al., 2018) is avoided.
Our empirical results show that SSL does not offer a significant advantage over fully fine-tuning VFMs with limited labeled data. In some cases, accuracy even decreases. We hypothesize that allowing SSL to update all parameters in VFMs may inadvertently reduce their built-in generalizability. Indeed, in practical applications of VFMs, parameter-efficient fine-tuning (PEFT)—which updates only a small fraction of parameters or introduces a lightweight learnable module while keeping the VFM largely frozen—often outperforms full fine-tuning, particularly when downstream labeled data is limited (Mai et al., 2024; Xin et al., 2024). When we integrate PEFT into SSL, performance improves significantly. However, even when PEFT alone is used to fine-tune VFMs, the performance gains often match or exceed those by SSL, all without requiring unlabeled data.
These findings convey two key messages. First, new SSL algorithms specifically designed for VFMs are needed. Second, the fact that supervised PEFT matches SSL accuracy suggests it can already generate high-quality pseudo-labels for unlabeled data. Appropriately leveraging these pseudo-labels could further improve model performance.
Building on these insights, we revisit self-training, a conceptually simple SSL baseline that uses the current model to pseudo-label unlabeled data for further training (Lee et al., 2013; Xie et al., 2020; Chen et al., 2020; Yalniz et al., 2019). A key challenge in self-training is ensuring the reliability of pseudo-labels—–incorrect labels can reinforce model biases and degrade performance. Traditional approaches rely on confidence thresholds or consistency constraints to filter pseudo-labels (Sohn et al., 2020; Nguyen & Yang, 2023), but these methods are often heuristic and suboptimal, limiting self-training performance compared to more advanced SSL methods.
To address this, we propose an alternative strategy that leverages the complementary behaviors of VFMs and PEFT. Different VFMs and PEFT algorithms often produce diverse predictions, even when their overall accuracies are  similar (Mai et al., 2024; Tong et al., 2024b).
This inherent diversity enables us to harness ensembling techniques (Zhou, 2012). By aggregating predictions from multiple widely used VFM backbones fine-tuned with different PEFT methods—and explicitly compensating for their varied confidence distributions—we obtain significantly more robust pseudo-labels, reducing the need for explicit pseudo-label selection.
As a result, our ensemble-based approach yields a more reliable yet simpler self-training pipeline (Figure 1) that effectively leverages labeled and unlabeled data as well as the strengths of VFMs and PEFT, leading to substantial improvements over existing SSL methods.
We extensively validate our approach on new benchmark datasets, where it outperforms existing SSL methods across most tasks, establishing it as a simple, effective, and competitive SSL baseline in the era of foundation models. Further analyses provide key insights for future improvements.
Table 1: Comparison of supervised linear probing performance between popular SSL datasets (CIFAR, Food101) and our benchmark, where N denotes the number of labeled samples per class. While frozen VFMs already excel in standard SSL datasets, they struggle in our benchmark—comprising diverse tasks, domains, and sizes—underscoring the potential of SSL for unleashing the full potential of VFMs.
CIFAR-10 CIFAR-100 FOOD-101 DTD SUN397 Resisc45 Retinopathy Clevr-C KITTI
N $4$ $100$ $10$ $6$ $6$ $2$ $80$ $20$ $10$
CLIP 85.0% 78.3% 80.2% 61.8% 63.7% 69.3% 35.9% 33.1% 51.1%
DINOv2 91.7% 88.1% 83.1% 66.7% 65.3% 52.4% 41.2% 30.5% 51.3%
Our key contributions are as follows:
  • We introduce new SSL benchmark datasets tailored for studying SSL in the context of foundation models.
  • We conduct a systematic evaluation of representative SSL methods, offering valuable insights for future research.
  • We propose a simple yet highly effective self-training-based SSL method leveraging the key properties of VFMs and PEFT, establishing a strong baseline for future work.
Remark. Self-training (Amini et al., 2025), pseudo-labeling (Lee et al., 2013), and ensembling (Dong et al., 2020) have been extensively studied in the ML literature, spanning both the pre-deep learning and deep learning eras. Our goal is not to compete with existing methods but to establish a simple yet effective semi-supervised learning baseline that harnesses their strengths while incorporating the unique properties of foundation models. By maintaining a clean and accessible approach, we aim to facilitate future research. While we acknowledge that integrating more sophisticated techniques could further enhance performance, we deliberately focus on simplicity.
A key step in ensembling is obtaining multiple base learners that are both diverse and equally performant. Our contribution lies in leveraging the complementary behaviors of multiple foundation models and PEFT methods—an approach tailored specifically for the foundation model era.
2 Related Works
Semi-Supervised Learning. In the past few years, numerous SSL algorithms have been developed, many of which focus on generating and selecting reliable pseudo-labels during the training process (Sohn et al., 2020; Berthelot et al., 2019, 2020). For example, FixMatch (Sohn et al., 2020) relies on a fixed threshold to select pseudo-labels, while more recent methods dynamically adjust this threshold as training proceeds. FlexMatch (Zhang et al., 2022) introduces class-specific thresholds to accommodate varying degrees of difficulty, SoftMatch (Chen et al., 2023) applies a soft confidence threshold to balance label quality and quantity. However, most existing approaches were originally designed to train models from scratch, leaving open questions about their compatibility with VFMs and the adjustment required for optimal performance. Another prominent direction in SSL is self-training (Xie et al., 2020; Lee et al., 2013), typically comprising three steps: (1) train a teacher model on labeled data, (2) generate pseudo-labels on unlabeled data using the teacher, and (3) train a student model on both labeled and pseudo data. A key challenge lies in ensuring the reliability of pseudo-labels. Conventional solutions employ unsupervised pre-training (Chen et al., 2020), confidence thresholds or consistency constraints (Sohn et al., 2020; Nguyen & Yang, 2023) to improve the pseudo-label quality. In contrast, we aim to leverage the power of VFMs to produce more robust pseudo-labels and turn self-training into a simple yet powerful baseline in the VFM era.
Vision Foundation Models. Vision Transformers (ViT) (Dosovitskiy et al., 2021) pre-trained on massive amounts of data, has become indispensable to modern AI development (Awais et al., 2025). These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks. For example, CLIP-ViT (Radford et al., 2021) with millions of image-text pairs show an unprecedented zero-shot capability, robustness to distribution shifts and serve as the encoders for various power generative models (Rombach et al., 2022; Liu et al., 2024)). Meanwhile, DINOv2 (Oquab et al., 2024) pre-trained with self-supervised objectives on extensive sets of well-curated images effectively captures fine-grained localization features. Despite this growing prominence of VFMs, limited work has explored how to integrate them with SSL. Although a recent study marks an initial attempt (Gan & Wei, 2024), a comprehensive study of how to effectively leverage VMFs in SSL is still lacking.
3 On Evaluation of SSL in the Era of VFMs
3.1 Problem Definition
We consider a $C$-class classification problem in SSL with a large unlabeled dataset $\mathcal{D}^{u}\coloneqq\{{\bm{x}}^{u}_{i}\}_{i=1}^{N^{u}}$ and a much smaller labeled dataset $\mathcal{D}^{l}\coloneqq\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{N^{l}}$, where ${\bm{x}}^{l}_{i}$ and ${\bm{x}}^{u}_{i}$ are training samples and $y^{l}_{i}$ is the ground-truth class label. $N^{u}$ and $N^{l}$ denote the unlabeled and labeled dataset size with $N^{u}\gg N^{l}$. SSL aims to learn a model $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ using $\mathcal{D}^{u}$ and $\mathcal{D}^{l}$. Unlike conventional SSL that initializes $\bm{\theta}$ randomly, our framework starts with a VFM (e.g., CLIP or DINOv2) and fine-tunes it for downstream tasks.
3.2 A Diverse and Challenging SSL Benchmark
Despite recent advancements in SSL, most studies (Gan & Wei, 2024; Chen et al., 2023; Zhang et al., 2022) still evaluate performance on classic datasets such as CIFAR-10/100 (Krizhevsky & Hinton, 2009), STL-10 (Coates et al., 2011), and Food101 (Bossard et al., 2014). These benchmarks were suitable when models were randomly initialized, but with the advent of VFMs, they have become far less challenging. As shown in Table 1, linear probing on top of frozen VFM backbones already achieves remarkable accuracy on these benchmarks, even with fairly limited labeled data. Moreover, these datasets primarily focus on natural images, providing a limited perspective on real-world scenarios where VFMs are deployed.
Table 2: Summary of our benchmark. Rem. Sen.: Remote sensing; Recog.: Recognition; $N^{l}/N^{u}$: # labeled/unlabeled training data.
Dataset Task Domain $N^{u}+N^{l}$ Classes $N^{l}/$Class
DTD Recog. Textual $3,008$ $47$ $3,6$
SUN397 Recog. Natural $49,601$ $973$,6363,63 , 6
Resisc45 Recog. Rem. Sen. $20,160$ $45$ $1,2$
Retinopathy Recog. Medical $36,825$ $5$ $40,80$
Clevr-C Count Synthetic $56,000$ $8$ $10,20$
KITTI Depth Auto Drive $5,416$ $4$ $5,10$
To effectively evaluate SSL methods in the VFM era, we propose a new benchmark designed to capture the complexities of real-world applications across different tasks, domains, and dataset sizes. Due to these intricacies, frozen VFMs face much greater challenges in our benchmark, as shown in Table 1, reinforcing the necessity of SSL to unleash the potential of VFMs. Specifically, we select six datasets from VTAB (Zhai et al., 2020) covering recognition, depth estimation, and counting tasks, drawn from textual, natural, remote sensing, medical, synthetic, and autonomous driving domains. We also vary label-per-class settings to assess the robustness of SSL methods. We summarize the datasets in Table 2 and provide detailed information inAppendix A. We believe our benchmark provides a more diverse and comprehensive evaluation of SSL methods in the VFM era.
3.3 A Fair Hyperparameter Tuning for SSL
Hyperparameter tuning has long been a persistent challenge in SSL (Oliver et al., 2018), remaining ambiguous and unstandardized throughout much of the existing literature. Due to the labeled data scarcity, standard train-validation split is often infeasible. Tuning on a held-out labeled validation or even test set can lead to significant data leakage, resulting in over-optimistic results and unfair comparisons. To tackle this issue and complement our benchmark, we establish a more rigorous protocol for hyperparameter tuning in SSL.
Our core insight is to harness the defining characteristic of SSL—an abundance of unlabeled training data—to tune parameters in an unsupervised manner and avoid the pitfalls of data leakage. One recent study has explored such an idea for domain adaptation, using unsupervised criteria such as RankMe (Garrido et al., 2023) and AMI (Vinh et al., 2009) to estimate the effectiveness of each hyperparameter configuration (Ericsson et al., 2023).
However, Ericsson et al. (2023) also revealed that no single criterion could reliably select suitable hyperparameters across all scenarios. Motivated by this finding, we propose integrating seven unsupervised criteria—five derived from features, AMI (Vinh et al., 2009), ARI (Halkidi et al., 2002), V-Measure (Rosenberg & Hirschberg, 2007), FMI (Fowlkes & Mallows, 1983) and BNM (Cui et al., 2020), and two from logits, RankMe (Garrido et al., 2023) and CHI (Caliński & Harabasz, 1974)—for more robust tuning. Concretely, for each hyperparameter configuration and its corresponding model, we compute all seven criteria using a held-out unlabeled validation set $\mathcal{D}^{v}\coloneqq\{{\bm{x}}^{v}_{i}\}_{i=1}^{N^{v}}$. Next, we rank every hyperparameter configuration based on each criterion and choose the one achieving the lowest average rank across all. We provide formal definitions and detailed procedures in Section C.1. By eliminating reliance on held-out labeled validation sets, our method mitigates data leakage and promotes a more practical and fair tuning protocol for SSL.
4 Systematic Evaluation of SSL with VFMs
Given the struggling performance of frozen VFMs in our benchmark (Table 1), we consider two straightforward strategies to enhance their effectiveness: (1) exploit the inherent generalizability of VFMs by fine-tuning only on labeled data; (2) employ SSL to leverage both labeled and unlabeled data. With the diverse benchmark and hyperparameter tuning protocol introduced in Section 3, we aim to investigate whether existing SSL algorithms remain effective when adopting VFMs as their backbone.
Evaluation Setup. We focus on two representative VFMs—ViT-B/16 CLIP (Radford et al., 2021) and ViT-B/14 DINOv2 (Oquab et al., 2024)—covering language-image contrastive pre-training and self-supervised pre-training strategies, respectively. We examine four representative SSL methods, including FixMatch (Sohn et al., 2020), FlexMatch (Zhang et al., 2022), SoftMatch (Chen et al., 2023), and FineSSL (Gan & Wei, 2024) and use labeled-only fine-tuning as the baseline. We employ the AdamW optimizer (Loshchilov & Hutter, 2019) with a batch size of 32 and weight decay $5\times 10^{-4}$ to fine-tune the model for 35 epochs. Learning rates and other hyperparameters are tuned with our proposed tuning protocol. The classification performance is evaluated using the Top-1 accuracy on the test dataset.
Figure 2: Average accuracy of SSL methods with full fine-tuning or PEFT (LoRA, AdaptFormer) across 12 settings in our benchmark. Under fair hyperparameter tuning, full fine-tuning with limited labeled data alone can surpass SSL methods. Although PEFT enhances SSL, PEFT with labeled data alone remains highly competitive, indicating limited benefit from unlabeled data in current SSL approaches when paired with VFMs. We present detailed results in Section B.3.
Surprising Effectiveness of Labeled-only Fine-tuning. Surprisingly, under fair comparison, full fine-tuning with even a few labeled images per class can match or surpass SSL methods, as illustrated in Figure 2. In other words, even with a large amount of additional unlabeled data, SSL provides little advantage over fine-tuning VFMs with limited labeled data. This finding is reminiscent of observations made in a comprehensive investigation conducted seven years ago (Oliver et al., 2018).
Due to the inherently noisy supervised signals associated with unlabeled data in SSL methods, we hypothesize that allowing SSL to update all parameters in VFMs may inadvertently reduce their built-in generalizability.
Can PEFT Come to the Rescue? When labeled downstream data are scarce, recent research has shown that parameter-efficient fine-tuning (PEFT)—which updates only a small subset of parameters or introduces a lightweight learnable module to frozen VFMs—often outperforms full fine-tuning on VFMs (Mai et al., 2024; Xin et al., 2024; Hu et al., 2021; Chen et al., 2022; Jia et al., 2022a; Tu et al., 2023). When considering SSL scenarios with limited labeled data, a natural question arises: is PEFT compatible with SSL? Although a recent study has made an initial attempt to apply PEFT in SSL (Gan & Wei, 2024), a comprehensive analysis of its compatibility with different VFM backbones, PEFT methods, and SSL approaches is still lacking. We, therefore, assess two commonly used PEFT strategies: LoRA (Hu et al., 2021), which trains additive low-rank matrices for Transformer layers to approximate weight updates, and AdaptFormer (Chen et al., 2022), an adapter-based approach proven effective in computer vision. We follow the same protocol outlined in Section 3 to tune the hyperparameters of these PEFT methods.
Our results in Figure 2 show that PEFT indeed improves SSL on VFMs. However, it also enhances labeled-only fine-tuning, with results remaining comparable to SSL. This highlights the limited effectiveness of using unlabeled data in existing SSL methods when paired with VFMs.
Discussion. Three key takeaways emerge from these findings. First, a well-tuned labeled-only PEFT serves as a competitive baseline for SSL with VFMs. Second, the ineffectual use of unlabeled data in existing SSL methods underscores the need for SSL approaches specifically designed for VFMs. Finally, given that labeled-only PEFT can achieve accuracy comparable to SSL, its predictions on unlabeled data generate pseudo-labels of sufficient quality. A promising direction is to leverage these pseudo-labels to further enhance model performance.
5 A Simple, Effective SSL Baseline for VFMs
To wrap up the discussion in Section 4, we propose a new semi-supervised learning method tailored to vision foundation models. Specifically, we build on self-training (Triguero et al., 2015; Lee et al., 2013; Blum & Mitchell, 1998; Balcan et al., 2005; McClosky et al., 2006b, a), a conceptually simple SSL baseline that uses pseudo-labels from unlabeled data as additional supervision to improve the model. In the following section, we first briefly review self-training.
5.1 Self-training
Let us denote the predicted posterior probability of class $c$ given ${\bm{x}}$ as $p(c|{\bm{x}};f_{\bm{\theta}})$. The core idea of self-training is to gradually assign pseudo-labels to unlabeled data $\mathcal{D}^{u}=\{{\bm{x}}^{u}_{i}\}_{i=1}^{N^{u}}$ based on high-confidence predictions
$\hat{y}^{u}_{i}\coloneqq\operatorname{arg\,max}_{c}p(c|{\bm{x}}^{u}_{i};f_{\bm{\theta}})\hskip 20.0pt\text{ if }\max_{c}p(c|{\bm{x}}^{u}_{i};f_{\bm{\theta}})\geq\tau,$
and add these pseudo-labeled data to the labeled set $\mathcal{D}^{l}=\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{N^{l}}$ for supervised learning. Here, $\bm{\theta}$ represents the weights of the current classifier, and $\tau$ is the confidence threshold. If (a) high-confidence predictions are correct and (b) the updated model gradually increases its confidence in the remaining data, self-training can be as effective as fully supervised learning on the entire dataset with true labels.
In our context, a VFM fine-tuned with PEFT using labeled $\mathcal{D}^{l}$ serves as the current model to generate pseudo-labels for $\mathcal{D}^{u}$, which are then used to further fine-tune the VFM.
Challenges. While self-training is conceptually simple, its practical implementation remains challenging, primarily due to the selection of $\tau$. Ideally, $\tau$ should exclude samples with incorrect pseudo-labels. However, if set too high, it limits the number of pseudo-labeled samples available for supervised learning, reducing effectiveness. Conversely, if $\tau$ is too low, erroneous pseudo-labels may be included, reinforcing model biases. Further complicating the issue, the optimal $\tau$ can vary across self-training iterations. While extensive research has been conducted on pseudo-label selection (Amini et al., 2025; Yang et al., 2022), existing methods are often heuristic or overly complex, inadvertently making self-training impractical or unreliable.
5.2 Our objective
To leverage the high-quality pseudo-labels from label-only PEFT while circumventing the aforementioned challenges, we aim to develop a self-training-based SSL algorithm that is both simple and effective in practice. Specifically, we seek to eliminate the need for complex pseudo-label selection by utilizing all available pseudo-labels (i.e., setting $\tau=0$). This also allows self-training to be completed in a single round, thereby removing an additional hyperparameter.
5.3 Ensembling of multiple PEFT and VFMs
A necessary step toward achieving our objective is to ensure pseudo-labels of even higher quality to prevent error propagation. To this end, we employ ensembling techniques (Zhou, 2012; Dietterich, 2000), which combine the predictions from multiple models to enhance the overall quality and reliability of pseudo-labels.
What to Ensemble? Unlike conventional methods that train multiple equally performant yet diverse base learners through bootstrapping, random initialization, or cyclic learning rate schedules (Huang et al., 2017), our approach exploits the unique properties of PEFT (Mai et al., 2024) and VFMs (Tong et al., 2024b, a). Specifically, it has been observed that different PEFT methods, while achieving similar overall accuracy on downstream tasks, generate diverse predictions for individual samples. Similarly, VFMs trained on different datasets with varying objective functions exhibit diverse capabilities, with no single model demonstrating clear overall dominance.
How to Ensemble? Given $K$ models $f_{\bm{\theta}_{1}},\dots,f_{\bm{\theta}_{K}}$, several strategies can be applied to integrate their predictions for an unlabeled image ${\bm{x}}^{u}_{i}$. Two common approaches are: (a) Mean Logits, where the logits for each class are averaged across the models; and (b) Mean Probabilities, where the predicted probabilities for each class are averaged. The assumption behind these methods is that the models produce logits or probabilities at similar scales, which is typically the case when the models share similar architectures and are trained on bootstrapped data.
In our context, since different VFMs were trained in distinct ways and adjusted using various PEFT methods, we find that the assumption of similar prediction scales often does not hold, as shown in Figure 3. As a result, some models may dominate the ensembling process, reducing its effectiveness. To address this, we introduce a simple solution called Mean Labels. In this approach, we first obtain “hard” pseudo-labels from each model $\hat{y}_{i,k}^{u}=\operatorname{arg\,max}_{c}p(c|{\bm{x}}_{i}^{u};f_{\bm{\theta}_{k}})$, convert them into one-hot encoding, and average them to obtain a “soft” pseudo-label $\bar{{\bm{e}}}_{i}$. This ensures uniform prediction scales across all models (since the result is always one-hot) while preserving each model’s prediction.
Figure 3: The average entropy of predicted probability distributions from different fine-tuned VFMs. This figure illustrates the gap in entropy among the pseudo-labels generated by different models, indicating that they are not well-calibrated.
5.4 Overall Algorithm (Figure 1)
We summarize our self-training-based SSL algorithm enhanced by ensembling multiple VFMs and PEFT as follows.
Stage 1. Let $n$ denote a specific VFM and $m$ denote a specific PEFT, we fine-tune $\bm{\theta}_{n}$ on $\mathcal{D}^{l}$ to obtain $\bm{\theta}_{n,m}$.
Stage 2. For an unlabeled image ${\bm{x}}_{i}^{u}\in\mathcal{D}^{u}$, we apply $\bm{\theta}_{n,m}$ to obtain its one-hot prediction ${\bm{e}}_{\hat{y}_{i,n,m}^{u}}\in\Delta^{C-1}$, where $\hat{y}_{i,n,m}^{u}=\operatorname{arg\,max}_{c}p(c|{\bm{x}}_{i}^{u};f_{\bm{\theta}_{n,m}})$.
Stage 3. We average the one-hot encoding across all fine-tuned models to obtain the soft pseudo-label
$\bar{{\bm{e}}}_{i}\coloneqq\cfrac{1}{N\times M}\sum_{n}\sum_{m}{\bm{e}}_{\hat{y}_{i,n,m}^{u}}\in\Delta^{C-1}.$ (1)
Stage 4. Let $\mathcal{D}^{p}=\{({\bm{x}}^{u}_{i},\bar{{\bm{e}}}_{i})\}_{i=1}^{N^{u}}$ denote the pseudo-labeled dataset, we re-fine tune a targeted VFM $n\star$ by a targeted PEFT $m\star$ on $\mathcal{D}^{p}$, using the Kullback-Leibler divergence loss function. The resulting $\bm{\theta}_{n\star,m\star}$ is the final SSL model.
6 Experiment
This section aims to answer the following questions: (1) How does the proposed method perform compared with the existing SSL methods? (2) Does the proposed method scale with more PEFT methods and VFMs? (3) How does the proposed ensemble method perform? (4) How effective is our proposed hyperparameter tuning protocol? We follow the setup described in Section 3 and Section 4.
DTD Sun397 Resisc45 Retinopathy Clevr-C KITTI Average
3 6 3 6 1 2 40 80 10 20 5 10
CLIP
LoRA Labeled Only 57.9% 64.1% 60.7% 66.8% 59.5% 72.1% 35.4% 42.2% 35.2% 50.8% 60.1% 64.3% 55.7%
Fixmatch 56.8% 68.6% 62.2% 72.8% 47.6% 77.4% 32.5% 39.1% 32.2% 42.6% 59.8% 52.6% 53.7%
FlexMatch 56.5% 67.0% 66.4% 71.9% 51.2% 78.4% 32.8% 38.7% 36.9% 75.4% 44.6% 55.0% 56.2%
SoftMatch 59.8% 68.0% 66.3% 70.9% 78.7% 83.7% 42.1% 42.9% 29.8% 71.2% 42.9% 60.5% 59.7%
FineSSL 63.1% 69.1% 56.3% 61.8% 74.2% 83.5% 35.0% 30.4% 33.6% 38.1% 53.7% 48.2% 53.9%
[…]
ST (PEFT) 63.4% 68.2% 65.6% 70.8% 67.5% 78.6% 36.4% 42.4% 38.3% 54.6% 60.3% 65.7% 59.3%
ST (PEFT + VFMs) 65.6% 71.7% 67.2% 72.8% 66.6% 77.2% 38.9% 51.1% 36.7% 58.5% 58.0% 62.3% 60.5%
AdaptFormer Labeled Only 61.5% 65.0% 61.1% 68.2% 61.0% 72.8% 34.4% 39.0% 37.8% 48.6% 57.8% 59.1% 55.6%
Fixmatch 61.2% 67.3% 64.6% 71.2% 63.1% 80.5% 29.0% 42.4% 33.1% 44.8% 30.0% 57.2% 53.7%
FlexMatch 60.3% 68.8% 64.3% 70.7% 58.1% 78.1% 31.0% 35.9% 38.5% 52.2% 34.2% 54.2% 53.9%
SoftMatch 61.3% 69.3% 65.3% 70.2% 65.8% 82.7% 38.2% 42.7% 34.5% 56.1% 33.9% 55.1% 56.3%
FineSSL 61.9% 68.1% 58.4% 64.9% 73.9% 84.2% 25.3% 29.8% 26.0% 36.9% 42.2% 47.1% 51.6%
[…]
ST (PEFT) 63.8% 68.9% 66.8% 71.7% 69.5% 80.2% 37.1% 41.5% 38.4% 55.8% 60.5% 61.7% 59.7%
ST (PEFT + VFMs) 65.7% 71.8% 67.8% 73.2% 67.9% 78.4% 38.9% 51.2% 37.0% 59.0% 58.2% 63.3% 61.0%
DINOv2
LoRA Labeled Only 61.0% 68.4% 59.7% 66.8% 48.6% 63.1% 40.7% 48.0% 35.9% 61.7% 46.7% 71.7% 56.0%
Fixmatch 57.4% 68.4% 56.6% 71.4% 36.6% 66.7% 38.1% 41.2% 35.7% 41.0% 14.8% 42.9% 47.6%
FlexMatch 52.8% 65.4% 55.7% 67.7% 33.3% 72.6% 36.6% 35.8% 37.1% 66.9% 26.2% 33.6% 48.6%
SoftMatch 53.6% 64.0% 60.4% 67.3% 46.6% 79.5% 32.1% 44.5% 33.6% 65.1% 27.4% 27.9% 50.2%
FineSSL 66.9% 72.7% 54.6% 62.1% 69.8% 87.4% 31.0% 24.6% 24.8% 31.7% 31.9% 39.8% 49.8%
[…]
ST (PEFT) 66.7% 72.2% 58.6% 65.3% 55.8% 70.6% 41.0% 54.3% 33.1% 56.9% 42.6% 63.3% 56.7%
ST (PEFT + VFMs) 67.8% 74.1% 67.1% 73.2% 66.0% 77.0% 39.2% 52.5% 36.8% 59.3% 58.2% 63.7% 61.2%
AdaptFormer Labeled Only 63.0% 67.6% 58.9% 67.0% 47.0% 57.4% 34.4% 53.1% 27.3% 41.8% 52.3% 49.8% 51.6%
Fixmatch 59.5% 70.4% 57.9% 69.7% 26.4% 62.3% 38.7% 45.6% 32.7% 38.9% 19.8% 50.4% 47.7%
FlexMatch 57.1% 65.9% 56.0% 67.5% 51.3% 64.0% 33.4% 40.2% 33.4% 42.2% 34.3% 27.9% 47.8%
SoftMatch 58.7% 66.2% 60.0% 66.7% 53.1% 78.1% 33.5% 32.2% 32.9% 45.0% 24.5% 30.9% 48.5%
FineSSL 64.0% 71.2% 56.0% 63.3% 62.3% 69.7% 17.7% 24.3% 25.8% 30.0% 31.8% 45.0% 46.8%
[…]
ST (PEFT) 67.2% 73.0% 63.5% 70.9% 56.2% 72.0% 40.2% 55.5% 33.1% 58.0% 45.2% 60.9% 58.0%
ST (PEFT + VFMs) 68.2% 74.2% 67.7% 73.4% 66.3% 78.0% 39.4% 52.5% 36.8% 60.0% 59.2% 61.7% 61.5%
Table 3: Performance comparison of baselines, existing SSL approaches, and our proposed methods on six diverse datasets (12 settings). The best result within each PEFT and VFM is highlighted in bold. We report the full results in Section B.2.
Figure 4: Incorporating more diverse pseudo-labels for ensemble, ST $\to$ ST(PEFT) $\to$ ST(PEFT+VFMs), boosts performance, highlighting the benefits of pseudo-label diversity.
6.1 Performance Comparison
Following Section 5.4, we use ST(PEFT+VFMs) to refer to ensembling over multiple PEFT methods and VFMs; ST(PEFT) to refer to ensembling over multiple PEFT methods but one VFM; and ST to refer to no ensembling. In all cases, when we fine-tune VFMs using pseudo-labels, we re-initialize the model from the original VFMs.
Section 6 shows the performance comparison between ST, ST(PEFT), ST(PEFT+VFMs) and existing SSL methods across all the 12 settings in our benchmark. Overall, ST(PEFT+VFMs) achieve superior performance compared with others, as shown in the Average column. To further analyze the results, Figure 5 visualize the ranking frequency, where each matrix entry $(i,j)$ indicates how often method $i$ is ranked $j$-th across 12 settings. We then compute the mean rank (number in the bracket) of each method and sort them accordingly. Although ST(PEFT+VFMs) does not claim first place in every single setting, it secures the top rank most frequently, showcasing the effectiveness of pseudo-label ensemble and establishing it as a simple yet powerful baseline in the VFM era.
Digging deeper into the three ST-based methods, we observe a consistent performance boost as more diverse pseudo-labels are introduced for the ensemble, i.e., ST $\to$ ST(PEFT) $\to$ ST(PEFT+VFMs), as shown in Figure 4. This underscores the importance of diversity among pseudo-labels. An open question emerging from these results is whether the benefits will continue to scale as even more PEFT methods and VFMs are integrated.
Figure 5: Ranking frequency of different SSL methods in the proposed benchmark. The number in $(i,j)$ indicates the frequency of method $i$ is ranked $j$-th across 12 settings. The number in the bracket indicates the average rank.
6.2 Scalability of Ensembling
While most of our experiments explore ensembles of LoRA and AdaptFormer, we also examine how well our method scales when integrating additional pseudo-label sources. In particular, we expand our evaluation across two dimensions—VFMs and PEFT methods—to assess the broader applicability of our ensembling strategy.
VFMs Beyond ViT-B (CLIP, DINOv2), we incorporate four more VFMs, for a total of six, including four ViT-B (CLIP, DINOv2, OpenCLIP, ImangeNet21k) (Cherti et al., 2023; Ridnik et al., 2021) and two ViT-L (CLIP, DINOv2).
PEFT Following the same setup, we start with LoRA and AdaptFormer, then expand our selection with four additional PEFT: ConvPass (Jie & Deng, 2022), BitFit (Zaken et al., 2022), VPT-Deep (Jia et al., 2022b), and Fact-TT (Jie & Deng, 2023). These methods span a broad range of PEFT approaches, including selective-based, adapter-based, and prompt-based techniques.
We evaluate CLIP on the DTD N3 dataset, training with LoRA, and examine ensemble performance across different numbers of pseudo-labels, ranging from 1 to 6. To gain deeper insights into the pattern, we examine two ensemble sources: (1) PEFT: Ensembling pseudo-labels generated by models trained with different PEFT methods while using the same pretrained VFM (CLIP) and (2) VFMs: Ensembling pseudo-labels generated by different VFMs, all trained using the same PEFT method (LoRA).
For each experiment, we report the average over 5 runs, each with a different randomly sampled set of pseudo-labels. As shown in Figure 6, we summarize the results as follows.
Both PEFT and VFMs Help. Whether we ensemble the pseudo-labels generated across different PEFT or VFMs, the performance is consistently improved. This indicates that both PEFT and VFMs contribute to the diversity of pseudo-labels and help to improve the ensemble robustness.
Diminishing Return. The performance improvement diminishes as we increase the number of ensemble sources. This suggests that the performance gain from ensemble is not linearly proportional to the number of sources. In order to achieve a balance between performance and computational cost, we recommend using a moderate number of ensemble sources, around 2 to 4 in this case.
6.3 Mean Label Consistently Outperforms
Furthermore, as also shown in Figure 6, we investigate three ensemble strategies: (1) Mean Label (our proposed method); (2) Mean Logits; and (3) Mean Probabilities.
Based on our observations from the previous section, we find that the Mean Label strategy is the most effective among the three. It consistently outperforms the other two strategies across different ensemble sources and ensemble sizes. This suggests that the Mean Label approach better captures the diversity of pseudo-labels and maximizes the information utilization from the ensemble sources.
6.4 Effectiveness of Hyperparameter Tuning
Table 4: Hyperparameter tuning comparison: our method versus baselines (random or chosen by one criterion). The absolute errors from oracle test accuracy (smaller the better) are reported averaged across 12 datasets, 5 SSL baselines, and 2 VFMs trained on 2 PEFT. The reported values are in the format: mean $\pm$ standard deviation.
RankMe AMI ARI V-Measure FMI CHI BNM Random Ours
7.9%$\pm$8.62.8%$\pm$5.02.8%$\pm$4.92.8%$\pm$4.92.9%$\pm$4.93.7%$\pm$7.012.0%$\pm$12.18.9%$\pm$8.12.6%$\pm$4.3%
We begin by investigating whether the proposed hyperparameter tuning strategy accurately gauges SSL methods’ performance. To illustrate its advantages, we compare it with two baselines, both measured via the absolute difference from the oracle test performance. Specifically, the oracle test performance is defined as the highest accuracy across the entire hyperparameter search space; each method selects a hyperparameter configuration, and we record how far its resulting accuracy is from this optimum. By taking the absolute value of this difference, we quantify how closely each tuning method approximates the best possible result.
Our baselines are: (1) Random expectation, which represents the average performance obtained by randomly choosing hyperparameters from the search space, and (2) Single criterion, which selects the best hyperparameter configuration using only one metric. Comparing with (1) offers insight into how well our method leverages the search space relative to a purely random choice, while comparing with (2) illustrates the advantage of integrating multiple evaluation criteria. As shown in Table 4, we measure the average absolute error between each method’s chosen accuracy and the oracle’s best accuracy (i.e., the maximum accuracy achievable in the search space) across 12 datasets, 5 SSL baselines, 2 VFMs (CLIP and DINOv2), and 2 PEFT methods (LoRA and AdaptFormer). Our method outperforms both baselines in most cases, suggesting that combining multiple metrics produces a more reliable and robust assessment of SSL methods. We hope future research adopts this approach to strengthen SSL evaluations.
Figure 6: Scaling analysis for ensembling with more PEFT methods and VFMs. The starting point is pseudo-labels generated by the model itself fine-tuned with LoRA. The performance improves as more diverse pseudo-labels are incorporated, but the gain diminishes as the number of ensemble sources increases.
6.5 Impact of Labeled-Only Initialization on SSL
Per discussion in Section 4, a natural question is whether the performance of SSL methods can be improved by initializing from a model trained on labeled data using PEFT. To explore this, we conduct experiments on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10). We first train DINOv2 models on labeled data using LoRA and then fine-tune them with SoftMatch. These datasets were chosen because their labeled-only performance exceeds that of SoftMatch by a large margin, as shown in Section 6.
The results, summarized in Table 5, suggest that even when starting from the labeled-only trained model, SSL performance still falls short. In some cases, it performs even worse than training from scratch.
Table 5: Performance comparison of starting from a model trained on labeled data and then fine-tuning using SoftMatch. The results are reported on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10). For consistency, we fix LoRA with dimension 4 for all the training.
Dataset Labeled-Only SoftMatch Labeled-Only $\rightarrow$ SoftMatch
DTD N6 68.4% 64.0% 63.5%
Diabetic Retinopathy N80 44.9% 44.5% 43.9%
KITTI N10 54.0% 27.9% 50.1%
7 Conclusion
As the machine learning community increasingly standardizes pre-trained foundation models as backbones across various applications, it becomes crucial to assess whether machine learning algorithms developed prior to the foundation model era remain effective. Specifically, given that vision foundation models (VFM) are pre-trained on millions of images, it remains unclear whether they would still benefit from abundant unlabeled data in conjunction with limited labeled data in a semi-supervised learning (SSL) setting. To address this gap, we conduct a comprehensive and systematic empirical study. Our findings reveal that VFMs can still benefit from SSL but require a specifically tailored approach. In particular, we propose a simple yet highly effective self-training algorithm, enhanced by pseudo-labels generated through the integration of multiple VFMs and parameter-efficient fine-tuning (PEFT) techniques. Overall, our study offers actionable insights into SSL with VFMs and lays the groundwork for more scalable and practical semi-supervised learning in the era of foundation models.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here. Specifically, compared to prior work in semi-supervised learning and vision foundation models, we do not see additional negative ethical and societal impacts that would be created by our work.
Appendix A Dataset Details
We select 6 datasets with a diverse ranges of applications for our experiments, including:
  • DTD (Cimpoi et al., 2014): The Describable Texture Dataset (DTD) is designed for texture pattern recognition. It includes a diverse collection of images from 47 distinct texture categories, offering rich and varied examples to evaluate texture feature extraction and classification methods under natural conditions.
  • SUN397 (Xiao et al., 2010): SUN397 is a large-scale scene recognition benchmark that covers 397 categories of both indoor and outdoor scenes. It provides a comprehensive testbed for scene classification algorithms.
  • Resisc45 (Cheng et al., 2017): Resisc45 is a remote sensing image dataset specifically created for scene classification tasks. It comprises thousands of images across 45 categories, representing diverse landscapes such as urban, rural, and industrial areas. This dataset is instrumental in developing and benchmarking algorithms for geographic information systems and environmental monitoring.
  • Diabetic-Retinopathy (Dugas et al., 2015): This dataset focuses on high-resolution retinal images used for detecting diabetic retinopathy. It includes images that span various stages of the disease, providing a robust resource for training and evaluating deep learning models aimed at early diagnosis and automated medical analysis in ophthalmology.
  • Clevr-Count (Johnson et al., 2016): Derived from the CLEVR family of datasets, Clevr-Count is centered on visual reasoning, particularly object counting tasks. Using synthetically generated images, it challenges models to accurately count objects within complex scenes, thereby assessing their ability to understand spatial relationships and compositional structures in visual data.
  • KITTI-Dist (Geiger et al., 2012): KITTI-Dist is based on the renowned KITTI benchmark and is tailored for depth estimation in vehicular contexts. It provides high-quality stereo images along with ground-truth depth maps, supporting research in autonomous driving, 3D reconstruction, and stereoscopic vision, and serves as a key resource for evaluating depth estimation algorithms.
The detailed statistics of the datasets, specifically the validation and test set sizes, are presented in Table 6.
Dataset $N^{v}$ Test Size
DTD 752 1880
SUN397 17401 21750
Resisc45 5040 6300
Retinopathy 9207 42670
Clevr-C 14000 15000
KITTI 1354 711
Table 6: Dataset statistics.
Appendix B Experiment Details
In this section, we provide additional details on the hyperparameters and computation settings used in our experiments.
B.1 Computation Details
Our experiments were conducted on a workstation equipped with eight NVIDIA RTX 6000 Ada GPUs, two AMD EPYC 9554 64-Core Processors, and 800GB of RAM. Additionally, we utilized NVIDIA Tesla V100, NVIDIA Tesla A100, and NVIDIA RTX H100 GPUs for certain experiments. All experiments were implemented using PyTorch (Paszke et al., 2019).
B.2 SSL and Labeled-Only Baseline Details
For all the experiments, we employ AdamW (Loshchilov & Hutter, 2019) optimizer with a cosine annealing learning rate scheduler (Loshchilov & Hutter, 2017) with warm-up period with $2.5\%$ of total iterations. We use a batch size of $32+32$ for all experiments, where the first $32$ corresponds to labeled data and the second $32$ corresponds to unlabeled data. We list the hyperparameter search spaces for the SSL, Labeled-Only, and ST settings in Table 7, including the drop path rate (dpr) (Huang et al., 2016), training augmentation (train-aug), LoRA dimension (lora_dim), adapter bottleneck size (adapter_bottleneck), weight decay, momentum, learning rate (lr), and number of epochs. For other unmentioned SSL algorithm related hyperparameters, we use the default values provided in the original papers.
Hyperparameter SSL Labeled-Only ST
dpr 0 [0, 0.2] 0
train-aug - [weak, strong] weak
lora_dim 4 [4, 16] 4
adapter_bottleneck 16 [4, 16] 16
weight_decay 5e-4 5e-4 5e-4
momentum 0.9 0.9 0.9
lr
Full FT: [1e-4, 1e-5, 1e-6]
PEFT: [1e-3, 1e-4, 1e-5]
Full FT: [1e-4, 1e-5, 1e-6]
PEFT: [1e-3, 1e-4, 1e-5]
Full FT: [1e-4, 1e-5, 1e-6]
PEFT: [1e-3, 1e-4, 1e-5]
epochs
Full FT: 35
PEFT: 30
Full FT: 60
PEFT: 50
Full FT: 35
PEFT: 30
Table 7: Hyperparameter search spaces for SSL, Labeled-Only, and ST settings.
As a completion of Section 6, we present the detailed results of the SSL baseline experiment with different SSL methods, including full fine-tuning, in Section B.2.
DTD Sun397 Resisc45 Retinopathy Clevr-C KITTI AVERAGE
3 6 3 6 1 2 40 80 10 20 5 10
CLIP
Full Labeled-Only 53.46% 62.82% 52.79% 59.49% 52.98% 65.51% 31.78% 32.69% 25.03% 29.62% 37.55% 65.54% 47.44%
Fixmatch 46.97% 60.37% 45.49% 60.97% 31.76% 74.48% 24.64% 45.95% 21.70% 21.05% 45.43% 45.15% 43.66%
FlexMatch 50.69% 61.06% 44.41% 61.35% 41.62% 79.05% 28.35% 38.12% 25.31% 20.75% 44.87% 47.40% 45.25%
SoftMatch 52.29% 62.55% 50.69% 60.70% 59.86% 84.67% 32.91% 24.15% 19.53% 20.02% 31.22% 57.52% 46.34%
FineSSL 7.02% 17.18% 1.46% 4.01% 7.78% 8.37% 39.64% 16.97% 19.53% 26.08% 41.49% 44.59% 19.51%
[…]
ST (PEFT) 63.19% 68.67% 63.28% 68.83% 68.32% 79.73% 39.46% 43.35% 37.33% 49.87% 60.34% 63.85% 58.85%
ST (PEFT + VFMs) 66.33% 70.64% 66.42% 72.23% 67.49% 79.11% 39.46% 55.01% 35.61% 54.88% 57.24% 65.12% 60.80%
LoRA Labeled-Only 57.87% 64.10% 60.66% 66.79% 59.48% 72.05% 35.38% 42.16% 35.21% 50.80% 60.06% 64.28% 55.74%
Fixmatch 56.81% 68.56% 62.17% 72.78% 47.56% 77.41% 32.45% 39.09% 32.20% 42.61% 59.77% 52.60% 53.67%
FlexMatch 56.49% 66.97% 66.39% 71.90% 51.24% 78.40% 32.84% 38.67% 36.89% 75.35% 44.59% 54.99% 56.23%
SoftMatch 59.79% 68.03% 66.31% 70.87% 78.70% 83.65% 42.05% 42.93% 29.78% 71.20% 42.90% 60.48% 59.72%
FineSSL 63.14% 69.10% 56.25% 61.78% 74.17% 83.48% 35.03% 30.42% 33.55% 38.07% 53.73% 48.24% 53.91%
[…]
ST (PEFT) 63.35% 68.24% 65.61% 70.80% 67.52% 78.56% 36.42% 42.39% 38.27% 54.63% 60.34% 65.68% 59.32%
ST (PEFT + VFMs) 65.59% 71.65% 67.19% 72.83% 66.56% 77.21% 38.85% 51.06% 36.69% 58.49% 57.95% 62.31% 60.53%
AdaptFormer Labeled-Only 61.54% 64.95% 61.14% 68.23% 60.98% 72.83% 34.38% 38.99% 37.80% 48.63% 57.81% 59.07% 55.53%
Fixmatch 61.22% 67.34% 64.63% 71.19% 63.06% 80.49% 28.96% 42.41% 33.13% 44.76% 29.96% 57.24% 53.70%
FlexMatch 60.27% 68.78% 64.32% 70.73% 58.13% 78.08% 31.02% 35.93% 38.50% 52.21% 34.18% 54.15% 53.86%
SoftMatch 61.33% 69.31% 65.33% 70.17% 65.78% 82.67% 38.16% 42.66% 34.48% 56.08% 33.90% 55.13% 56.25%
FineSSL 61.91% 68.14% 58.37% 64.86% 73.95% 84.17% 25.27% 29.78% 25.97% 36.87% 42.19% 47.12% 51.55%
[…]
ST (PEFT) 63.78% 68.94% 66.78% 71.71% 69.54% 80.16% 37.14% 41.51% 38.45% 55.78% 60.48% 61.74% 59.67%
ST (PEFT + VFMs) 65.69% 71.76% 67.81% 73.23% 67.92% 78.43% 38.90% 51.22% 36.96% 59.01% 58.23% 63.29% 61.04%
DINOv2
Full Labeled-Only 46.44% 61.76% 47.58% 58.83% 5.22% 56.95% 34.24% 41.54% 26.47% 26.94% 53.87% 62.31% 43.51%
Fixmatch 49.57% 63.19% 42.98% 64.73% 20.97% 55.16% 26.96% 20.55% 16.75% 16.82% 31.08% 50.49% 38.27%
FlexMatch 44.04% 60.80% 44.97% 58.23% 24.71% 63.75% 23.78% 26.09% 17.14% 15.80% 40.37% 30.24% 37.49%
SoftMatch 27.02% 57.55% 48.73% 59.06% 51.52% 77.41% 26.28% 36.17% 16.54% 17.71% 16.17% 46.55% 40.06%
FineSSL 23.78% 48.40% 2.76% 16.57% 10.54% 22.05% 22.25% 21.17% 20.75% 23.52% 39.94% 35.44% 23.93%
[…]
ST (PEFT) 64.41% 69.84% 59.94% 66.87% 55.76% 69.62% 34.34% 56.55% 31.40% 55.02% 46.98% 64.98% 56.31%
ST (PEFT + VFMs) 66.97% 73.78% 65.53% 71.48% 67.32% 79.25% 36.55% 52.98% 35.93% 56.55% 55.98% 64.70% 60.59%
LoRA Labeled-Only 61.01% 68.40% 59.71% 66.78% 48.59% 63.08% 40.69% 48.02% 35.87% 61.72% 46.69% 71.73% 56.02%
Fixmatch 57.39% 68.40% 56.62% 71.38% 36.60% 66.65% 38.11% 41.16% 35.65% 40.99% 14.77% 42.90% 47.55%
FlexMatch 52.82% 65.37% 55.69% 67.65% 33.32% 72.63% 36.56% 35.81% 37.13% 66.85% 26.16% 33.61% 48.63%
SoftMatch 53.62% 64.04% 60.41% 67.34% 46.60% 79.54% 32.05% 44.48% 33.57% 65.07% 27.43% 27.85% 50.17%
FineSSL 66.97% 72.66% 54.62% 62.07% 69.78% 87.38% 30.99% 24.55% 24.77% 31.73% 31.93% 39.80% 49.77%
[…]
ST (PEFT) 66.65% 72.18% 58.57% 65.29% 55.79% 70.59% 40.99% 54.28% 33.11% 56.85% 42.62% 63.29% 56.68%
ST (PEFT + VFMs) 67.77% 74.10% 67.12% 73.24% 66.03% 77.03% 39.20% 52.48% 36.77% 59.25% 58.23% 63.71% 61.24%
AdaptFormer Labeled-Only 62.98% 67.61% 58.87% 67.03% 47.00% 57.41% 34.42% 53.07% 27.33% 41.79% 52.32% 49.79% 51.64%
Fixmatch 59.47% 70.37% 57.85% 69.68% 26.41% 62.30% 38.73% 45.55% 32.65% 38.89% 19.83% 50.35% 47.67%
FlexMatch 57.07% 65.90% 56.01% 67.53% 51.29% 64.02% 33.43% 40.20% 33.35% 42.16% 34.32% 27.85% 47.76%
SoftMatch 58.67% 66.17% 59.97% 66.65% 53.06% 78.06% 33.47% 32.17% 32.91% 45.00% 24.47% 30.94% 48.46%
FineSSL 64.04% 71.17% 55.96% 63.34% 62.32% 69.73% 17.66% 24.30% 25.79% 30.03% 31.79% 45.01% 46.76%
[…]
ST (PEFT) 67.18% 72.98% 63.47% 70.93% 56.21% 72.00% 40.18% 55.53% 33.05% 57.95% 45.15% 60.90% 57.96%
ST (PEFT + VFMs) 68.24% 74.15% 67.65% 73.43% 66.32% 77.98% 39.39% 52.53% 36.77% 59.97% 59.21% 61.74% 61.45%
Table 8: Complete results of the SSL baseline experiment with different SSL methods.
B.3 PEFT Experiment Details
As a completion of Figure 2, we present the detailed results of the PEFT experiment with different SSL methods in Table 9. We present the detailed results of the PEFT experiment with different SSL methods in Table 9.
DTD Sun397 Resisc45 Retinopathy Clevr-C KITTI
N3 N6 N3 N6 N1 N2 N40 N80 N10 N20 N5 N10
CLIP Labeled-Only Full 53.46% 62.82% 52.79% 59.49% 52.98% 65.51% 31.78% 32.69% 25.03% 29.62% 37.55% 65.54%
LoRA 57.87% 64.10% 60.66% 66.79% 59.48% 72.05% 35.38% 42.16% 35.21% 50.80% 60.06% 64.28%
AdaptFormer 61.54% 64.95% 61.14% 68.23% 60.98% 72.83% 34.38% 38.99% 37.80% 48.63% 57.81% 59.07%
$\Delta$ 6.25% 1.70% 8.11% 8.02% 7.25% 6.93% 3.10% 7.88% 11.47% 20.10% 21.38% -3.87%
FixMatch Full 46.97% 60.37% 45.49% 60.97% 31.76% 74.48% 24.64% 45.95% 21.70% 21.05% 45.43% 45.15%
LoRA 56.81% 68.56% 62.17% 72.78% 47.56% 77.41% 32.45% 39.09% 32.20% 42.61% 59.77% 52.60%
AdaptFormer 61.22% 67.34% 64.63% 71.19% 63.06% 80.49% 28.96% 42.41% 33.13% 44.76% 29.96% 57.24%
$\Delta$ 12.05% 7.58% 17.92% 11.01% 23.55% 4.48% 6.07% -5.20% 10.97% 22.63% -0.56% 9.77%
FlexMatch Full 50.69% 61.06% 44.41% 61.35% 41.62% 79.05% 28.35% 38.12% 25.31% 20.75% 44.87% 47.40%
LoRA 56.49% 66.97% 66.39% 71.90% 51.24% 78.40% 32.84% 38.67% 36.89% 75.35% 44.59% 54.99%
AdaptFormer 60.27% 68.78% 64.32% 70.73% 58.13% 78.08% 31.02% 35.93% 38.50% 52.21% 34.18% 54.15%
$\Delta$ 7.69% 6.81% 20.94% 9.97% 13.06% -0.81% 3.58% -0.82% 12.39% 43.03% -5.49% 7.17%
SoftMatch Full 52.29% 62.55% 50.69% 60.70% 59.86% 84.67% 32.91% 24.15% 19.53% 20.02% 31.22% 57.52%
LoRA 59.79% 68.03% 66.31% 70.87% 78.70% 83.65% 42.05% 42.93% 29.78% 71.20% 42.90% 60.48%
AdaptFormer 61.33% 69.31% 65.33% 70.17% 65.78% 82.67% 38.16% 42.66% 34.48% 56.08% 33.90% 55.13%
$\Delta$ 8.27% 6.12% 15.13% 9.82% 12.38% -1.51% 7.19% 18.65% 12.60% 43.62% 7.17% 0.28%
FineSSL Full 7.02% 17.18% 1.46% 4.01% 7.78% 8.37% 39.64% 16.97% 19.53% 26.08% 41.49% 44.59%
LoRA 63.14% 69.10% 56.25% 61.78% 74.17% 83.48% 35.03% 30.42% 33.55% 38.07% 53.73% 48.24%
AdaptFormer 61.91% 68.14% 58.37% 64.86% 73.95% 84.17% 25.27% 29.78% 25.97% 36.87% 42.19% 47.12%
$\Delta$ 55.51% 51.44% 55.85% 59.31% 66.29% 75.46% -9.49% 13.13% 10.23% 11.39% 6.47% 3.09%
DINOv2 Labeled-Only Full 46.44% 61.76% 47.58% 58.83% 5.22% 56.95% 34.24% 41.54% 26.47% 26.94% 53.87% 62.31%
LoRA 61.01% 68.40% 59.71% 66.78% 48.59% 63.08% 40.69% 48.02% 35.87% 61.72% 46.69% 71.73%
AdaptFormer 62.98% 67.61% 58.87% 67.03% 47.00% 57.41% 34.42% 53.07% 27.33% 41.79% 52.32% 49.79%
$\Delta$ 15.56% 6.25% 11.71% 8.07% 42.57% 3.29% 3.32% 9.00% 5.13% 24.82% -4.36% -1.55%
FixMatch Full 49.57% 63.19% 42.98% 64.73% 20.97% 55.16% 26.96% 20.55% 16.75% 16.82% 31.08% 50.49%
LoRA 57.39% 68.40% 56.62% 71.38% 36.60% 66.65% 38.11% 41.16% 35.65% 40.99% 14.77% 42.90%
AdaptFormer 59.47% 70.37% 57.85% 69.68% 26.41% 62.30% 38.73% 45.55% 32.65% 38.89% 19.83% 50.35%
$\Delta$ 8.86% 6.20% 14.25% 5.80% 10.54% 9.32% 11.45% 22.81% 17.40% 23.12% -13.78% -3.87%
FlexMatch Full 44.04% 60.80% 44.97% 58.23% 24.71% 63.75% 23.78% 26.09% 17.14% 15.80% 40.37% 30.24%
LoRA 52.82% 65.37% 55.69% 67.65% 33.32% 72.63% 36.56% 35.81% 37.13% 66.85% 26.16% 33.61%
AdaptFormer 57.07% 65.90% 56.01% 67.53% 51.29% 64.02% 33.43% 40.20% 33.35% 42.16% 34.32% 27.85%
$\Delta$ 10.90% 4.84% 10.89% 9.36% 17.59% 4.58% 11.21% 11.92% 18.10% 38.71% -10.13% 0.49%
SoftMatch Full 27.02% 57.55% 48.73% 59.06% 51.52% 77.41% 26.28% 36.17% 16.54% 17.71% 16.17% 46.55%
LoRA 53.62% 64.04% 60.41% 67.34% 46.60% 79.54% 32.05% 44.48% 33.57% 65.07% 27.43% 27.85%
AdaptFormer 58.67% 66.17% 59.97% 66.65% 53.06% 78.06% 33.47% 32.17% 32.91% 45.00% 24.47% 30.94%
$\Delta$ 29.12% 7.55% 11.47% 7.94% -1.69% 1.39% 6.48% 2.16% 16.70% 37.32% 9.77% -17.16%
FineSSL Full 23.78% 48.40% 2.76% 16.57% 10.54% 22.05% 22.25% 21.17% 20.75% 23.52% 39.94% 35.44%
LoRA 66.97% 72.66% 54.62% 62.07% 69.78% 87.38% 30.99% 24.55% 24.77% 31.73% 31.93% 39.80%
AdaptFormer 64.04% 71.17% 55.96% 63.34% 62.32% 69.73% 17.66% 24.30% 25.79% 30.03% 31.79% 45.01%
$\Delta$ 41.73% 23.51% 52.53% 46.13% 55.51% 56.51% 2.08% 3.25% 4.53% 7.36% -8.09% 6.96%
Table 9: Performance comparison of various SSL methods on our proposed benchmark datasets. For each dataset, model, and SSL baseline, we report the arithmetic mean of test accuracy across different PEFT, subtracted by the test accuracy of full fine-tuning. A positive value indicates that PEFT outperform full fine-tuning in the given scenario. Overall, the results show that PEFT consistently outperform full fine-tuning, with very few exceptions. This demonstrates the effectiveness of PEFT in enhancing the performance of VFMs within SSL frameworks.
Appendix C Hyperparameter Tuning Details
We provide the detailed hyperparameter tuning procedure in this section.
C.1 Hyperparameter Tuning Procedure
Concretely, given a set of hyperparameters $\{\phi_{1},\phi_{2},\ldots,\phi_{h}\}$, and models trained with them $\{f_{(\theta,\phi_{1})},f_{(\theta,\phi_{2})},\ldots,f_{(\theta,\phi_{h})}\}$, a held-out unlabled validation set $\mathcal{D}^{v}=\{(x^{v}_{i})\}_{i=1}^{N^{v}}$ is use to calculate the 7 unsupervised criteria. Then we calculate the rank of each criterion for each hyperparameter set. We will choose the hyperparameter set with the lowest average rank $\frac{1}{7}\{R\}$ across all criteria.
Problem Formulation Formally, the objective is to find a suitable function,
[…]
where $h$ serves as a feature extractor, and $g$ is a linear projector. The aim is to use $\mathcal{L}$ and $\mathcal{U}$ to learn $h$ and $g$ that minimize the generalization error on the test set $\mathcal{T}$.
For any dataset $\mathcal{D}\coloneqq{x_{i}}_{i=1}^{|\mathcal{D}|}$, regardless of whether it is labeled, we denote the output of $h$ as features, the output of $g$ as logits, and the argmax of the logits as predictions.
$\phi_{\mathcal{D}}$$\coloneqq\{h(x_{i})\}_{i=1}^{|\mathcal{D}|}\quad$ $(\text{features}),$
$\ell_{\mathcal{D}}$$\coloneqq\{g(h(x_{i}))\}_{i=1}^{|\mathcal{D}|}\quad$ $(\text{logits}),$
$\pi_{\mathcal{D}}$$\coloneqq\{\operatorname{arg\,max}g(h(x_{i}))\}_{i=1}^{|\mathcal{D}|}\quad$ $(\text{predictions}).$
Tuning without Labels The primary challenge in tuning hyperparameters within an unsupervised setting lies in effectively adjusting them without access to labels. Inspired by the approach for standardizing hyperparameter tuning in Unsupervised Domain Adaptation (UDA) (Ericsson et al., 2023), we propose tuning the hyperparameters using $\phi_{\mathcal{D}_{v}}$, $\ell_{\mathcal{D}_{v}}$ and $\pi_{\mathcal{D}_{v}}$. Specifically, we define function $v$ as a validator:
$v$ $\coloneqq(\phi_{\mathcal{D}_{v}},\ell_{\mathcal{D}_{v}},\pi_{\mathcal{D}_{v}})\rightarrow\mathbb{R}.$
who receives the features, logits, and predictions of the validation set $\mathcal{D}_{v}$ and outputs a scalar value. When given a set of hyperparameters $\{\theta_{1},\theta_{2},\ldots,\theta_{h}\}$, and models trained with them as $\{f_{\theta_{1}},f_{\theta_{2}},\ldots,f_{\theta_{h}}\}$. We first extract the features, logits, and predictions of the validation set $\mathcal{D}_{v}$ using each model,
$\phi_{\theta_{i}}$$\coloneqq\{h(x_{i})\}_{i=1}^{N^{v}},$
$\ell_{\theta_{i}}$$\coloneqq\{g(h(x_{i}))\}_{i=1}^{N^{v}},$
$\pi_{\theta_{i}}$$\coloneqq\{\operatorname{arg\,max}g(h(x_{i}))\}_{i=1}^{N^{v}},\forall i\in\{1,2,\ldots,h\}.$
Then, given a set of validators $\{v_{1},v_{2},\ldots,v_{n}\}$, we evaluate the performance of each model on the validation set $\mathcal{D}_{v}$ using the validators. Specifically, given extracted features, logits, and predictions of the validation set $\mathcal{D}_{v}$ for each model as:
$\{(\phi_{\theta_{1}},\ell_{\theta_{1}},\pi_{\theta_{1}}),(\phi_{\theta_{2}},\ell_{\theta_{2}},\pi_{\theta_{2}}),\ldots,(\phi_{\theta_{h}},\ell_{\theta_{h}},\pi_{\theta_{h}})\}$, we calculate the performance of each model using each validator as,
$v_{1}(\phi_{\theta_{1}},\ell_{\theta_{1}},\pi_{\theta_{1}}),$
$v_{2}(\phi_{\theta_{2}},\ell_{\theta_{2}},\pi_{\theta_{2}}),$
$\ldots$,… ,
$v_{n}(\phi_{\theta_{h}},\ell_{\theta_{h}},\pi_{\theta_{h}}).$
Suppose we have $n$ validators and $h$ models. We denote the performance of each model on the validation set $\mathcal{D}_{v}$ using each validator as a matrix $M$:
$M$ $\coloneqq\begin{bmatrix}v_{1}(\phi_{\theta_{1}},\ell_{\theta_{1}},\pi_{\theta_{1}})&\ldots&v_{n}(\phi_{\theta_{1}},\ell_{\theta_{1}},\pi_{\theta_{1}})\\ v_{1}(\phi_{\theta_{2}},\ell_{\theta_{2}},\pi_{\theta_{2}})&\ldots&v_{n}(\phi_{\theta_{2}},\ell_{\theta_{2}},\pi_{\theta_{2}})\\ \vdots&\ddots&\vdots\\ v_{1}(\phi_{\theta_{h}},\ell_{\theta_{h}},\pi_{\theta_{h}})&\ldots&v_{n}(\phi_{\theta_{h}},\ell_{\theta_{h}},\pi_{\theta_{h}})\\ \end{bmatrix}.$
Where each row represents the performance of a model under different validators. And each column represents the performance of different models under the same validator. Then across every column, we calculate the rank of each model based on the performance under the corresponding validator. We then get a rank matrix $R$:
$R$ $\coloneqq\begin{bmatrix}r_{11}&r_{12}&\ldots&r_{1h}\\ r_{21}&r_{22}&\ldots&r_{2h}\\ \vdots&\vdots&\ddots&\vdots\\ r_{n1}&r_{n2}&\ldots&r_{nh}\\ \end{bmatrix}.$
Here each row represents the rank of a model under different validators. Then we can approximate the performance of each model by calculating the average rank across all validators. We denote the average rank of each model as a vector $A$:
$A$ $\coloneqq\begin{bmatrix}\frac{1}{n}\sum_{j=1}^{n}r_{1j}\\ \frac{1}{n}\sum_{j=1}^{n}r_{2j}\\ \vdots\\ \frac{1}{n}\sum_{j=1}^{n}r_{hj}\\ \end{bmatrix}.$
To get the best performance model, we select the model with lowest value in $A$. We put the validators we use and more detailed examples in Section C.2.
C.2 Validators
We deploys 7 semi-supervised learning validators, including
RankMe Score (Garrido et al., 2023) We adapt RankMe Score to compute the feature matrix rank of the pre-trained data over both source and target data domains.
Adjusted Mutual Information (AMI) (Vinh et al., 2009) We adapt AMI to compute the adjusted mutual information between predicted and cluster labels of the validation set.
Adjusted Rand Index (ARI) (Halkidi et al., 2002) ARI is a widely used metric in cluster analysis by measuring the similarity between two clustering solutions. Given $RI$ is the Rank Index between the true and predicted clustering, ARI is defined by:
$ARI\coloneqq\frac{RI-\text{Expected RI}}{\max(RI)-\text{Expected RI}}$
V-Measure (Rosenberg & Hirschberg, 2007) Beside AMI, we also adapt V-Measure to measure the harmonic mean between homogeneity and completeness over the clustering labels and prediction.
Fowlkes-Mallows Index (FMI) (Fowlkes & Mallows, 1983) Measures the similarity between two clusterings, FMI is defined by:
$FMI\coloneqq\sqrt{\frac{TP}{TP+FP}\cdot\frac{TP}{TP+FN}}$
Calinski-Harabasz Index (CHI) (Caliński & Harabasz, 1974) CHI is an internal clustering measurement. Given $n$ sets of data point $\{x\}_{1}^{n}$ grouped into $k$ clusters $\{C\}_{1}^{k}$ with centroids $\{c\}_{i}^{k}$ and $\bm{c}$ is the overall centroid, CHI is defined by:
$CHI\coloneqq\frac{\left[\sum_{i=1}^{k}n_{i}||c_{i}-c||^{2}\right]/(k-1)}{\left[\sum_{i=1}^{k}\sum_{x\in C_{i}}||x-c_{i}||^{2}\right](n-k)}$
Batch Nuclear-norm Minimization (BNM) (Cui et al., 2020) An unsupervised domain adaptation method aims to generate predictions that favor both diversion and confidence; BNM maximizes the nuclear norm of the prediction matrix in a batch and repurposes it as a validation criterion.
C.3 Hyperparameter Tuning Results
DTD Sun397 Resisc45 Retinopathy Clevr-C KITTI
3 6 3 6 1 2 40 80 10 20 5 10
CLIP
LoRA Oracle Labeled-Only 59.04% 65.80% 61.55% 67.59% 59.48% 72.17% 40.10% 58.15% 36.07% 55.95% 67.51% 77.92%
Fixmatch 56.81% 68.56% 62.17% 72.78% 51.41% 79.51% 35.59% 44.09% 37.11% 49.53% 59.77% 61.88%
FlexMatch 59.68% 66.97% 66.39% 72.68% 57.08% 82.76% 37.29% 41.95% 36.89% 75.35% 52.88% 63.15%
SoftMatch 59.79% 68.03% 66.31% 70.87% 78.70% 83.65% 42.05% 43.50% 36.30% 71.20% 50.35% 65.12%
FineSSL 63.14% 69.10% 56.25% 61.78% 74.17% 84.73% 35.03% 31.58% 33.55% 40.26% 54.85% 51.48%
RankMe Labeled-Only 52.45% 62.18% 58.29% 57.04% 55.71% 59.73% 30.27% 48.43% 35.21% 44.01% 60.06% 64.28%
Fixmatch 51.54% 59.20% 51.17% 62.03% 35.25% 62.37% 33.01% 44.09% 32.20% 42.61% 59.77% 52.60%
FlexMatch 59.68% 66.33% 65.36% 52.01% 51.24% 82.76% 31.82% 39.90% 30.94% 27.98% 44.59% 55.41%
SoftMatch 58.99% 64.79% 63.07% 56.18% 62.16% 78.25% 42.05% 42.93% 29.78% 28.81% 33.47% 60.48%
FineSSL 58.51% 64.89% 54.26% 58.17% 57.60% 73.48% 16.95% 22.64% 23.93% 38.07% 48.52% 51.48%
Random Labeled-Only 33.14% 43.03% 35.59% 47.73% 26.18% 36.69% 29.67% 37.88% 23.73% 30.81% 48.38% 51.78%
Fixmatch 54.34% 64.57% 58.07% 68.34% 44.74% 73.10% 33.68% 41.39% 30.48% 40.82% 48.95% 52.46%
FlexMatch 52.00% 57.62% 58.07% 65.53% 52.38% 79.52% 33.98% 40.17% 30.59% 48.37% 46.23% 57.85%
SoftMatch 53.56% 58.12% 60.11% 64.99% 67.42% 81.32% 36.93% 42.32% 30.21% 49.76% 42.24% 59.21%
FineSSL 46.61% 51.06% 52.82% 57.12% 47.60% 80.56% 26.31% 28.22% 29.29% 37.48% 52.37% 50.26%
Ours Labeled-Only 57.87% 64.10% 60.66% 66.79% 59.48% 72.05% 35.38% 42.16% 35.21% 50.80% 60.06% 64.28%
Fixmatch 56.81% 68.56% 62.17% 72.78% 47.56% 77.41% 32.45% 39.09% 32.20% 42.61% 59.77% 52.60%
FlexMatch 56.49% 66.97% 66.39% 71.90% 51.24% 78.40% 32.84% 38.67% 36.89% 75.35% 44.59% 54.99%
SoftMatch 59.79% 68.03% 66.31% 70.87% 78.70% 83.65% 42.05% 42.93% 29.78% 71.20% 42.90% 60.48%
FineSSL 63.14% 69.10% 56.25% 61.78% 74.17% 83.48% 35.03% 30.42% 33.55% 38.07% 53.73% 48.24%
AdaptFormer Oracle Labeled-Only 61.54% 67.50% 61.14% 68.26% 60.98% 72.83% 37.98% 46.94% 37.80% 48.82% 57.81% 59.07%
Fixmatch 61.22% 67.50% 64.63% 71.19% 63.06% 80.49% 38.88% 46.37% 34.27% 46.41% 55.13% 60.34%
FlexMatch 60.27% 68.78% 64.32% 71.78% 58.13% 78.08% 37.67% 57.29% 38.50% 54.54% 51.05% 58.37%
SoftMatch 61.33% 69.31% 65.33% 70.17% 65.81% 82.67% 38.16% 42.66% 36.66% 56.08% 47.12% 57.24%
FineSSL 61.91% 68.14% 61.97% 67.83% 73.95% 84.17% 26.55% 34.22% 25.97% 36.87% 54.15% 52.60%
RankMe Labeled-Only 56.86% 65.43% 61.14% 67.35% 52.98% 70.43% 31.02% 39.28% 30.43% 41.95% 55.13% 56.68%
Fixmatch 61.22% 67.34% 50.47% 64.62% 46.84% 74.46% 38.88% 42.41% 34.27% 44.76% 29.96% 60.34%
FlexMatch 59.95% 67.13% 38.34% 49.39% 50.10% 73.56% 37.67% 38.30% 38.50% 43.43% 34.18% 54.15%
SoftMatch 59.52% 67.50% 47.72% 55.94% 54.94% 80.40% 38.16% 38.04% 36.66% 51.21% 45.15% 55.13%
61.91 FineSSL 61.91% 68.14% 58.37% 64.86% 73.95% 84.17% 25.27% 29.78% 25.97% 36.87% 42.19% 47.12%
Random Labeled-Only 31.40% 41.17% 37.04% 47.87% 22.73% 34.31% 21.31% 26.43% 20.56% 26.57% 44.75% 45.08%
Fixmatch 57.04% 64.02% 59.29% 68.67% 54.72% 74.62% 34.25% 42.70% 31.64% 43.01% 40.18% 58.27%
FlexMatch 51.91% 58.12% 55.59% 63.96% 55.30% 75.92% 34.53% 43.84% 33.32% 50.06% 43.08% 55.98%
SoftMatch 52.50% 58.71% 58.19% 64.76% 62.17% 81.61% 34.45% 40.74% 33.64% 48.07% 42.05% 55.51%
FineSSL 44.33% 48.28% 54.62% 59.75% 66.98% 77.64% 25.80% 31.07% 24.75% 33.32% 48.57% 49.41%
Ours Labeled-Only 61.54% 64.95% 61.14% 68.23% 60.98% 72.83% 34.38% 38.99% 37.80% 48.63% 57.81% 59.07%
Fixmatch 61.22% 67.34% 64.63% 71.19% 63.06% 80.49% 28.96% 42.41% 33.13% 44.76% 29.96% 57.24%
FlexMatch 60.27% 68.78% 64.32% 70.73% 58.13% 78.08% 31.02% 35.93% 38.50% 52.21% 34.18% 54.15%
SoftMatch 61.33% 69.31% 65.33% 70.17% 65.78% 82.67% 38.16% 42.66% 34.48% 56.08% 33.90% 55.13%
FineSSL 61.91% 68.14% 58.37% 64.86% 73.95% 84.17% 25.27% 29.78% 25.97% 36.87% 42.19% 47.12%
DinoV2
LoRA Oracle Labeled-Only 61.01% 69.04% 59.85% 66.84% 49.62% 63.08% 52.14% 53.40% 37.31% 61.72% 54.43% 71.73%
Fixmatch 57.39% 68.40% 58.72% 71.38% 36.60% 66.65% 38.11% 44.67% 35.65% 48.18% 37.27% 48.80%
FlexMatch 56.01% 65.37% 55.69% 67.65% 40.06% 72.63% 36.56% 41.06% 37.13% 66.85% 28.27% 35.58%
SoftMatch 53.62% 65.59% 60.41% 67.34% 54.79% 81.46% 38.64% 44.48% 36.59% 65.07% 27.43% 43.60%
FineSSL 66.97% 72.66% 56.14% 62.07% 69.78% 87.38% 45.23% 28.80% 27.60% 34.03% 46.55% 49.65%
RankMe Labeled-Only 46.76% 63.56% 55.34% 60.97% 42.22% 43.33% 33.35% 47.85% 31.75% 39.69% 46.69% 56.40%
Fixmatch 49.41% 64.84% 56.62% 59.04% 16.59% 50.78% 36.70% 36.13% 35.65% 28.30% 13.92% 42.90%
FlexMatch 56.01% 63.83% 55.69% 50.05% 33.32% 72.14% 35.86% 28.86% 32.84% 33.20% 28.27% 33.61%
SoftMatch 51.38% 64.04% 56.45% 53.54% 46.60% 75.76% 32.48% 36.66% 33.57% 31.61% 20.25% 29.54%
FineSSL 42.18% 45.53% 51.60% 58.65% 57.90% 70.95% 30.99% 24.55% 19.40% 32.33% 31.93% 39.80%
Random Labeled-Only 50.35% 60.99% 54.79% 61.49% 46.11% 78.92% 34.39% 41.27% 32.14% 47.64% 23.25% 33.66%
Fixmatch 54.06% 64.18% 55.65% 67.00% 23.78% 56.07% 36.01% 40.66% 31.43% 39.16% 21.99% 38.07%
FlexMatch 51.44% 60.71% 50.92% 61.35% 33.56% 66.64% 33.26% 35.24% 30.90% 48.62% 24.14% 32.82%
SoftMatch 50.35% 60.99% 54.79% 61.49% 46.11% 78.92% 34.39% 41.27% 32.14% 47.64% 23.25% 33.66%
FineSSL 58.30% 63.28% 54.12% 60.55% 45.20% 79.85% 34.20% 24.00% 23.92% 32.70% 39.10% 44.49%
Ours Labeled-Only 61.01% 68.40% 59.71% 66.78% 48.59% 63.08% 40.69% 48.02% 35.87% 61.72% 46.69% 71.73%
Fixmatch 57.39% 68.40% 56.62% 71.38% 36.60% 66.65% 38.11% 41.16% 35.65% 40.99% 14.77% 42.90%
FlexMatch 52.82% 65.37% 55.69% 67.65% 33.32% 72.63% 36.56% 35.81% 37.13% 66.85% 26.16% 33.61%
SoftMatch 53.62% 64.04% 60.41% 67.34% 46.60% 79.54% 32.05% 44.48% 33.57% 65.07% 27.43% 27.85%
FineSSL 66.97% 72.66% 54.62% 62.07% 69.78% 87.38% 30.99% 24.55% 24.77% 31.73% 31.93% 39.80%
AdaptFormer Oracle Labeled-Only 63.09% 69.63% 59.83% 67.03% 48.05% 63.86% 43.25% 59.78% 34.73% 46.51% 52.46% 58.23%
Fixmatch 59.47% 70.37% 60.41% 69.68% 27.56% 62.30% 38.73% 45.55% 32.65% 39.25% 33.33% 50.35%
FlexMatch 57.07% 65.90% 56.01% 67.53% 51.29% 64.02% 33.87% 42.18% 33.35% 47.67% 34.32% 40.51%
SoftMatch 58.67% 66.17% 59.97% 66.65% 53.06% 78.06% 34.65% 40.80% 33.37% 46.55% 25.60% 37.97%
FineSSL 66.38% 71.17% 56.17% 63.34% 62.38% 70.19% 28.75% 31.61% 25.79% 30.03% 31.79% 48.52%
RankMe Labeled-Only 58.51% 4.73% 56.57% 60.78% 44.71% 54.22% 35.82% 37.82% 29.67% 40.65% 52.32% 55.41%
Fixmatch 51.65% 66.28% 45.86% 59.07% 27.56% 58.54% 38.73% 45.55% 26.99% 29.78% 19.83% 50.35%
FlexMatch 52.13% 62.34% 41.94% 51.05% 45.79% 62.10% 27.50% 33.58% 25.49% 27.55% 34.32% 27.85%
SoftMatch 47.71% 61.12% 43.25% 53.20% 49.71% 70.83% 30.02% 40.34% 26.05% 28.76% 23.91% 29.54%
FineSSL 42.87% 45.32% 53.55% 60.31% 62.32% 69.73% 17.66% 23.04% 25.79% 28.73% 31.65% 46.41%
Random Labeled-Only 34.69% 45.25% 38.53% 49.57% 20.67% 29.72% 29.12% 35.31% 19.95% 25.34% 42.13% 45.34%
Fixmatch 55.55% 65.37% 54.71% 65.96% 23.97% 58.65% 35.31% 42.21% 29.48% 35.97% 23.58% 34.74%
FlexMatch 53.10% 61.76% 51.12% 61.99% 43.72% 60.70% 31.60% 38.66% 30.46% 39.13% 28.50% 31.60%
SoftMatch 52.34% 61.68% 52.93% 60.77% 51.27% 74.91% 32.71% 37.77% 30.78% 40.10% 24.66% 32.82%
FineSSL 57.77% 62.41% 55.23% 61.34% 60.01% 67.41% 24.13% 26.32% 25.39% 29.44% 31.22% 46.65%
Ours Labeled-Only 62.98% 67.61% 58.87% 67.03% 47.00% 57.41% 34.42% 53.07% 27.33% 41.79% 52.32% 49.79%
Fixmatch 59.47% 70.37% 57.85% 69.68% 26.41% 62.30% 38.73% 45.55% 32.65% 38.89% 19.83% 50.35%
FlexMatch 57.07% 65.90% 56.01% 67.53% 51.29% 64.02% 33.43% 40.20% 33.35% 42.16% 34.32% 27.85%
SoftMatch 58.67% 66.17% 59.97% 66.65% 53.06% 78.06% 33.47% 32.17% 32.91% 45.00% 24.47% 30.94%
FineSSL 64.04% 71.17% 55.96% 63.34% 62.32% 69.73% 17.66% 24.30% 25.79% 30.03% 31.79% 45.01%
Table 10: Complete hyperparameter tuning results for all datasets, models, and SSL baselines. For single validators we only include RankMe.

### v2

@@
 Abstract
-Semi-supervised learning (SSL) leverages abundant unlabeled data alongside limited labeled data to enhance learning.
-As vision foundation models (VFMs) increasingly serve as the backbone of vision applications, it remains unclear how SSL interacts with these pre-trained models.
-To address this gap, we develop new SSL benchmark datasets where frozen VFMs underperform and systematically evaluate representative SSL methods.
-We make a surprising observation: parameter-efficient fine-tuning (PEFT) using only labeled data often matches SSL performance, even without leveraging unlabeled data.
-This motivates us to revisit self-training, a conceptually simple SSL baseline, where we use the supervised PEFT model to pseudo-label unlabeled data for further training.
-To overcome the notorious issue of noisy pseudo-labels, we propose ensembling multiple PEFT approaches and VFM backbones to produce more robust pseudo-labels.
-Empirical results validate the effectiveness of this simple yet powerful approach, providing actionable insights into SSL with VFMs and paving the way for more scalable and practical semi-supervised learning in the era of foundation models.
-[^1] Our code is available at https://github.com/OSU-MLB/SSL-Foundation-Models.
-semi-supervised learning, self-training, vision foundation models, foundation models, benchmark, parameter-efficient fine-tuning
+Semi-supervised learning (SSL) enhances model performance by leveraging abundant unlabeled data alongside limited labeled data.
+As vision foundation models (VFMs) become central to modern vision applications, this paper revisits SSL in the context of these powerful pre-trained models.
+We conduct a systematic study on tasks where frozen VFMs underperform and reveal several key insights when fine-tuning them.
+First, parameter-efficient fine-tuning (PEFT) using only labeled data often surpasses traditional SSL methods—even without access to unlabeled data.
+Second, pseudo-labels generated by PEFT models offer valuable supervisory signals for unlabeled data, and different PEFT techniques yield complementary pseudo-labels.
+These findings motivate a simple yet effective SSL baseline for the VFM era: ensemble pseudo-labeling across diverse PEFT methods and VFM backbones.
+Extensive experiments validate the effectiveness of this approach, offering actionable insights into SSL with VFMs and paving the way for more scalable and robust semi-supervised learning in the foundation model era.
+footnotetext: *Equal contribution.footnotetext: Our code is available at https://github.com/OSU-MLB/SSL-Foundation-Models.
 1 Introduction
-The quality of machine learning models is often closely tied to the amount of labeled data available, but annotation can be costly or labor-intensive.
-As a result, semi-supervised learning (SSL), which leverages abundant unlabeled data alongside limited labeled data, has emerged as a promising paradigm for developing ML models without the need for extensive labeling (Zhu, 2008; Chapelle et al., 2006; Van Engelen & Hoos, 2020; Yang et al., 2022).
+The quality of machine learning (ML) models is often closely tied to the amount of labeled data available, but annotation can be costly or labor-intensive.
+Semi-supervised learning (SSL), which leverages abundant unlabeled data alongside limited labeled data, has thus emerged as a promising paradigm for developing ML models without the need for extensive labeling zhu2008semi ; chapelle2006semi ; van2020survey ; yang2022survey .
 Over the past few decades, numerous SSL algorithms have been developed to advance this field.
-Exemplar methods from the deep learning era include Mean Teacher (Tarvainen & Valpola, 2018), MixMatch (Berthelot et al., 2019), and FixMatch (Sohn et al., 2020), which dynamically impose objective functions on unlabeled data based on distillation and consistency, thereby enhancing learning performance.
-It is worth noting that many of these methods were originally designed to train neural networks from scratch.
-Meanwhile, pre-training on external labeled or unlabeled data has become the de facto standard across many machine learning application domains (Gui et al., 2024).
-For example, in computer vision, many recent algorithms are built upon vision foundation models (VFMs) such as CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2024).
-These models were pre-trained on massive datasets—–millions, if not billions, of data points.
-As a result, they have demonstrated remarkable generalizability across a wide range of tasks, often requiring only slight fine-tuning or, in some cases, functioning effectively as a frozen backbone (Zhang et al., 2024).
-Figure 1: Illustration of our proposed SSL pipeline.
-To effectively leverage abundant unlabeled data alongside scarce labeled data in the era of VFMs, our approach follows four phases: (1) Supervised Parameter-Efficient Fine-Tuning, where we harness labeled data by fine-tuning pre-trained VFMs using various PEFT algorithms; (2) Pseudo-Label Generation, where we exploit fine-tuned VFMs’ generalization ability to generate pseudo-labels for unlabeled data; (3) Pseudo-Label Ensemble, where we enhance robustness by aggregating pseudo-labels from multiple fine-tuned VFMs; and (4) Self-Training, where we consolidate all knowledge into one model.
+Exemplar methods from the deep learning era include Mean Teacher tarvainen2018meanteachersbetterrole , MixMatch berthelot2019mixmatch , and FixMatch sohn2020fixmatch , which dynamically impose objective functions on unlabeled data based on distillation and consistency, thereby enhancing learning performance.
+It is worth noting that many of these methods were originally designed to train neural networks “from scratch.”
+Figure 1: Left: The Venn diagram of the top 20% highest-confidence predictions from various VFMs and PEFT under DTD 3-shot, illustrating their intrinsic property of producing a diverse range of high-confidence predictions.
+Right: Ensembling these diverse pseudo-label predictions progressively boosts downstream performance for increasingly more ensembles (Self-Training → PETV-PET), highlighting the quality improvements from diversity.
+Results are averaged across 12 settings in our benchmark.
+Recently, pre-training on external labeled or unlabeled data has become the de facto standard across many machine learning application domains gui2024survey .
+For example, in computer vision, many recent algorithms are built upon vision foundation models (VFMs) such as CLIP radford2021learningtransferablevisualmodels and DINOv2 oquab2024dinov2learningrobustvisual .
+These models were pre-trained on massive datasets—millions, if not billions, of data points.
+As a result, they have demonstrated remarkable generalizability across a wide range of tasks, often requiring only slight fine-tuning or, in some cases, functioning effectively as a frozen backbone zhang2024vision .
 Given the promising advancements in both fields, we explore their interaction in this paper.
@@
 Finally, can we leverage the power of VFMs to develop more effective, yet simpler, SSL algorithms?
-To this end, we introduce new SSL benchmark datasets based on the Visual Task Adaptation Benchmark (VTAB) (Zhai et al., 2020), a diverse suite of classification tasks designed to evaluate visual representations.
+Study design.
+To this end, we introduce new SSL benchmark datasets based on the Visual Task Adaptation Benchmark (VTAB) zhai2020largescalestudyrepresentationlearning , a diverse suite of classification tasks designed to evaluate visual representations.
 Our focus is on tasks where frozen VFMs underperform, indicating the need for further fine-tuning, and where SSL could offer a beneficial solution.
-We then systematically evaluate three recent representative SSL methods—FixMatch (Sohn et al., 2020), FlexMatch (Zhang et al., 2022), SoftMatch (Chen et al., 2023) and FineSSL (Gan & Wei, 2024).
-Hyperparameters are carefully selected using techniques proposed for unsupervised domain adaptation (Ericsson et al., 2023), ensuring that the data leakage issue discussed in (Oliver et al., 2018) is avoided.
-Our empirical results show that SSL does not offer a significant advantage over fully fine-tuning VFMs with limited labeled data.
-In some cases, accuracy even decreases.
-We hypothesize that allowing SSL to update all parameters in VFMs may inadvertently reduce their built-in generalizability.
-Indeed, in practical applications of VFMs, parameter-efficient fine-tuning (PEFT)—which updates only a small fraction of parameters or introduces a lightweight learnable module while keeping the VFM largely frozen—often outperforms full fine-tuning, particularly when downstream labeled data is limited (Mai et al., 2024; Xin et al., 2024).
-When we integrate PEFT into SSL, performance improves significantly.
-However, even when PEFT alone is used to fine-tune VFMs, the performance gains often match or exceed those by SSL, all without requiring unlabeled data.
-These findings convey two key messages.
-First, new SSL algorithms specifically designed for VFMs are needed.
-Second, the fact that supervised PEFT matches SSL accuracy suggests it can already generate high-quality pseudo-labels for unlabeled data.
-Appropriately leveraging these pseudo-labels could further improve model performance.
-Building on these insights, we revisit self-training, a conceptually simple SSL baseline that uses the current model to pseudo-label unlabeled data for further training (Lee et al., 2013; Xie et al., 2020; Chen et al., 2020; Yalniz et al., 2019).
-A key challenge in self-training is ensuring the reliability of pseudo-labels—–incorrect labels can reinforce model biases and degrade performance.
-Traditional approaches rely on confidence thresholds or consistency constraints to filter pseudo-labels (Sohn et al., 2020; Nguyen & Yang, 2023), but these methods are often heuristic and suboptimal, limiting self-training performance compared to more advanced SSL methods.
-To address this, we propose an alternative strategy that leverages the complementary behaviors of VFMs and PEFT.
-Different VFMs and PEFT algorithms often produce diverse predictions, even when their overall accuracies are  similar (Mai et al., 2024; Tong et al., 2024b).
-This inherent diversity enables us to harness ensembling techniques (Zhou, 2012).
-By aggregating predictions from multiple widely used VFM backbones fine-tuned with different PEFT methods—and explicitly compensating for their varied confidence distributions—we obtain significantly more robust pseudo-labels, reducing the need for explicit pseudo-label selection.
-As a result, our ensemble-based approach yields a more reliable yet simpler self-training pipeline (Figure 1) that effectively leverages labeled and unlabeled data as well as the strengths of VFMs and PEFT, leading to substantial improvements over existing SSL methods.
-We extensively validate our approach on new benchmark datasets, where it outperforms existing SSL methods across most tasks, establishing it as a simple, effective, and competitive SSL baseline in the era of foundation models.
-Further analyses provide key insights for future improvements.
+We then systematically evaluate three representative SSL methods—FixMatch sohn2020fixmatch , FlexMatch zhang2021flexmatch and SoftMatch chen2023softmatch .
+Hyperparameters are carefully selected using techniques proposed for unsupervised domain adaptation ericsson2023betterpracticesdomainadaptation , ensuring that the data leakage issue discussed in oliver2018realistic is avoided.
+Key insights.
+Our empirical results highlight two main findings.
+First, fine-tuning VFMs with representative SSL algorithms offers limited advantages over using only labeled data for fine-tuning.
+Second, parameter-efficient fine-tuning (PEFT) mai2024lessonslearnedunifyingempirical ; vpetlbench2024 —which updates a small subset of parameters or adds lightweight learnable modules while keeping the VFM largely frozen—consistently yields substantial performance gains, regardless of the learning paradigm.
+These observations imply two important directions.
+First, there is a need for SSL methods tailored specifically for VFMs.
+Second, since PEFT models trained only on labeled data already match the performance of standard SSL approaches, effectively leveraging their predictions—i.e., pseudo-labels—offers a promising path to further improve performance in semi-supervised settings.
+An SSL baseline in the VFM era.
+Building on these insights, we introduce a simple yet effective SSL approach that leverages VFMs as backbones.
+Our method is grounded in the principle of self-training—a straightforward SSL strategy in which the model generates pseudo-labels for unlabeled data to guide further training lee2013pseudo ; xie2020selftrainingnoisystudentimproves ; chen2020bigselfsupervisedmodelsstrong ; yalniz2019billionscalesemisupervisedlearningimage .
+While traditional self-training often struggles with low-quality pseudo-labels, we address this by exploiting two key properties of VFMs and PEFT methods: their strong initial performance and complementary behaviors.
+Specifically, as observed in mai2024lessonslearnedunifyingempirical ; tong2024eyes , different VFM backbones and PEFT methods frequently produce diverse predictions—even when their overall accuracies are similar.
+As shown in Figure 1, this diversity motivates ensembling predictions from multiple VFM-PEFT pairs zhou2012ensemble .
+By explicitly compensating for their varied confidence distributions, we obtain significantly more robust pseudo-labels—without requiring explicit filtering sohn2020fixmatch ; nguyen2023boostingsemisupervisedlearningbridging .
+The result is a simpler yet more reliable self-training pipeline (Figure 2) that effectively leverages both labeled and unlabeled data, achieving substantial gains over existing SSL methods.
+We extensively validate our approach, VFM-PEFT Ensemble Training (V-PET), on newly proposed benchmark datasets.
+V-PET consistently outperforms existing SSL methods across most tasks, including the recently proposed FineSSL gan2024erasingbiasfinetuningfoundation , which also builds on VFMs.
+These results establish PET as a simple, effective, and competitive SSL baseline for the foundation model era.
+Figure 2: Illustration of V-PET.
+To effectively leverage abundant unlabeled data alongside scarce labeled data in the era of VFMs, our approach follows four phases: (a) Supervised Parameter-Efficient Fine-Tuning, where we harness labeled data by fine-tuning pre-trained VFMs using various PEFT algorithms; (b) Pseudo-Label Generation, where we exploit fine-tuned VFMs’ generalization ability to generate pseudo-labels for unlabeled data; (c) Pseudo-Label Ensemble, where we enhance robustness by aggregating pseudo-labels from multiple fine-tuned VFMs; and (d) Self-Training, where we consolidate all knowledge into one model.
+Remark.
+Self-training amini2025self ; assran2022maskedsiamesenetworkslabelefficient , pseudo-labeling lee2013pseudo , and ensemble methods dong2020survey have been extensively studied in the ML literature, spanning both the pre-deep learning and deep learning eras.
+Our goal is not to compete with existing methods but to establish a simple yet effective semi-supervised learning baseline that harnesses their strengths while incorporating the unique properties of foundation models.
+We note that a key step in ensembling is obtaining multiple base learners that are diverse and equally performant.
+Our contribution lies in leveraging the complementary behaviors of multiple foundation models and PEFT methods—an approach tailored specifically for the foundation model era.
+Consistent with representative SSL studies berthelot2019mixmatch ; sohn2020fixmatch ; zhang2021flexmatch , our primary focus is on classification.
+Nonetheless, we believe that our insights are transferable to various tasks and hope they will inspire future research on other downstream SSL tasks, such as segmentation and detection.
+2 Related Works
+Semi-Supervised Learning (SSL).
+In recent years, many SSL methods have centered on generating and selecting reliable pseudo-labels sohn2020fixmatch ; berthelot2019mixmatch ; berthelot2020remixmatch .
+FixMatch sohn2020fixmatch uses a fixed confidence threshold, while FlexMatch zhang2021flexmatch adopts class-specific thresholds and SoftMatch chen2023softmatch applies a soft threshold to balance label quality and quantity.
+However, these approaches were designed for training from scratch, leaving open questions about their compatibility with VFMs.
+Another key paradigm is self-training xie2020selftrainingnoisystudentimproves ; lee2013pseudo , which (1) trains a teacher on labeled data, (2) generates pseudo-labels for unlabeled samples, and (3) trains a student on both.
+Ensuring pseudo-label reliability typically involves unsupervised pre-training chen2020bigselfsupervisedmodelsstrong , confidence thresholds, or consistency constraints sohn2020fixmatch ; nguyen2023boostingsemisupervisedlearningbridging .
+We instead leverage VFMs to produce more robust pseudo-labels, questioning whether scratch-oriented SSL algorithms remain effective in the foundation-model era.
+Transfer Learning & Self-Supervised Learning.
+Transfer learning zhuang2020comprehensivesurveytransferlearning —and specifically PEFT (or PETL) for efficiently adapting foundation models han2024parameterefficientfinetuninglargemodels —has long leveraged pretrained models to boost downstream tasks.
+In our approach, we fine-tune VFMs via PEFT on labeled data, then use the adapted models to generate pseudo-labels for unlabeled examples, thus bridging transfer learning with semi-supervised methods.
+A related paradigm, self-supervised learning, also exploits unlabeled data and often serves as a baseline in SSL comparisons chen2020bigselfsupervisedmodelsstrong ; assran2022maskedsiamesenetworkslabelefficient , but it assumes abundant unlabeled data—an assumption that may fail in low-resource settings.
+Vision Foundation Models.
+Vision Transformers (ViT) dosovitskiy2021imageworth16x16words pre-trained on massive amounts of data, has become indispensable to modern AI development awais2025foundation .
+These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks.
+For example, CLIP-ViT radford2021learningtransferablevisualmodels trained with millions of image-text pairs shows an unprecedented zero-shot capability, robustness to distribution shifts, and serves as the encoders for various powerful generative models rombach2022high ; liu2024visual .
+Meanwhile, DINOv2 oquab2024dinov2learningrobustvisual pre-trained with self-supervised objectives on extensive sets of well-curated images effectively captures fine-grained localization features.
+Moreover, it is increasingly becoming a common sense that strategically fusing multiple VFMs can unlock synergistic gains mai2024lessonslearnedunifyingempirical ; tong2024eyes , leading to state-of-the-art performance across tasks from visual question answering to object detection.
+Leveraging VFMs for Semi-Supervised Learning.
+With the growing prominence of VFMs, few recent studies have explored their usage in SSL.
+The most relevant to ours is FineSSL gan2024erasingbiasfinetuningfoundation , which specifically investigates the use of CLIP vision backbone in an SSL setting, accompanied by pseudo-labels refinement using a balanced margin softmax.
+The evaluation was limited to simple, small-scale datasets like CIFAR-10 krizhevsky2009learning , where frozen VFMs can already achieve high accuracy.
+A few other works constrained the study only to VFMs with zero-shot capabilities, such as the full CLIP model with both vision and text encoders menghini2024enhancingclipclipexploring ; zhang2024candidatepseudolabellearningenhancing .
+Our work, in contrast, extends the scope in both the diversity of VFMs and the evaluation benchmarks, establishing the first comprehensive study of SSL in the era of foundation models.
 Table 1: Comparison of supervised linear probing performance between popular SSL datasets (CIFAR, Food101) and our benchmark, where N denotes the number of labeled samples per class.
@@
 DINOv2 91.7% 88.1% 83.1% 66.7% 65.3% 52.4% 41.2% 30.5% 51.3%
-Our key contributions are as follows:
-  • We introduce new SSL benchmark datasets tailored for studying SSL in the context of foundation models.
-  • We conduct a systematic evaluation of representative SSL methods, offering valuable insights for future research.
-  • We propose a simple yet highly effective self-training-based SSL method leveraging the key properties of VFMs and PEFT, establishing a strong baseline for future work.
-Remark.
-Self-training (Amini et al., 2025), pseudo-labeling (Lee et al., 2013), and ensembling (Dong et al., 2020) have been extensively studied in the ML literature, spanning both the pre-deep learning and deep learning eras.
-Our goal is not to compete with existing methods but to establish a simple yet effective semi-supervised learning baseline that harnesses their strengths while incorporating the unique properties of foundation models.
-By maintaining a clean and accessible approach, we aim to facilitate future research.
-While we acknowledge that integrating more sophisticated techniques could further enhance performance, we deliberately focus on simplicity.
-A key step in ensembling is obtaining multiple base learners that are both diverse and equally performant.
-Our contribution lies in leveraging the complementary behaviors of multiple foundation models and PEFT methods—an approach tailored specifically for the foundation model era.
-2 Related Works
-Semi-Supervised Learning.
-In the past few years, numerous SSL algorithms have been developed, many of which focus on generating and selecting reliable pseudo-labels during the training process (Sohn et al., 2020; Berthelot et al., 2019, 2020).
-For example, FixMatch (Sohn et al., 2020) relies on a fixed threshold to select pseudo-labels, while more recent methods dynamically adjust this threshold as training proceeds.
-FlexMatch (Zhang et al., 2022) introduces class-specific thresholds to accommodate varying degrees of difficulty, SoftMatch (Chen et al., 2023) applies a soft confidence threshold to balance label quality and quantity.
-However, most existing approaches were originally designed to train models from scratch, leaving open questions about their compatibility with VFMs and the adjustment required for optimal performance.
-Another prominent direction in SSL is self-training (Xie et al., 2020; Lee et al., 2013), typically comprising three steps: (1) train a teacher model on labeled data, (2) generate pseudo-labels on unlabeled data using the teacher, and (3) train a student model on both labeled and pseudo data.
-A key challenge lies in ensuring the reliability of pseudo-labels.
-Conventional solutions employ unsupervised pre-training (Chen et al., 2020), confidence thresholds or consistency constraints (Sohn et al., 2020; Nguyen & Yang, 2023) to improve the pseudo-label quality.
-In contrast, we aim to leverage the power of VFMs to produce more robust pseudo-labels and turn self-training into a simple yet powerful baseline in the VFM era.
-Vision Foundation Models.
-Vision Transformers (ViT) (Dosovitskiy et al., 2021) pre-trained on massive amounts of data, has become indispensable to modern AI development (Awais et al., 2025).
-These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks.
-For example, CLIP-ViT (Radford et al., 2021) with millions of image-text pairs show an unprecedented zero-shot capability, robustness to distribution shifts and serve as the encoders for various power generative models (Rombach et al., 2022; Liu et al., 2024)).
-Meanwhile, DINOv2 (Oquab et al., 2024) pre-trained with self-supervised objectives on extensive sets of well-curated images effectively captures fine-grained localization features.
-Despite this growing prominence of VFMs, limited work has explored how to integrate them with SSL.
-Although a recent study marks an initial attempt (Gan & Wei, 2024), a comprehensive study of how to effectively leverage VMFs in SSL is still lacking.
 3 On Evaluation of SSL in the Era of VFMs
 3.1 Problem Definition
-We consider a $C$-class classification problem in SSL with a large unlabeled dataset $\mathcal{D}^{u}\coloneqq\{{\bm{x}}^{u}_{i}\}_{i=1}^{N^{u}}$ and a much smaller labeled dataset $\mathcal{D}^{l}\coloneqq\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{N^{l}}$, where ${\bm{x}}^{l}_{i}$ and ${\bm{x}}^{u}_{i}$ are training samples and $y^{l}_{i}$ is the ground-truth class label.
-$N^{u}$ and $N^{l}$ denote the unlabeled and labeled dataset size with $N^{u}\gg N^{l}$.
-SSL aims to learn a model $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ using $\mathcal{D}^{u}$ and $\mathcal{D}^{l}$.
-Unlike conventional SSL that initializes $\bm{\theta}$ randomly, our framework starts with a VFM (e.g., CLIP or DINOv2) and fine-tunes it for downstream tasks.
-3.2 A Diverse and Challenging SSL Benchmark
-Despite recent advancements in SSL, most studies (Gan & Wei, 2024; Chen et al., 2023; Zhang et al., 2022) still evaluate performance on classic datasets such as CIFAR-10/100 (Krizhevsky & Hinton, 2009), STL-10 (Coates et al., 2011), and Food101 (Bossard et al., 2014).
-These benchmarks were suitable when models were randomly initialized, but with the advent of VFMs, they have become far less challenging.
-As shown in Table 1, linear probing on top of frozen VFM backbones already achieves remarkable accuracy on these benchmarks, even with fairly limited labeled data.
-Moreover, these datasets primarily focus on natural images, providing a limited perspective on real-world scenarios where VFMs are deployed.
-Table 2: Summary of our benchmark. Rem. Sen.: Remote sensing; Recog.: Recognition; $N^{l}/N^{u}$: # labeled/unlabeled training data.
-Dataset Task Domain $N^{u}+N^{l}$ Classes $N^{l}/$Class
-DTD Recog. Textual $3,008$ $47$ $3,6$
-SUN397 Recog. Natural $49,601$ $973$,6363,63 , 6
-Resisc45 Recog. Rem. Sen. $20,160$ $45$ $1,2$
-Retinopathy Recog. Medical $36,825$ $5$ $40,80$
-Clevr-C Count Synthetic $56,000$ $8$ $10,20$
-KITTI Depth Auto Drive $5,416$ $4$ $5,10$
-To effectively evaluate SSL methods in the VFM era, we propose a new benchmark designed to capture the complexities of real-world applications across different tasks, domains, and dataset sizes.
-Due to these intricacies, frozen VFMs face much greater challenges in our benchmark, as shown in Table 1, reinforcing the necessity of SSL to unleash the potential of VFMs.
-Specifically, we select six datasets from VTAB (Zhai et al., 2020) covering recognition, depth estimation, and counting tasks, drawn from textual, natural, remote sensing, medical, synthetic, and autonomous driving domains.
-We also vary label-per-class settings to assess the robustness of SSL methods.
-We summarize the datasets in Table 2 and provide detailed information inAppendix A.
-We believe our benchmark provides a more diverse and comprehensive evaluation of SSL methods in the VFM era.
-3.3 A Fair Hyperparameter Tuning for SSL
-Hyperparameter tuning has long been a persistent challenge in SSL (Oliver et al., 2018), remaining ambiguous and unstandardized throughout much of the existing literature.
-Due to the labeled data scarcity, standard train-validation split is often infeasible.
+We consider a $C$-class classification problem in SSL with a large unlabeled dataset $\mathcal{U}\coloneqq\{{\bm{x}}^{u}_{i}\}_{i=1}^{|\mathcal{U}|}$ and a much smaller labeled dataset $\mathcal{L}\coloneqq\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{|\mathcal{L}|}$, where ${\bm{x}}^{l}_{i}$ and ${\bm{x}}^{u}_{i}$ are training samples and $y^{l}_{i}$ is the ground-truth class label.
+$|\mathcal{U}|$ and $|\mathcal{L}|$ denote the unlabeled and labeled dataset size with $|\mathcal{U}|\gg|\mathcal{L}|$.
+SSL aims to learn a model $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ using $\mathcal{U}$ and $\mathcal{L}$.
+Unlike conventional SSL that initializes $\bm{\theta}$ randomly, our framework starts with a VFM (e.g., , CLIP or DINOv2) and fine-tunes it for downstream tasks.
+3.2 A Comprehensive SSL Image Classification Benchmark
+Despite recent advancements in SSL, most studies gan2024erasingbiasfinetuningfoundation ; chen2023softmatch ; zhang2021flexmatch continue to evaluate on classic datasets such as CIFAR-10/100 krizhevsky2009learning , STL-10 pmlr-v15-coates11a , and Food101 bossard2014food .
+However, these benchmarks exhibit two key limitations: Diminishing difficulty under VFMs.
+Following prior PEFT-on-VFM work jie2022convolutionalbypassesbettervision ; jia2022visual ; tu2023visual ; mai2024lessonslearnedunifyingempirical , we use linear probing as our primary performance indicator.
+As shown in Table 1, linear probing on frozen VFM backbones already delivers remarkable accuracy—even with limited labeled data.
+We also report results for various classification-head sizes in section B.4, and Narrow domain coverage.
+Because these benchmarks focus mainly on natural images, they offer only a narrow view of real-world VFM applications.
+To effectively evaluate SSL methods in the VFM era—focusing exclusively on semi-supervised image classification—we propose a new benchmark designed to capture the complexities of real-world applications across different domains and dataset sizes.
+Dataset & Regime We choose six classification datasets from VTAB zhai2020largescalestudyrepresentationlearning , spanning remote sensing, medical imaging, synthetic data, and autonomous driving tasks.
+To test the robustness of SSL methods, we vary the number of labels per class and employ linear probing as our performance metric, selecting shot counts that keep each task sufficiently challenging for frozen VFMs.
+These settings impose substantial difficulty on frozen VFMs (Table 1), underscoring the importance of SSL for unlocking their full potential.
+A summary of the datasets appears in Table 2, with full details provided in Appendix A.
+Together, this benchmark offers a more diverse and comprehensive foundation-model-era evaluation of semi-supervised image classification.
+3.3 Fair Hyperparameter Tuning for SSL
+Hyperparameter tuning has long been a persistent challenge in SSL oliver2018realistic , remaining ambiguous and unstandardized throughout much of the existing literature.
+Due to the labeled data scarcity, a standard train-validation split is often infeasible.
 Tuning on a held-out labeled validation or even test set can lead to significant data leakage, resulting in over-optimistic results and unfair comparisons.
 To tackle this issue and complement our benchmark, we establish a more rigorous protocol for hyperparameter tuning in SSL.
+Table 2: Summary of our benchmark. Rem. Sen.: Remote sensing; Recog.: Recognition; $|\mathcal{L}|$: # labeled training data; $|\mathcal{U}|$: # unlabeled training data.
+Dataset Task Domain $|\mathcal{L}|+|\mathcal{U}|$ Classes $|\mathcal{L}|/$Class
+DTD Recog. Textual 3,008 47 3, 6
+SUN397 Recog. Natural 49,601 397 3, 6
+Resisc45 Recog. Rem. Sen. 20,160 45 1, 2
+Retinopathy Recog. Medical 36,825 5 4, 8
+Clevr-C Count Synthetic 56,000 8 1, 2
+KITTI Depth Auto Drive 5,416 4 5, 10
 Our core insight is to harness the defining characteristic of SSL—an abundance of unlabeled training data—to tune parameters in an unsupervised manner and avoid the pitfalls of data leakage.
-One recent study has explored such an idea for domain adaptation, using unsupervised criteria such as RankMe (Garrido et al., 2023) and AMI (Vinh et al., 2009) to estimate the effectiveness of each hyperparameter configuration (Ericsson et al., 2023).
-However, Ericsson et al.
-(2023) also revealed that no single criterion could reliably select suitable hyperparameters across all scenarios.
-Motivated by this finding, we propose integrating seven unsupervised criteria—five derived from features, AMI (Vinh et al., 2009), ARI (Halkidi et al., 2002), V-Measure (Rosenberg & Hirschberg, 2007), FMI (Fowlkes & Mallows, 1983) and BNM (Cui et al., 2020), and two from logits, RankMe (Garrido et al., 2023) and CHI (Caliński & Harabasz, 1974)—for more robust tuning.
-Concretely, for each hyperparameter configuration and its corresponding model, we compute all seven criteria using a held-out unlabeled validation set $\mathcal{D}^{v}\coloneqq\{{\bm{x}}^{v}_{i}\}_{i=1}^{N^{v}}$.
+One recent study has explored such an idea for domain adaptation, using unsupervised criteria such as RankMe garrido2023rankmeassessingdownstreamperformance and AMI vinh2009information to estimate the effectiveness of each hyperparameter configuration ericsson2023betterpracticesdomainadaptation .
+However, ericsson2023betterpracticesdomainadaptation also revealed that no single criterion could reliably select suitable hyperparameters across all scenarios.
+Motivated by this finding, we propose integrating seven unsupervised criteria—five derived from features, AMI vinh2009information , ARI halkidi2002cluster , V-Measure rosenberg2007v , FMI fowlkes1983method and BNM cui2020towards , and two from logits, RankMe garrido2023rankmeassessingdownstreamperformance and CHI calinski1974dendrite —for more robust tuning.
+Concretely, for each hyperparameter configuration and its corresponding model, we compute all seven criteria using a held-out unlabeled validation set $\mathcal{V}\coloneqq\{{\bm{x}}^{v}_{i}\}_{i=1}^{|\mathcal{V}|}$.
 Next, we rank every hyperparameter configuration based on each criterion and choose the one achieving the lowest average rank across all.
-We provide formal definitions and detailed procedures in Section C.1.
+We provide formal definitions and detailed procedures in section C.1 and the effectiveness of the proposed method in section B.7.
 By eliminating reliance on held-out labeled validation sets, our method mitigates data leakage and promotes a more practical and fair tuning protocol for SSL.
@@
 Evaluation Setup.
-We focus on two representative VFMs—ViT-B/16 CLIP (Radford et al., 2021) and ViT-B/14 DINOv2 (Oquab et al., 2024)—covering language-image contrastive pre-training and self-supervised pre-training strategies, respectively.
-We examine four representative SSL methods, including FixMatch (Sohn et al., 2020), FlexMatch (Zhang et al., 2022), SoftMatch (Chen et al., 2023), and FineSSL (Gan & Wei, 2024) and use labeled-only fine-tuning as the baseline.
-We employ the AdamW optimizer (Loshchilov & Hutter, 2019) with a batch size of 32 and weight decay $5\times 10^{-4}$ to fine-tune the model for 35 epochs.
+We focus on two representative VFMs—ViT-B/16 CLIP radford2021learningtransferablevisualmodels and ViT-B/14 DINOv2 oquab2024dinov2learningrobustvisual —covering language-image contrastive pre-training and self-supervised pre-training strategies, respectively.
+We examine four representative SSL methods, including FixMatch sohn2020fixmatch , FlexMatch zhang2021flexmatch , SoftMatch chen2023softmatch , and FineSSL gan2024erasingbiasfinetuningfoundation and use labeled-only fine-tuning as the baseline.
+We employ the AdamW optimizer Loshchilov2019Decoupled with a batch size of 32 and weight decay $5\times 10^{-4}$ to fine-tune the model for 35 epochs.
 Learning rates and other hyperparameters are tuned with our proposed tuning protocol.
 The classification performance is evaluated using the Top-1 accuracy on the test dataset.
-Figure 2: Average accuracy of SSL methods with full fine-tuning or PEFT (LoRA, AdaptFormer) across 12 settings in our benchmark.
-Under fair hyperparameter tuning, full fine-tuning with limited labeled data alone can surpass SSL methods.
-Although PEFT enhances SSL, PEFT with labeled data alone remains highly competitive, indicating limited benefit from unlabeled data in current SSL approaches when paired with VFMs.
-We present detailed results in Section B.3.
 Surprising Effectiveness of Labeled-only Fine-tuning.
@@
 In other words, even with a large amount of additional unlabeled data, SSL provides little advantage over fine-tuning VFMs with limited labeled data.
-This finding is reminiscent of observations made in a comprehensive investigation conducted seven years ago (Oliver et al., 2018).
+This finding is reminiscent of observations made in a comprehensive investigation conducted seven years ago oliver2018realistic .
 Due to the inherently noisy supervised signals associated with unlabeled data in SSL methods, we hypothesize that allowing SSL to update all parameters in VFMs may inadvertently reduce their built-in generalizability.
 Can PEFT Come to the Rescue?
-When labeled downstream data are scarce, recent research has shown that parameter-efficient fine-tuning (PEFT)—which updates only a small subset of parameters or introduces a lightweight learnable module to frozen VFMs—often outperforms full fine-tuning on VFMs (Mai et al., 2024; Xin et al., 2024; Hu et al., 2021; Chen et al., 2022; Jia et al., 2022a; Tu et al., 2023).
+When labeled downstream data are scarce, recent research has shown that parameter-efficient fine-tuning (PEFT)—which updates only a small subset of parameters or introduces a lightweight learnable module to frozen VFMs—often outperforms full fine-tuning on VFMs mai2024lessonslearnedunifyingempirical ; vpetlbench2024 ; hu2021loralowrankadaptationlarge ; chen2022adaptformeradaptingvisiontransformers ; jia2022visual ; tu2023visual .
 When considering SSL scenarios with limited labeled data, a natural question arises: is PEFT compatible with SSL?
-Although a recent study has made an initial attempt to apply PEFT in SSL (Gan & Wei, 2024), a comprehensive analysis of its compatibility with different VFM backbones, PEFT methods, and SSL approaches is still lacking.
-We, therefore, assess two commonly used PEFT strategies: LoRA (Hu et al., 2021), which trains additive low-rank matrices for Transformer layers to approximate weight updates, and AdaptFormer (Chen et al., 2022), an adapter-based approach proven effective in computer vision.
+Although a recent study has made an initial attempt to apply PEFT in SSL gan2024erasingbiasfinetuningfoundation , a comprehensive analysis of its compatibility with different VFM backbones, PEFT methods, and SSL approaches is still lacking.
+We, therefore, assess two commonly used PEFT strategies: LoRA hu2021loralowrankadaptationlarge , which trains additive low-rank matrices for Transformer layers to approximate weight updates, and AdaptFormer chen2022adaptformeradaptingvisiontransformers , an adapter-based approach proven effective in computer vision.
 We follow the same protocol outlined in section 3 to tune the hyperparameters of these PEFT methods.
@@
 To wrap up the discussion in section 4, we propose a new semi-supervised learning method tailored to vision foundation models.
-Specifically, we build on self-training (Triguero et al., 2015; Lee et al., 2013; Blum & Mitchell, 1998; Balcan et al., 2005; McClosky et al., 2006b, a), a conceptually simple SSL baseline that uses pseudo-labels from unlabeled data as additional supervision to improve the model.
+Specifically, we build on self-training triguero2015self ; lee2013pseudo ; blum1998combining ; balcan2005co ; mcclosky2006reranking ; mcclosky2006effective , a conceptually simple SSL baseline that uses pseudo-labels from unlabeled data as additional supervision to improve the model.
 In the following section, we first briefly review self-training.
 5.1 Self-training
-Let us denote the predicted posterior probability of class $c$ given ${\bm{x}}$ as $p(c|{\bm{x}};f_{\bm{\theta}})$.
-The core idea of self-training is to gradually assign pseudo-labels to unlabeled data $\mathcal{D}^{u}=\{{\bm{x}}^{u}_{i}\}_{i=1}^{N^{u}}$ based on high-confidence predictions
-$\hat{y}^{u}_{i}\coloneqq\operatorname{arg\,max}_{c}p(c|{\bm{x}}^{u}_{i};f_{\bm{\theta}})\hskip 20.0pt\text{ if }\max_{c}p(c|{\bm{x}}^{u}_{i};f_{\bm{\theta}})\geq\tau,$
-and add these pseudo-labeled data to the labeled set $\mathcal{D}^{l}=\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{N^{l}}$ for supervised learning.
+Let us denote the predicted posterior probability of class $c$ given ${\bm{x}}$ as $p(c\big{|}{\bm{x}};f({\bm{\theta}}))$.
+The core idea of self-training is to gradually assign pseudo-labels to unlabeled data $\mathcal{U}=\{{\bm{x}}^{u}_{i}\}_{i=1}^{|\mathcal{U}|}$ based on high-confidence predictions
+[…]
+and add these pseudo-labeled data to the labeled set =$\mathcal{L}=\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{|\mathcal{L}|}$ for supervised learning.
 Here, $\bm{\theta}$ represents the weights of the current classifier, and $\tau$ is the confidence threshold.
 If (a) high-confidence predictions are correct and (b) the updated model gradually increases its confidence in the remaining data, self-training can be as effective as fully supervised learning on the entire dataset with true labels.
-In our context, a VFM fine-tuned with PEFT using labeled $\mathcal{D}^{l}$ serves as the current model to generate pseudo-labels for $\mathcal{D}^{u}$, which are then used to further fine-tune the VFM.
+In our context, a VFM fine-tuned with PEFT using labeled $\mathcal{L}$ serves as the current model to generate pseudo-labels for $\mathcal{U}$, which are then used to further fine-tune the VFM.
 Challenges.
-While self-training is conceptually simple, its practical implementation remains challenging, primarily due to the selection of $\tau$.
-Ideally, $\tau$ should exclude samples with incorrect pseudo-labels.
-However, if set too high, it limits the number of pseudo-labeled samples available for supervised learning, reducing effectiveness.
-Conversely, if $\tau$ is too low, erroneous pseudo-labels may be included, reinforcing model biases.
-Further complicating the issue, the optimal $\tau$ can vary across self-training iterations.
-While extensive research has been conducted on pseudo-label selection (Amini et al., 2025; Yang et al., 2022), existing methods are often heuristic or overly complex, inadvertently making self-training impractical or unreliable.
-5.2 Our objective
+Although self-training is conceptually simple, choosing the confidence threshold $\tau$ is difficult.
+A high $\tau$ excludes wrongly labeled samples but leaves too few for learning; a low $\tau$ admits errors that reinforce bias.
+Moreover, the optimal $\tau$ shifts across iterations.
+Despite extensive study of pseudo-label selection amini2025self ; yang2022survey , most approaches remain heuristic or overly complex, undermining self-training’s practicality and reliability.
+Our objective.
 To leverage the high-quality pseudo-labels from label-only PEFT while circumventing the aforementioned challenges, we aim to develop a self-training-based SSL algorithm that is both simple and effective in practice.
@@
 5.2 Ensembling of multiple PEFT and VFMs
+Input :
+Labeled dataset $\mathcal{L}$ and unlabeled dataset $\mathcal{U}$
+PEFT methods indexed by $n\in[1,N]$
+VFMs indexed by $m\in[1,M]$
+Initialized parameters of VFM on PEFT ${\bm{\theta}}_{n,m}$
+Output : Optimal parameter $\bm{\theta}^{*}$
+// (a) Supervised Parameter Efficient Fine-Tuning
+for $n\in[1,N],\,m\in[1,M]$ do
+       Fine-tune ${\bm{\theta}}_{n,m}$ on $\mathcal{L}$ to obtain $\tilde{\bm{\theta}}_{n,m}$
+// (b) Pseudo-Label Generation
+for $n\in[1,N],\,m\in[1,M]$ do
+       Compute one-hot pseudo-label set $\mathcal{P}_{n,m}$ as
+[…]
+// (c) Pseudo-Label Ensemble
+Ensemble pseudo-label $\mathcal{P}\coloneqq\{\bar{p}_{i}\}_{i}^{|\mathcal{U}|}$, where:
+$\bar{p}_{i}=\frac{1}{N\times M}\sum_{\bar{n}\in[1,N],\bar{m}\in[1,M]}\mathcal{P}_{\bar{n},\bar{m}}[i]$
+// (d) Self-Training
+Choose $n^{\star}\in[1,N],\,m^{\star}\in[1,M]$ and fine-tune ${\bm{\theta}}_{n\star,m\star}$ on $\mathcal{P}$ to get $\bm{\theta}^{*}$
+Algorithm 1 V-PET (Figure 2)
 A necessary step toward achieving our objective is to ensure pseudo-labels of even higher quality to prevent error propagation.
-To this end, we employ ensembling techniques (Zhou, 2012; Dietterich, 2000), which combine the predictions from multiple models to enhance the overall quality and reliability of pseudo-labels.
+To this end, we employ ensembling techniques zhou2012ensemble ; dietterich2000ensemble , which combine the predictions from multiple models to enhance the overall quality and reliability of pseudo-labels.
 What to Ensemble?
-Unlike conventional methods that train multiple equally performant yet diverse base learners through bootstrapping, random initialization, or cyclic learning rate schedules (Huang et al., 2017), our approach exploits the unique properties of PEFT (Mai et al., 2024) and VFMs (Tong et al., 2024b, a).
+Unlike conventional methods that train multiple equally performant yet diverse base learners through bootstrapping, random initialization, or cyclic learning rate schedules huang2017snapshot , our approach exploits the unique properties of PEFT mai2024lessonslearnedunifyingempirical and VFMs tong2024eyes ; tongcambrian .
 Specifically, it has been observed that different PEFT methods, while achieving similar overall accuracy on downstream tasks, generate diverse predictions for individual samples.
@@
 How to Ensemble?
-Given $K$ models $f_{\bm{\theta}_{1}},\dots,f_{\bm{\theta}_{K}}$, several strategies can be applied to integrate their predictions for an unlabeled image ${\bm{x}}^{u}_{i}$.
-Two common approaches are: (a) Mean Logits, where the logits for each class are averaged across the models; and (b) Mean Probabilities, where the predicted probabilities for each class are averaged.
-The assumption behind these methods is that the models produce logits or probabilities at similar scales, which is typically the case when the models share similar architectures and are trained on bootstrapped data.
-In our context, since different VFMs were trained in distinct ways and adjusted using various PEFT methods, we find that the assumption of similar prediction scales often does not hold, as shown in Figure 3.
-As a result, some models may dominate the ensembling process, reducing its effectiveness.
+Common strategies for ensembling different predictions are (a) Mean Logits—averaging class logits, and (b) Mean Probabilities—averaging class probabilities, both assuming similar output scales (typical with identical architectures and bootstrapped training).
+However, VFMs trained differently and fine-tuned via varied PEFT methods produce inconsistent scales (Figure 4), causing some models to dominate and weakening the ensemble.
 To address this, we introduce a simple solution called Mean Labels.
-In this approach, we first obtain “hard” pseudo-labels from each model $\hat{y}_{i,k}^{u}=\operatorname{arg\,max}_{c}p(c|{\bm{x}}_{i}^{u};f_{\bm{\theta}_{k}})$, convert them into one-hot encoding, and average them to obtain a “soft” pseudo-label $\bar{{\bm{e}}}_{i}$.
-This ensures uniform prediction scales across all models (since the result is always one-hot) while preserving each model’s prediction.
-Figure 3: The average entropy of predicted probability distributions from different fine-tuned VFMs.
-This figure illustrates the gap in entropy among the pseudo-labels generated by different models, indicating that they are not well-calibrated.
-5.4 Overall Algorithm (Figure 1)
-We summarize our self-training-based SSL algorithm enhanced by ensembling multiple VFMs and PEFT as follows.
-Stage 1. Let $n$ denote a specific VFM and $m$ denote a specific PEFT, we fine-tune $\bm{\theta}_{n}$ on $\mathcal{D}^{l}$ to obtain $\bm{\theta}_{n,m}$.
-Stage 2.
-For an unlabeled image ${\bm{x}}_{i}^{u}\in\mathcal{D}^{u}$, we apply $\bm{\theta}_{n,m}$ to obtain its one-hot prediction ${\bm{e}}_{\hat{y}_{i,n,m}^{u}}\in\Delta^{C-1}$, where $\hat{y}_{i,n,m}^{u}=\operatorname{arg\,max}_{c}p(c|{\bm{x}}_{i}^{u};f_{\bm{\theta}_{n,m}})$.
-Stage 3. We average the one-hot encoding across all fine-tuned models to obtain the soft pseudo-label
-$\bar{{\bm{e}}}_{i}\coloneqq\cfrac{1}{N\times M}\sum_{n}\sum_{m}{\bm{e}}_{\hat{y}_{i,n,m}^{u}}\in\Delta^{C-1}.$ (1)
-Stage 4.
-Let $\mathcal{D}^{p}=\{({\bm{x}}^{u}_{i},\bar{{\bm{e}}}_{i})\}_{i=1}^{N^{u}}$ denote the pseudo-labeled dataset, we re-fine tune a targeted VFM $n\star$ by a targeted PEFT $m\star$ on $\mathcal{D}^{p}$, using the Kullback-Leibler divergence loss function.
-The resulting $\bm{\theta}_{n\star,m\star}$ is the final SSL model.
+We first obtain predictions from each model, convert them into one-hot encoding to ensure uniformity, and average them to obtain a soft pseudo-label.
+Finally, we generate the augmented dataset $\mathcal{P}$ using averaged soft pseudo-label and fine-tune PEFT-based VFM models.
+We present our proposed pipeline in algorithm 1.
 6 Experiment
-This section aims to answer the following questions: (1) How does the proposed method perform compared with the existing SSL methods?
-(2) Does the proposed method scale with more PEFT methods and VFMs?
-(3) How does the proposed ensemble method perform?
-(4) How effective is our proposed hyperparameter tuning protocol?
-We follow the setup described in Section 3 and Section 4.
+We conduct experiments to address four questions—(1) How does our method compare to existing SSL?
+(2) Does it scale with more PEFTs and VFMs?
+(3) How effective is the ensemble strategy?
+Following the setups in section 3 and section 4, we evaluate two variants: V-PET (ensembles over PEFT methods and VFMs) and PET (ensembles over PEFT methods on a single VFM).
+We use ST (self-training without ensembling) as a baseline and, in all cases, reinitialize from the original pretrained VFM before fine-tuning with pseudo-labels.
+Figure 3: Average SSL accuracy with full fine-tuning or PEFT across 12 settings shows that, with fair hyper-parameter tuning, fine-tuning on limited labels can outperform SSL; PEFT boosts SSL yet matches labeled-only performance, indicating minimal unlabeled-data benefit in current VFM-based SSL (see section B.3).
+Figure 4: The average entropy of predicted probability distributions from different fine-tuned VFMs, highlighting the entropy gap among pseudo-labels, indicating poor calibration.
+6.1 Performance Comparison
+We summarize the performance comparison between ST, PET, V-PET and existing SSL methods across all the 12 settings in subsection 6.1.
+Overall, V-PET achieve superior performance compared with others, as shown in the Average column.
+To further analyze the results, Figure 6 visualize the ranking frequency, where each matrix entry $(i,j)$ indicates how often method $i$ is ranked $j$-th across 12 settings.
+We then compute the mean rank (number in the bracket) of each method and sort them accordingly.
+Although V-PET does not claim first place in every single setting, it secures the top rank most frequently, showcasing the effectiveness of pseudo-label ensemble and establishing it as a simple yet powerful baseline in the VFM era.
 DTD Sun397 Resisc45 Retinopathy Clevr-C KITTI Average
@@
 […]
-ST (PEFT) 63.4% 68.2% 65.6% 70.8% 67.5% 78.6% 36.4% 42.4% 38.3% 54.6% 60.3% 65.7% 59.3%
-ST (PEFT + VFMs) 65.6% 71.7% 67.2% 72.8% 66.6% 77.2% 38.9% 51.1% 36.7% 58.5% 58.0% 62.3% 60.5%
+PET 63.4% 68.2% 65.6% 70.8% 67.5% 78.6% 36.4% 42.4% 38.3% 54.6% 60.3% 65.7% 59.3%
+V-PET 65.6% 71.7% 67.2% 72.8% 66.6% 77.2% 38.9% 51.1% 36.7% 58.5% 58.0% 62.3% 60.5%
 AdaptFormer Labeled Only 61.5% 65.0% 61.1% 68.2% 61.0% 72.8% 34.4% 39.0% 37.8% 48.6% 57.8% 59.1% 55.6%
@@
 […]
-ST (PEFT) 63.8% 68.9% 66.8% 71.7% 69.5% 80.2% 37.1% 41.5% 38.4% 55.8% 60.5% 61.7% 59.7%
-ST (PEFT + VFMs) 65.7% 71.8% 67.8% 73.2% 67.9% 78.4% 38.9% 51.2% 37.0% 59.0% 58.2% 63.3% 61.0%
+PET 63.8% 68.9% 66.8% 71.7% 69.5% 80.2% 37.1% 41.5% 38.4% 55.8% 60.5% 61.7% 59.7%
+V-PET 65.7% 71.8% 67.8% 73.2% 67.9% 78.4% 38.9% 51.2% 37.0% 59.0% 58.2% 63.3% 61.0%
 DINOv2
@@
 […]
-ST (PEFT) 66.7% 72.2% 58.6% 65.3% 55.8% 70.6% 41.0% 54.3% 33.1% 56.9% 42.6% 63.3% 56.7%
-ST (PEFT + VFMs) 67.8% 74.1% 67.1% 73.2% 66.0% 77.0% 39.2% 52.5% 36.8% 59.3% 58.2% 63.7% 61.2%
+PET 66.7% 72.2% 58.6% 65.3% 55.8% 70.6% 41.0% 54.3% 33.1% 56.9% 42.6% 63.3% 56.7%
+V-PET 67.8% 74.1% 67.1% 73.2% 66.0% 77.0% 39.2% 52.5% 36.8% 59.3% 58.2% 63.7% 61.2%
 AdaptFormer Labeled Only 63.0% 67.6% 58.9% 67.0% 47.0% 57.4% 34.4% 53.1% 27.3% 41.8% 52.3% 49.8% 51.6%
@@
 […]
-ST (PEFT) 67.2% 73.0% 63.5% 70.9% 56.2% 72.0% 40.2% 55.5% 33.1% 58.0% 45.2% 60.9% 58.0%
-ST (PEFT + VFMs) 68.2% 74.2% 67.7% 73.4% 66.3% 78.0% 39.4% 52.5% 36.8% 60.0% 59.2% 61.7% 61.5%
+PET 67.2% 73.0% 63.5% 70.9% 56.2% 72.0% 40.2% 55.5% 33.1% 58.0% 45.2% 60.9% 58.0%
+V-PET 68.2% 74.2% 67.7% 73.4% 66.3% 78.0% 39.4% 52.5% 36.8% 60.0% 59.2% 61.7% 61.5%
 Table 3: Performance comparison of baselines, existing SSL approaches, and our proposed methods on six diverse datasets (12 settings).
@@
 We report the full results in section B.2.
-Figure 4: Incorporating more diverse pseudo-labels for ensemble, ST $\to$ ST(PEFT) $\to$ ST(PEFT+VFMs), boosts performance, highlighting the benefits of pseudo-label diversity.
-6.1 Performance Comparison
-Following Section 5.4, we use ST(PEFT+VFMs) to refer to ensembling over multiple PEFT methods and VFMs; ST(PEFT) to refer to ensembling over multiple PEFT methods but one VFM; and ST to refer to no ensembling.
-In all cases, when we fine-tune VFMs using pseudo-labels, we re-initialize the model from the original VFMs.
-Section 6 shows the performance comparison between ST, ST(PEFT), ST(PEFT+VFMs) and existing SSL methods across all the 12 settings in our benchmark.
-Overall, ST(PEFT+VFMs) achieve superior performance compared with others, as shown in the Average column.
-To further analyze the results, Figure 5 visualize the ranking frequency, where each matrix entry $(i,j)$ indicates how often method $i$ is ranked $j$-th across 12 settings.
-We then compute the mean rank (number in the bracket) of each method and sort them accordingly.
-Although ST(PEFT+VFMs) does not claim first place in every single setting, it secures the top rank most frequently, showcasing the effectiveness of pseudo-label ensemble and establishing it as a simple yet powerful baseline in the VFM era.
-Digging deeper into the three ST-based methods, we observe a consistent performance boost as more diverse pseudo-labels are introduced for the ensemble, i.e., ST $\to$ ST(PEFT) $\to$ ST(PEFT+VFMs), as shown in Figure 4.
+Digging deeper into the three ST-based methods, we observe a consistent performance boost as more diverse pseudo-labels are introduced for the ensemble, i.e., ST $\to$ PET $\to$ V-PET, as shown in Figure 1.
 This underscores the importance of diversity among pseudo-labels.
-An open question emerging from these results is whether the benefits will continue to scale as even more PEFT methods and VFMs are integrated.
-Figure 5: Ranking frequency of different SSL methods in the proposed benchmark.
-The number in $(i,j)$ indicates the frequency of method $i$ is ranked $j$-th across 12 settings.
-The number in the bracket indicates the average rank.
 6.2 Scalability of Ensembling
@@
 In particular, we expand our evaluation across two dimensions—VFMs and PEFT methods—to assess the broader applicability of our ensembling strategy.
-VFMs Beyond ViT-B (CLIP, DINOv2), we incorporate four more VFMs, for a total of six, including four ViT-B (CLIP, DINOv2, OpenCLIP, ImangeNet21k) (Cherti et al., 2023; Ridnik et al., 2021) and two ViT-L (CLIP, DINOv2).
-PEFT Following the same setup, we start with LoRA and AdaptFormer, then expand our selection with four additional PEFT: ConvPass (Jie & Deng, 2022), BitFit (Zaken et al., 2022), VPT-Deep (Jia et al., 2022b), and Fact-TT (Jie & Deng, 2023).
+Figure 5: Scaling analysis for ensembling more PEFT methods and VFMs. Starting from LoRA fine-tuning, performance improves with diverse pseudo-labels, but gain diminishes with more ensemble sources.
+VFMs Beyond ViT-B (CLIP, DINOv2), we incorporate four more VFMs, for a total of six, including four ViT-B (CLIP, DINOv2, OpenCLIP, ImangeNet21k) cherti2023reproducible ; ridnik2021imagenet21kpretrainingmasses and two ViT-L (CLIP, DINOv2).
+PEFT Following the same setup, we start with LoRA and AdaptFormer, then expand our selection with four additional PEFT: ConvPass jie2022convolutionalbypassesbettervision , BitFit zaken2022bitfitsimpleparameterefficientfinetuning , VPT-Deep jia2022visualprompttuning , and Fact-TT jie2023factfactortuninglightweightadaptation .
 These methods span a broad range of PEFT approaches, including selective-based, adapter-based, and prompt-based techniques.
-We evaluate CLIP on the DTD N3 dataset, training with LoRA, and examine ensemble performance across different numbers of pseudo-labels, ranging from 1 to 6.
-To gain deeper insights into the pattern, we examine two ensemble sources: (1) PEFT: Ensembling pseudo-labels generated by models trained with different PEFT methods while using the same pretrained VFM (CLIP) and (2) VFMs: Ensembling pseudo-labels generated by different VFMs, all trained using the same PEFT method (LoRA).
-For each experiment, we report the average over 5 runs, each with a different randomly sampled set of pseudo-labels. As shown in Figure 6, we summarize the results as follows.
-Both PEFT and VFMs Help.
-Whether we ensemble the pseudo-labels generated across different PEFT or VFMs, the performance is consistently improved.
-This indicates that both PEFT and VFMs contribute to the diversity of pseudo-labels and help to improve the ensemble robustness.
-Diminishing Return.
-The performance improvement diminishes as we increase the number of ensemble sources.
-This suggests that the performance gain from ensemble is not linearly proportional to the number of sources.
-In order to achieve a balance between performance and computational cost, we recommend using a moderate number of ensemble sources, around 2 to 4 in this case.
-6.3 Mean Label Consistently Outperforms
-Furthermore, as also shown in Figure 6, we investigate three ensemble strategies: (1) Mean Label (our proposed method); (2) Mean Logits; and (3) Mean Probabilities.
-Based on our observations from the previous section, we find that the Mean Label strategy is the most effective among the three.
+We train CLIP with LoRA on DTD N3, generate pseudo-labels, and ensemble groups of 1–6 labels.
+We compare (1) PEFT: ensemble of pseudo-labels from models using different PEFT methods on the same CLIP backbone, and (2) VFMs: ensemble of pseudo-labels from different VFMs all fine-tuned with LoRA.
+Results are averaged over five runs with distinct random pseudo-label sets (see Figure 5).
+Mean Label Consistently Outperforms.
+As also shown in Figure 5, we investigate three ensemble strategies: (1) Mean Label (our proposed); (2) Mean Logits; and (3) Mean Probabilities.
+We find that the Mean Label strategy is the most effective among the three.
 It consistently outperforms the other two strategies across different ensemble sources and ensemble sizes.
 This suggests that the Mean Label approach better captures the diversity of pseudo-labels and maximizes the information utilization from the ensemble sources.
-6.4 Effectiveness of Hyperparameter Tuning
-Table 4: Hyperparameter tuning comparison: our method versus baselines (random or chosen by one criterion).
-The absolute errors from oracle test accuracy (smaller the better) are reported averaged across 12 datasets, 5 SSL baselines, and 2 VFMs trained on 2 PEFT.
-The reported values are in the format: mean $\pm$ standard deviation.
-RankMe AMI ARI V-Measure FMI CHI BNM Random Ours
-7.9%$\pm$8.62.8%$\pm$5.02.8%$\pm$4.92.8%$\pm$4.92.9%$\pm$4.93.7%$\pm$7.012.0%$\pm$12.18.9%$\pm$8.12.6%$\pm$4.3%
-We begin by investigating whether the proposed hyperparameter tuning strategy accurately gauges SSL methods’ performance.
-To illustrate its advantages, we compare it with two baselines, both measured via the absolute difference from the oracle test performance.
-Specifically, the oracle test performance is defined as the highest accuracy across the entire hyperparameter search space; each method selects a hyperparameter configuration, and we record how far its resulting accuracy is from this optimum.
-By taking the absolute value of this difference, we quantify how closely each tuning method approximates the best possible result.
-Our baselines are: (1) Random expectation, which represents the average performance obtained by randomly choosing hyperparameters from the search space, and (2) Single criterion, which selects the best hyperparameter configuration using only one metric.
-Comparing with (1) offers insight into how well our method leverages the search space relative to a purely random choice, while comparing with (2) illustrates the advantage of integrating multiple evaluation criteria.
-As shown in Table 4, we measure the average absolute error between each method’s chosen accuracy and the oracle’s best accuracy (i.e., the maximum accuracy achievable in the search space) across 12 datasets, 5 SSL baselines, 2 VFMs (CLIP and DINOv2), and 2 PEFT methods (LoRA and AdaptFormer).
-Our method outperforms both baselines in most cases, suggesting that combining multiple metrics produces a more reliable and robust assessment of SSL methods.
-We hope future research adopts this approach to strengthen SSL evaluations.
-Figure 6: Scaling analysis for ensembling with more PEFT methods and VFMs.
-The starting point is pseudo-labels generated by the model itself fine-tuned with LoRA.
-The performance improves as more diverse pseudo-labels are incorporated, but the gain diminishes as the number of ensemble sources increases.
-6.5 Impact of Labeled-Only Initialization on SSL
-Per discussion in Section 4, a natural question is whether the performance of SSL methods can be improved by initializing from a model trained on labeled data using PEFT.
-To explore this, we conduct experiments on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10).
-We first train DINOv2 models on labeled data using LoRA and then fine-tune them with SoftMatch.
-These datasets were chosen because their labeled-only performance exceeds that of SoftMatch by a large margin, as shown in Section 6.
-The results, summarized in Table 5, suggest that even when starting from the labeled-only trained model, SSL performance still falls short.
-In some cases, it performs even worse than training from scratch.
-Table 5: Performance comparison of starting from a model trained on labeled data and then fine-tuning using SoftMatch.
-The results are reported on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10).
-For consistency, we fix LoRA with dimension 4 for all the training.
-Dataset Labeled-Only SoftMatch Labeled-Only $\rightarrow$ SoftMatch
-DTD N6 68.4% 64.0% 63.5%
-Diabetic Retinopathy N80 44.9% 44.5% 43.9%
-KITTI N10 54.0% 27.9% 50.1%
+Both PEFT and VFMs Help.
+Whether we ensemble the pseudo-labels generated across different PEFT or VFMs, the performance is consistently improved, indicating that both PEFT and VFMs contribute to the diversity of pseudo-labels and improve the ensemble robustness.
+Diminishing Return.
+The performance improvement diminishes as we increase the number of ensemble sources, suggesting the performance gain from ensemble is not linearly scalable.
+In order to achieve a balance between performance and computational cost, we recommend using around 2 to 4 in this case.
+Impact of Pseudo-Label Quality.
+When applying V-PET, one might wonder: if the performances of different VFMs vary significantly, can their pseudo-labels still be ensembled to boost downstream results?
+The answer is yes, we demonstrate the result in section B.6.
+Figure 6: Ranking frequency across SSL methods by the proposed benchmark.
+The number in $(i,j)$ indicates the frequency of method $i$ is ranked $j$-th across 12 settings.
+The number in bracket indicates the average rank.
 7 Conclusion
@@
 Specifically, given that vision foundation models (VFM) are pre-trained on millions of images, it remains unclear whether they would still benefit from abundant unlabeled data in conjunction with limited labeled data in a semi-supervised learning (SSL) setting.
-To address this gap, we conduct a comprehensive and systematic empirical study.
-Our findings reveal that VFMs can still benefit from SSL but require a specifically tailored approach.
-In particular, we propose a simple yet highly effective self-training algorithm, enhanced by pseudo-labels generated through the integration of multiple VFMs and parameter-efficient fine-tuning (PEFT) techniques.
+To this end, we conduct a comprehensive and systematic empirical study, revealing that VFMs can still benefit from SSL but require a specifically tailored approach.
+In particular, we propose a simple yet highly effective self-training algorithm that ensembles pseudo-labels from multiple VFMs and parameter-efficient fine-tuning (PEFT) techniques.
 Overall, our study offers actionable insights into SSL with VFMs and lays the groundwork for more scalable and practical semi-supervised learning in the era of foundation models.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
-Specifically, compared to prior work in semi-supervised learning and vision foundation models, we do not see additional negative ethical and societal impacts that would be created by our work.
 Appendix A Dataset Details
 We select 6 datasets with a diverse ranges of applications for our experiments, including:
-  • DTD (Cimpoi et al., 2014): The Describable Texture Dataset (DTD) is designed for texture pattern recognition.
+  • DTD [16]: The Describable Texture Dataset (DTD) is designed for texture pattern recognition.
 It includes a diverse collection of images from 47 distinct texture categories, offering rich and varied examples to evaluate texture feature extraction and classification methods under natural conditions.
-  • SUN397 (Xiao et al., 2010): SUN397 is a large-scale scene recognition benchmark that covers 397 categories of both indoor and outdoor scenes.
+  • SUN397 [64]: SUN397 is a large-scale scene recognition benchmark that covers 397 categories of both indoor and outdoor scenes.
 It provides a comprehensive testbed for scene classification algorithms.
-  • Resisc45 (Cheng et al., 2017): Resisc45 is a remote sensing image dataset specifically created for scene classification tasks.
+  • Resisc45 [14]: Resisc45 is a remote sensing image dataset specifically created for scene classification tasks.
 It comprises thousands of images across 45 categories, representing diverse landscapes such as urban, rural, and industrial areas.
 This dataset is instrumental in developing and benchmarking algorithms for geographic information systems and environmental monitoring.
-  • Diabetic-Retinopathy (Dugas et al., 2015): This dataset focuses on high-resolution retinal images used for detecting diabetic retinopathy.
+  • Diabetic-Retinopathy [22]: This dataset focuses on high-resolution retinal images used for detecting diabetic retinopathy.
 It includes images that span various stages of the disease, providing a robust resource for training and evaluating deep learning models aimed at early diagnosis and automated medical analysis in ophthalmology.
-  • Clevr-Count (Johnson et al., 2016): Derived from the CLEVR family of datasets, Clevr-Count is centered on visual reasoning, particularly object counting tasks.
+  • Clevr-Count [38]: Derived from the CLEVR family of datasets, Clevr-Count is centered on visual reasoning, particularly object counting tasks.
 Using synthetically generated images, it challenges models to accurately count objects within complex scenes, thereby assessing their ability to understand spatial relationships and compositional structures in visual data.
-  • KITTI-Dist (Geiger et al., 2012): KITTI-Dist is based on the renowned KITTI benchmark and is tailored for depth estimation in vehicular contexts.
+  • KITTI-Dist [27]: KITTI-Dist is based on the renowned KITTI benchmark and is tailored for depth estimation in vehicular contexts.
 It provides high-quality stereo images along with ground-truth depth maps, supporting research in autonomous driving, 3D reconstruction, and stereoscopic vision, and serves as a key resource for evaluating depth estimation algorithms.
@@
 Dataset $N^{v}$ Test Size
-DTD 752 1880
-SUN397 17401 21750
-Resisc45 5040 6300
-Retinopathy 9207 42670
-Clevr-C 14000 15000
-KITTI 1354 711
+DTD 752 1,880
+SUN397 17,401 21,750
+Resisc45 5,040 6,300
+Retinopathy 9,207 42670
+Clevr-C 14,000 15,000
+KITTI 1,354 711
 Table 4: Dataset statistics.
@@
 Additionally, we utilized NVIDIA Tesla V100, NVIDIA Tesla A100, and NVIDIA RTX H100 GPUs for certain experiments.
-All experiments were implemented using PyTorch (Paszke et al., 2019).
+All experiments were implemented using PyTorch [51].
 B.2 SSL and Labeled-Only Baseline Details
-For all the experiments, we employ AdamW (Loshchilov & Hutter, 2019) optimizer with a cosine annealing learning rate scheduler (Loshchilov & Hutter, 2017) with warm-up period with $2.5\%$ of total iterations.
+For all the experiments, we employ AdamW [43] optimizer with a cosine annealing learning rate scheduler [42] with warm-up period with $2.5\%$ of total iterations.
 We use a batch size of $32+32$ for all experiments, where the first $32$ corresponds to labeled data and the second $32$ corresponds to unlabeled data.
-We list the hyperparameter search spaces for the SSL, Labeled-Only, and ST settings in Table 7, including the drop path rate (dpr) (Huang et al., 2016), training augmentation (train-aug), LoRA dimension (lora_dim), adapter bottleneck size (adapter_bottleneck), weight decay, momentum, learning rate (lr), and number of epochs.
+We list the hyperparameter search spaces for the SSL, Labeled-Only, and ST settings in Table 5, including the drop path rate (dpr) [33], training augmentation (train-aug), LoRA dimension (lora_dim), adapter bottleneck size (adapter_bottleneck), weight decay, momentum, learning rate (lr), and number of epochs.
 For other unmentioned SSL algorithm related hyperparameters, we use the default values provided in the original papers.
@@
 Table 5: Hyperparameter search spaces for SSL, Labeled-Only, and ST settings.
-As a completion of Section 6, we present the detailed results of the SSL baseline experiment with different SSL methods, including full fine-tuning, in Section B.2.
+As a completion of subsection 6.1, we present the detailed results of the SSL baseline experiment with different SSL methods, including full fine-tuning, in subsection B.2.
 DTD Sun397 Resisc45 Retinopathy Clevr-C KITTI AVERAGE
@@
 […]
-ST (PEFT) 63.19% 68.67% 63.28% 68.83% 68.32% 79.73% 39.46% 43.35% 37.33% 49.87% 60.34% 63.85% 58.85%
-ST (PEFT + VFMs) 66.33% 70.64% 66.42% 72.23% 67.49% 79.11% 39.46% 55.01% 35.61% 54.88% 57.24% 65.12% 60.80%
+PET 63.19% 68.67% 63.28% 68.83% 68.32% 79.73% 39.46% 43.35% 37.33% 49.87% 60.34% 63.85% 58.85%
+V-PET 66.33% 70.64% 66.42% 72.23% 67.49% 79.11% 39.46% 55.01% 35.61% 54.88% 57.24% 65.12% 60.80%
 LoRA Labeled-Only 57.87% 64.10% 60.66% 66.79% 59.48% 72.05% 35.38% 42.16% 35.21% 50.80% 60.06% 64.28% 55.74%
@@
 […]
-ST (PEFT) 63.35% 68.24% 65.61% 70.80% 67.52% 78.56% 36.42% 42.39% 38.27% 54.63% 60.34% 65.68% 59.32%
-ST (PEFT + VFMs) 65.59% 71.65% 67.19% 72.83% 66.56% 77.21% 38.85% 51.06% 36.69% 58.49% 57.95% 62.31% 60.53%
+PET 63.35% 68.24% 65.61% 70.80% 67.52% 78.56% 36.42% 42.39% 38.27% 54.63% 60.34% 65.68% 59.32%
+V-PET 65.59% 71.65% 67.19% 72.83% 66.56% 77.21% 38.85% 51.06% 36.69% 58.49% 57.95% 62.31% 60.53%
 AdaptFormer Labeled-Only 61.54% 64.95% 61.14% 68.23% 60.98% 72.83% 34.38% 38.99% 37.80% 48.63% 57.81% 59.07% 55.53%
@@
 […]
-ST (PEFT) 63.78% 68.94% 66.78% 71.71% 69.54% 80.16% 37.14% 41.51% 38.45% 55.78% 60.48% 61.74% 59.67%
-ST (PEFT + VFMs) 65.69% 71.76% 67.81% 73.23% 67.92% 78.43% 38.90% 51.22% 36.96% 59.01% 58.23% 63.29% 61.04%
+PET 63.78% 68.94% 66.78% 71.71% 69.54% 80.16% 37.14% 41.51% 38.45% 55.78% 60.48% 61.74% 59.67%
+V-PET 65.69% 71.76% 67.81% 73.23% 67.92% 78.43% 38.90% 51.22% 36.96% 59.01% 58.23% 63.29% 61.04%
 DINOv2
@@
 […]
-ST (PEFT) 64.41% 69.84% 59.94% 66.87% 55.76% 69.62% 34.34% 56.55% 31.40% 55.02% 46.98% 64.98% 56.31%
-ST (PEFT + VFMs) 66.97% 73.78% 65.53% 71.48% 67.32% 79.25% 36.55% 52.98% 35.93% 56.55% 55.98% 64.70% 60.59%
+PET 64.41% 69.84% 59.94% 66.87% 55.76% 69.62% 34.34% 56.55% 31.40% 55.02% 46.98% 64.98% 56.31%
+V-PET 66.97% 73.78% 65.53% 71.48% 67.32% 79.25% 36.55% 52.98% 35.93% 56.55% 55.98% 64.70% 60.59%
 LoRA Labeled-Only 61.01% 68.40% 59.71% 66.78% 48.59% 63.08% 40.69% 48.02% 35.87% 61.72% 46.69% 71.73% 56.02%
@@
 […]
-ST (PEFT) 66.65% 72.18% 58.57% 65.29% 55.79% 70.59% 40.99% 54.28% 33.11% 56.85% 42.62% 63.29% 56.68%
-ST (PEFT + VFMs) 67.77% 74.10% 67.12% 73.24% 66.03% 77.03% 39.20% 52.48% 36.77% 59.25% 58.23% 63.71% 61.24%
+PET 66.65% 72.18% 58.57% 65.29% 55.79% 70.59% 40.99% 54.28% 33.11% 56.85% 42.62% 63.29% 56.68%
+V-PET 67.77% 74.10% 67.12% 73.24% 66.03% 77.03% 39.20% 52.48% 36.77% 59.25% 58.23% 63.71% 61.24%
 AdaptFormer Labeled-Only 62.98% 67.61% 58.87% 67.03% 47.00% 57.41% 34.42% 53.07% 27.33% 41.79% 52.32% 49.79% 51.64%
@@
 […]
-ST (PEFT) 67.18% 72.98% 63.47% 70.93% 56.21% 72.00% 40.18% 55.53% 33.05% 57.95% 45.15% 60.90% 57.96%
-ST (PEFT + VFMs) 68.24% 74.15% 67.65% 73.43% 66.32% 77.98% 39.39% 52.53% 36.77% 59.97% 59.21% 61.74% 61.45%
+PET 67.18% 72.98% 63.47% 70.93% 56.21% 72.00% 40.18% 55.53% 33.05% 57.95% 45.15% 60.90% 57.96%
+V-PET 68.24% 74.15% 67.65% 73.43% 66.32% 77.98% 39.39% 52.53% 36.77% 59.97% 59.21% 61.74% 61.45%
+5.‘
 Table 6: Complete results of the SSL baseline experiment with different SSL methods.
@@
 This demonstrates the effectiveness of PEFT in enhancing the performance of VFMs within SSL frameworks.
+B.4 Linear Probing with Different Classification Heads
+We provide results of frozen backbone with different classification heads on our benchmark in Table 8.
+DTD SUN397 Resisc45 Retinopathy Clevr-C KITTI
+N $6$ $6$ $2$ $80$ $20$ $10$
+CLIP-LINEAR 61.80% 63.70% 69.30% 35.90% 33.10% 51.10%
+CLIP-MLP-96 63.09% 60.29% 67.89% 36.91% 38.11% 50.77%
+CLIP-MLP-768 62.29% 62.52% 69.62% 39.67% 39.51% 48.10%
+CLIP LoRA 64.10% 66.79% 72.05% 42.16% 50.80% 64.28%
+DINOv2-LINEAR 66.70% 65.30% 52.40% 41.20% 30.50% 51.30%
+DINOv2-MLP-96 66.97% 62.41% 50.30% 44.33% 36.37% 50.91%
+DINOv2-MLP-768 67.98% 61.16% 52.06% 35.66% 35.21% 50.77%
+DINOv2 LoRA 68.40% 66.78% 63.08% 48.02% 61.72% 71.73%
+Table 8: Linear probing results compared to LoRA under various scenarios, where LINEAR denote a single linear layer and MLP-x denotes an MLP with a single hidden layer of size x.
+The results demonstrate that LoRA is a more effective approach for fine-tuning vision foundation models.
+B.5 Impact of Labeled-Only Initialization on SSL
+Table 9: Performance of starting from a labeled-only model then fine-tuning with SoftMatch. Results on DTD (N6), Retinopathy (N80), and KITTI (N10). LoRA dim=4 is fixed.
+Dataset Labeled-Only SoftMatch Labeled-Only $\rightarrow$ SoftMatch
+DTD N6 68.4% 64.0% 63.5%
+Retinopathy N80 44.9% 44.5% 43.9%
+KITTI N10 54.0% 27.9% 50.1%
+Per discussion in section 4, a natural question is whether the performance of SSL methods can be improved by initializing from a model trained on labeled data using PEFT.
+To explore this, we conduct experiments on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10).
+We first train DINOv2 models on labeled data using LoRA and then fine-tune them with SoftMatch.
+These datasets were chosen because their labeled-only performance exceeds that of SoftMatch by a large margin, as shown in subsection 6.1.
+The results, summarized in Table 9, suggest that even when starting from the labeled-only trained model, SSL performance still falls short.
+In some cases, it performs even worse than training from scratch.
+B.6 The Effectiveness of VFMs Ensemble
+Table 10: We use pseudo-labels from different source to train a CLIP-B model. Demonstrating when performing ensembling training, even weaker pseudo-labels will boost performance.
+Model Test Accuracy
+CLIP 56.17%
+ViT-B-IN21k 25.27%
+CLIP + ViT-B-IN21k 58.62%
+To assess the benefit of pseudo-label ensembling across VFMs, even when their individual accuracies differ significantly, we performed an ablation study on the CLIP-B model.
+We trained three variants on different pseudo-label sets: (1) labels generated by CLIP itself, (2) labels from ViT-B-IN21k, a comparatively weaker VFM, and (3) the ensemble of both sources.
+Results in Table 10 show that, despite ViT-B-IN21k’s lower standalone accuracy, combining its labels with CLIP’s improves overall performance.
+B.7 The Effectiveness of Hyperparameter Tuning
+Table 11: Hyperparameter tuning comparison: our method versus baselines (random or chosen by one criterion).
+The absolute errors from oracle test accuracy (smaller the better) are reported averaged across 12 datasets, 5 SSL baselines, and 2 VFMs trained on 2 PEFT.
+The reported values are in the format: mean $\pm$ standard deviation.
+RankMe AMI ARI V-Measure FMI CHI BNM Random Ours
+7.9%$\pm$8.62.8%$\pm$5.02.8%$\pm$4.92.8%$\pm$4.92.9%$\pm$4.93.7%$\pm$7.012.0%$\pm$12.18.9%$\pm$8.12.6%$\pm$4.3%
+We begin by investigating whether the proposed hyperparameter tuning strategy accurately gauges SSL methods’ performance.
+To illustrate its advantages, we compare it with two baselines, both measured via the absolute difference from the oracle test performance.
+Specifically, the oracle test performance is defined as the highest accuracy across the entire hyperparameter search space; each method selects a hyperparameter configuration, and we record how far its resulting accuracy is from this optimum.
+By taking the absolute value of this difference, we quantify how closely each tuning method approximates the best possible result.
+Our baselines are: (1) Random expectation, which represents the average performance obtained by randomly choosing hyperparameters from the search space, and (2) Single criterion, which selects the best hyperparameter configuration using only one metric.
+Comparing with (1) offers insight into how well our method leverages the search space relative to a purely random choice, while comparing with (2) illustrates the advantage of integrating multiple evaluation criteria.
+As shown in Table 11, we measure the average absolute error between each method’s chosen accuracy and the oracle’s best accuracy (i.e., the maximum accuracy achievable in the search space) across 12 datasets, 5 SSL baselines, 2 VFMs (CLIP and DINOv2), and 2 PEFT methods (LoRA and AdaptFormer).
+Our method outperforms both baselines in most cases, suggesting that combining multiple metrics produces a more reliable and robust assessment of SSL methods.
+We hope future research adopts this approach to strengthen SSL evaluations.
 Appendix C Hyperparameter Tuning Details
@@
 Tuning without Labels The primary challenge in tuning hyperparameters within an unsupervised setting lies in effectively adjusting them without access to labels.
-Inspired by the approach for standardizing hyperparameter tuning in Unsupervised Domain Adaptation (UDA) (Ericsson et al., 2023), we propose tuning the hyperparameters using $\phi_{\mathcal{D}_{v}}$, $\ell_{\mathcal{D}_{v}}$ and $\pi_{\mathcal{D}_{v}}$.
+Inspired by the approach for standardizing hyperparameter tuning in Unsupervised Domain Adaptation (UDA) [23], we propose tuning the hyperparameters using $\phi_{\mathcal{D}_{v}}$, $\ell_{\mathcal{D}_{v}}$, and $\pi_{\mathcal{D}_{v}}$.
 Specifically, we define function $v$ as a validator:
@@
 $A$ $\coloneqq\begin{bmatrix}\frac{1}{n}\sum_{j=1}^{n}r_{1j}\\ \frac{1}{n}\sum_{j=1}^{n}r_{2j}\\ \vdots\\ \frac{1}{n}\sum_{j=1}^{n}r_{hj}\\ \end{bmatrix}.$
-To get the best performance model, we select the model with lowest value in $A$. We put the validators we use and more detailed examples in Section C.2.
+To get the best performance model, we select the model with the lowest value in $A$. We put the validators we use and more detailed examples in section C.2.
 C.2 Validators
-We deploys 7 semi-supervised learning validators, including
-RankMe Score (Garrido et al., 2023) We adapt RankMe Score to compute the feature matrix rank of the pre-trained data over both source and target data domains.
-Adjusted Mutual Information (AMI) (Vinh et al., 2009) We adapt AMI to compute the adjusted mutual information between predicted and cluster labels of the validation set.
-Adjusted Rand Index (ARI) (Halkidi et al., 2002) ARI is a widely used metric in cluster analysis by measuring the similarity between two clustering solutions.
+We deploy 7 semi-supervised learning validators, including
+RankMe Score [26] We adapt RankMe Score to compute the feature matrix rank of the pre-trained data over both source and target data domains.
+Adjusted Mutual Information (AMI) [63] We adapt AMI to compute the adjusted mutual information between predicted and cluster labels of the validation set.
+Adjusted Rand Index (ARI) [29] ARI is a widely used metric in cluster analysis by measuring the similarity between two clustering solutions.
 Given $RI$ is the Rank Index between the true and predicted clustering, ARI is defined by:
 $ARI\coloneqq\frac{RI-\text{Expected RI}}{\max(RI)-\text{Expected RI}}$
-V-Measure (Rosenberg & Hirschberg, 2007) Beside AMI, we also adapt V-Measure to measure the harmonic mean between homogeneity and completeness over the clustering labels and prediction.
-Fowlkes-Mallows Index (FMI) (Fowlkes & Mallows, 1983) Measures the similarity between two clusterings, FMI is defined by:
+V-Measure [55] Beside AMI, we also adapt V-Measure to measure the harmonic mean between homogeneity and completeness over the clustering labels and prediction.
+Fowlkes-Mallows Index (FMI) [24] Measures the similarity between two clusterings. FMI is defined by:
 $FMI\coloneqq\sqrt{\frac{TP}{TP+FP}\cdot\frac{TP}{TP+FN}}$
-Calinski-Harabasz Index (CHI) (Caliński & Harabasz, 1974) CHI is an internal clustering measurement.
+Calinski-Harabasz Index (CHI) [9] CHI is an internal clustering measurement.
 Given $n$ sets of data point $\{x\}_{1}^{n}$ grouped into $k$ clusters $\{C\}_{1}^{k}$ with centroids $\{c\}_{i}^{k}$ and $\bm{c}$ is the overall centroid, CHI is defined by:
 $CHI\coloneqq\frac{\left[\sum_{i=1}^{k}n_{i}||c_{i}-c||^{2}\right]/(k-1)}{\left[\sum_{i=1}^{k}\sum_{x\in C_{i}}||x-c_{i}||^{2}\right](n-k)}$
-Batch Nuclear-norm Minimization (BNM) (Cui et al., 2020) An unsupervised domain adaptation method aims to generate predictions that favor both diversion and confidence; BNM maximizes the nuclear norm of the prediction matrix in a batch and repurposes it as a validation criterion.
+Batch Nuclear-norm Minimization (BNM) [18] An unsupervised domain adaptation method aims to generate predictions that favor both diversion and confidence; BNM maximizes the nuclear norm of the prediction matrix in a batch and repurposes it as a validation criterion.
 C.3 Hyperparameter Tuning Results
+We present complete hyperparameter tuning results in Table 12.
 DTD Sun397 Resisc45 Retinopathy Clevr-C KITTI

### v3

@@
 Results are averaged across 12 settings in our benchmark.
-Recently, pre-training on external labeled or unlabeled data has become the de facto standard across many machine learning application domains gui2024survey .
+Recently, pre-training on external labeled or unlabeled data has become the de facto standard across many machine learning application domains gui2024survey ; mai2024fine .
 For example, in computer vision, many recent algorithms are built upon vision foundation models (VFMs) such as CLIP radford2021learningtransferablevisualmodels and DINOv2 oquab2024dinov2learningrobustvisual .
@@
 Transfer Learning & Self-Supervised Learning.
-Transfer learning zhuang2020comprehensivesurveytransferlearning —and specifically PEFT (or PETL) for efficiently adapting foundation models han2024parameterefficientfinetuninglargemodels —has long leveraged pretrained models to boost downstream tasks.
+Transfer learning zhuang2020comprehensivesurveytransferlearning ; tu2023holistic —and specifically PEFT (or PETL) for efficiently adapting foundation models han2024parameterefficientfinetuninglargemodels —has long leveraged pretrained models to boost downstream tasks.
 In our approach, we fine-tune VFMs via PEFT on labeled data, then use the adapted models to generate pseudo-labels for unlabeled examples, thus bridging transfer learning with semi-supervised methods.
@@
 Vision Foundation Models.
-Vision Transformers (ViT) dosovitskiy2021imageworth16x16words pre-trained on massive amounts of data, has become indispensable to modern AI development awais2025foundation .
+Vision Transformers (ViT) dosovitskiy2021imageworth16x16words pre-trained on massive amounts of data, has become indispensable to modern AI development awais2025foundation ; mai2025ava ; gu2025bioclip .
 These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks.
@@
 CIFAR-10 CIFAR-100 FOOD-101 DTD SUN397 Resisc45 Retinopathy Clevr-C KITTI
-N $4$ $100$ $10$ $6$ $6$ $2$ $80$ $20$ $10$
+N 44 $100$ $10$ 66 66 22 $80$ 2020 $10$
 CLIP 85.0% 78.3% 80.2% 61.8% 63.7% 69.3% 35.9% 33.1% 51.1%
@@
 However, these benchmarks exhibit two key limitations: Diminishing difficulty under VFMs.
-Following prior PEFT-on-VFM work jie2022convolutionalbypassesbettervision ; jia2022visual ; tu2023visual ; mai2024lessonslearnedunifyingempirical , we use linear probing as our primary performance indicator.
+Following prior PEFT-on-VFM work jie2022convolutionalbypassesbettervision ; jia2022visual ; tu2023visual ; mai2024lessonslearnedunifyingempirical ; xiao2025visual ; zhang2025ot , we use linear probing as our primary performance indicator.
 As shown in Table 1, linear probing on frozen VFM backbones already delivers remarkable accuracy—even with limited labeled data.
-We also report results for various classification-head sizes in section B.4, and Narrow domain coverage.
+We also report results for various classification-head sizes in section˜B.4, and Narrow domain coverage.
 Because these benchmarks focus mainly on natural images, they offer only a narrow view of real-world VFM applications.
-To effectively evaluate SSL methods in the VFM era—focusing exclusively on semi-supervised image classification—we propose a new benchmark designed to capture the complexities of real-world applications across different domains and dataset sizes.
+To effectively evaluate SSL methods in the VFM era—focusing exclusively on semi‐supervised image classification—we propose a new benchmark designed to capture the complexities of real‐world applications across different domains and dataset sizes.
 Dataset & Regime We choose six classification datasets from VTAB zhai2020largescalestudyrepresentationlearning , spanning remote sensing, medical imaging, synthetic data, and autonomous driving tasks.
@@
 A summary of the datasets appears in Table 2, with full details provided in Appendix A.
-Together, this benchmark offers a more diverse and comprehensive foundation-model-era evaluation of semi-supervised image classification.
+Together, this benchmark offers a more diverse and comprehensive foundation‐model‐era evaluation of semi‐supervised image classification.
 3.3 Fair Hyperparameter Tuning for SSL
@@
 Table 2: Summary of our benchmark. Rem. Sen.: Remote sensing; Recog.: Recognition; $|\mathcal{L}|$: # labeled training data; $|\mathcal{U}|$: # unlabeled training data.
-Dataset Task Domain $|\mathcal{L}|+|\mathcal{U}|$ Classes $|\mathcal{L}|/$Class
+[…]
 DTD Recog. Textual 3,008 47 3, 6
@@
 Next, we rank every hyperparameter configuration based on each criterion and choose the one achieving the lowest average rank across all.
-We provide formal definitions and detailed procedures in section C.1 and the effectiveness of the proposed method in section B.7.
+We provide formal definitions and detailed procedures in section˜C.1 and the effectiveness of the proposed method in section˜B.7.
 By eliminating reliance on held-out labeled validation sets, our method mitigates data leakage and promotes a more practical and fair tuning protocol for SSL.
@@
 Given the struggling performance of frozen VFMs in our benchmark (Table 1), we consider two straightforward strategies to enhance their effectiveness: (1) exploit the inherent generalizability of VFMs by fine-tuning only on labeled data; (2) employ SSL to leverage both labeled and unlabeled data.
-With the diverse benchmark and hyperparameter tuning protocol introduced in section 3, we aim to investigate whether existing SSL algorithms remain effective when adopting VFMs as their backbone.
+With the diverse benchmark and hyperparameter tuning protocol introduced in section˜3, we aim to investigate whether existing SSL algorithms remain effective when adopting VFMs as their backbone.
 Evaluation Setup.
@@
 We, therefore, assess two commonly used PEFT strategies: LoRA hu2021loralowrankadaptationlarge , which trains additive low-rank matrices for Transformer layers to approximate weight updates, and AdaptFormer chen2022adaptformeradaptingvisiontransformers , an adapter-based approach proven effective in computer vision.
-We follow the same protocol outlined in section 3 to tune the hyperparameters of these PEFT methods.
+We follow the same protocol outlined in section˜3 to tune the hyperparameters of these PEFT methods.
 Our results in Figure 4 show that PEFT indeed improves SSL on VFMs.
@@
 The core idea of self-training is to gradually assign pseudo-labels to unlabeled data $\mathcal{U}=\{{\bm{x}}^{u}_{i}\}_{i=1}^{|\mathcal{U}|}$ based on high-confidence predictions
-[…]
-and add these pseudo-labeled data to the labeled set =$\mathcal{L}=\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{|\mathcal{L}|}$ for supervised learning.
+$\hat{y}^{u}_{i}\coloneqq\underset{c}{\operatorname{arg\,max}}\ p\left(c\big|{\bm{x}}^{u}_{i};f_{\bm{\theta}}\right)\hskip 20.0pt\text{ if }\ \ \underset{c}{\max}\ p(c\big|{\bm{x}}^{u}_{i};f_{\bm{\theta}})\geq\tau,$
+and add these pseudo-labeled data to the labeled set $\mathcal{L}=\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{|\mathcal{L}|}$ for supervised learning.
 Here, $\bm{\theta}$ represents the weights of the current classifier, and $\tau$ is the confidence threshold.
@@
   Compute one-hot pseudo-label set $\mathcal{P}_{n,m}$ as
-[…]
+$\mathcal{P}_{n,m}=\left\{\text{one\_hot}\bigl(\underset{c}{\operatorname{arg\,max}}f_{\tilde{\bm{\theta}}_{n,m}}(u)\bigr)\;\middle|\;\forall u\in\mathcal{U}\right\}.$
 // (c) Pseudo-Label Ensemble
@@
 We use ST (self-training without ensembling) as a baseline and, in all cases, reinitialize from the original pretrained VFM before fine-tuning with pseudo-labels.
-Figure 3: Average SSL accuracy with full fine-tuning or PEFT across 12 settings shows that, with fair hyper-parameter tuning, fine-tuning on limited labels can outperform SSL; PEFT boosts SSL yet matches labeled-only performance, indicating minimal unlabeled-data benefit in current VFM-based SSL (see section B.3).
+Figure 3: Average SSL accuracy with full fine-tuning or PEFT across 12 settings shows that, with fair hyper-parameter tuning, fine-tuning on limited labels can outperform SSL; PEFT boosts SSL yet matches labeled-only performance, indicating minimal unlabeled-data benefit in current VFM-based SSL (see section˜B.3).
 Figure 4: The average entropy of predicted probability distributions from different fine-tuned VFMs, highlighting the entropy gap among pseudo-labels, indicating poor calibration.
 6.1 Performance Comparison
-We summarize the performance comparison between ST, PET, V-PET and existing SSL methods across all the 12 settings in subsection 6.1.
+We summarize the performance comparison between ST, PET, V-PET and existing SSL methods across all the 12 settings in Table 3.
 Overall, V-PET achieve superior performance compared with others, as shown in the Average column.
@@
 FineSSL 63.1% 69.1% 56.3% 61.8% 74.2% 83.5% 35.0% 30.4% 33.6% 38.1% 53.7% 48.2% 53.9%
-[…]
+ST 58.9% 63.8% 49.2% 52.7% 64.1% 77.4% 35.7% 42.8% 36.5% 51.6% 59.8% 64.3% 54.7%
 PET 63.4% 68.2% 65.6% 70.8% 67.5% 78.6% 36.4% 42.4% 38.3% 54.6% 60.3% 65.7% 59.3%
@@
 FineSSL 61.9% 68.1% 58.4% 64.9% 73.9% 84.2% 25.3% 29.8% 26.0% 36.9% 42.2% 47.1% 51.6%
-[…]
+ST 62.0% 67.2% 58.1% 61.3% 68.8% 79.3% 34.4% 39.6% 39.0% 50.3% 58.2% 61.9% 56.7%
 PET 63.8% 68.9% 66.8% 71.7% 69.5% 80.2% 37.1% 41.5% 38.4% 55.8% 60.5% 61.7% 59.7%
@@
 FineSSL 66.9% 72.7% 54.6% 62.1% 69.8% 87.4% 31.0% 24.6% 24.8% 31.7% 31.9% 39.8% 49.8%
-[…]
+ST 52.3% 56.9% 45.4% 50.5% 54.7% 72.4% 42.6% 49.6% 36.3% 62.7% 33.2% 77.8% 52.9%
 PET 66.7% 72.2% 58.6% 65.3% 55.8% 70.6% 41.0% 54.3% 33.1% 56.9% 42.6% 63.3% 56.7%
@@
 FineSSL 64.0% 71.2% 56.0% 63.3% 62.3% 69.7% 17.7% 24.3% 25.8% 30.0% 31.8% 45.0% 46.8%
-[…]
+ST 62.3% 67.0% 59.0% 67.1% 54.1% 66.2% 34.9% 56.0% 27.4% 43.1% 43.9% 48.7% 52.5%
 PET 67.2% 73.0% 63.5% 70.9% 56.2% 72.0% 40.2% 55.5% 33.1% 58.0% 45.2% 60.9% 58.0%
@@
 The best result within each PEFT and VFM is highlighted in bold.
-We report the full results in section B.2.
+We report the full results in section˜B.2.
 Digging deeper into the three ST-based methods, we observe a consistent performance boost as more diverse pseudo-labels are introduced for the ensemble, i.e., ST $\to$ PET $\to$ V-PET, as shown in Figure 1.
@@
 When applying V-PET, one might wonder: if the performances of different VFMs vary significantly, can their pseudo-labels still be ensembled to boost downstream results?
-The answer is yes, we demonstrate the result in section B.6.
+The answer is yes, we demonstrate the result in section˜B.6.
 Figure 6: Ranking frequency across SSL methods by the proposed benchmark.
@@
 Overall, our study offers actionable insights into SSL with VFMs and lays the groundwork for more scalable and practical semi-supervised learning in the era of foundation models.
+Acknowledgment
+This research is supported by grants from the National Science Foundation (ICICLE: OAC-2112606). We appreciate the generous support of computational resources from the Ohio Supercomputer Center (OSC).
 Appendix A Dataset Details
@@
 Table 5: Hyperparameter search spaces for SSL, Labeled-Only, and ST settings.
-As a completion of subsection 6.1, we present the detailed results of the SSL baseline experiment with different SSL methods, including full fine-tuning, in subsection B.2.
+As a completion of Table 3, we present the detailed results of the SSL baseline experiment with different SSL methods, including full fine-tuning, in Table 6.
 DTD Sun397 Resisc45 Retinopathy Clevr-C KITTI AVERAGE
@@
 FineSSL 7.02% 17.18% 1.46% 4.01% 7.78% 8.37% 39.64% 16.97% 19.53% 26.08% 41.49% 44.59% 19.51%
-[…]
+ST 54.10% 61.60% 52.88% 59.30% 55.95% 69.90% 34.69% 32.95% 24.65% 29.78% 34.46% 65.40% 47.97%
 PET 63.19% 68.67% 63.28% 68.83% 68.32% 79.73% 39.46% 43.35% 37.33% 49.87% 60.34% 63.85% 58.85%
@@
 FineSSL 63.14% 69.10% 56.25% 61.78% 74.17% 83.48% 35.03% 30.42% 33.55% 38.07% 53.73% 48.24% 53.91%
-[…]
+ST 58.94% 63.83% 49.24% 52.68% 64.05% 77.43% 35.73% 42.77% 36.45% 51.63% 59.77% 64.28% 54.73%
 PET 63.35% 68.24% 65.61% 70.80% 67.52% 78.56% 36.42% 42.39% 38.27% 54.63% 60.34% 65.68% 59.32%
@@
 FineSSL 61.91% 68.14% 58.37% 64.86% 73.95% 84.17% 25.27% 29.78% 25.97% 36.87% 42.19% 47.12% 51.55%
-[…]
+ST 61.97% 67.23% 58.10% 61.29% 68.84% 79.33% 34.42% 39.60% 39.01% 50.33% 58.23% 61.88% 56.69%
 PET 63.78% 68.94% 66.78% 71.71% 69.54% 80.16% 37.14% 41.51% 38.45% 55.78% 60.48% 61.74% 59.67%
@@
 FineSSL 23.78% 48.40% 2.76% 16.57% 10.54% 22.05% 22.25% 21.17% 20.75% 23.52% 39.94% 35.44% 23.93%
-[…]
+ST 49.47% 60.43% 46.83% 56.59% 5.38% 61.79% 35.68% 43.38% 27.17% 26.81% 60.20% 60.90% 44.55%
 PET 64.41% 69.84% 59.94% 66.87% 55.76% 69.62% 34.34% 56.55% 31.40% 55.02% 46.98% 64.98% 56.31%
@@
 FineSSL 66.97% 72.66% 54.62% 62.07% 69.78% 87.38% 30.99% 24.55% 24.77% 31.73% 31.93% 39.80% 49.77%
-[…]
+ST 52.34% 56.86% 45.39% 50.47% 54.73% 72.44% 42.58% 49.56% 36.29% 62.74% 33.19% 77.78% 52.86%
 PET 66.65% 72.18% 58.57% 65.29% 55.79% 70.59% 40.99% 54.28% 33.11% 56.85% 42.62% 63.29% 56.68%
@@
 FineSSL 64.04% 71.17% 55.96% 63.34% 62.32% 69.73% 17.66% 24.30% 25.79% 30.03% 31.79% 45.01% 46.76%
-[…]
+ST 62.29% 67.02% 59.00% 67.11% 54.06% 66.17% 34.87% 55.96% 27.41% 43.13% 43.88% 48.66% 52.46%
 PET 67.18% 72.98% 63.47% 70.93% 56.21% 72.00% 40.18% 55.53% 33.05% 57.95% 45.15% 60.90% 57.96%
@@
 DTD SUN397 Resisc45 Retinopathy Clevr-C KITTI
-N $6$ $6$ $2$ $80$ $20$ $10$
+N 66 66 22 $80$ 2020 $10$
 CLIP-LINEAR 61.80% 63.70% 69.30% 35.90% 33.10% 51.10%
@@
 KITTI N10 54.0% 27.9% 50.1%
-Per discussion in section 4, a natural question is whether the performance of SSL methods can be improved by initializing from a model trained on labeled data using PEFT.
+Per discussion in section˜4, a natural question is whether the performance of SSL methods can be improved by initializing from a model trained on labeled data using PEFT.
 To explore this, we conduct experiments on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10).
 We first train DINOv2 models on labeled data using LoRA and then fine-tune them with SoftMatch.
-These datasets were chosen because their labeled-only performance exceeds that of SoftMatch by a large margin, as shown in subsection 6.1.
+These datasets were chosen because their labeled-only performance exceeds that of SoftMatch by a large margin, as shown in Table 3.
 The results, summarized in Table 9, suggest that even when starting from the labeled-only trained model, SSL performance still falls short.
@@
 RankMe AMI ARI V-Measure FMI CHI BNM Random Ours
-7.9%$\pm$8.62.8%$\pm$5.02.8%$\pm$4.92.8%$\pm$4.92.9%$\pm$4.93.7%$\pm$7.012.0%$\pm$12.18.9%$\pm$8.12.6%$\pm$4.3%
+[…]
 We begin by investigating whether the proposed hyperparameter tuning strategy accurately gauges SSL methods’ performance.
@@
 Problem Formulation Formally, the objective is to find a suitable function,
-[…]
-[…]
-[…]
+$f$ $\coloneqq g\circ h\in\mathbb{R}^{i}\rightarrow\mathbb{R}^{j}\quad$ $\text{(model)},$
+$h$ $\in\mathbb{R}^{i}\rightarrow\mathbb{R}^{k}\quad$ $\text{(feature extractor)},$
+$g$ $\in\mathbb{R}^{k}\rightarrow\mathbb{R}^{j}\quad$ $\text{(linear projector)}.$
 where $h$ serves as a feature extractor, and $g$ is a linear projector.
@@
 $v_{2}(\phi_{\theta_{2}},\ell_{\theta_{2}},\pi_{\theta_{2}}),$
-$\ldots$,… ,
+$\ldots,$
 $v_{n}(\phi_{\theta_{h}},\ell_{\theta_{h}},\pi_{\theta_{h}}).$
@@
 $A$ $\coloneqq\begin{bmatrix}\frac{1}{n}\sum_{j=1}^{n}r_{1j}\\ \frac{1}{n}\sum_{j=1}^{n}r_{2j}\\ \vdots\\ \frac{1}{n}\sum_{j=1}^{n}r_{hj}\\ \end{bmatrix}$.
-To get the best performance model, we select the model with the lowest value in $A$. We put the validators we use and more detailed examples in section C.2.
+To get the best performance model, we select the model with the lowest value in $A$. We put the validators we use and more detailed examples in section˜C.2.
 C.2 Validators
@@
 Given $n$ sets of data point $\{x\}_{1}^{n}$ grouped into $k$ clusters $\{C\}_{1}^{k}$ with centroids $\{c\}_{i}^{k}$ and $\bm{c}$ is the overall centroid, CHI is defined by:
-$CHI\coloneqq\frac{\left[\sum_{i=1}^{k}n_{i}||c_{i}-c||^{2}\right]/(k-1)}{\left[\sum_{i=1}^{k}\sum_{x\in C_{i}}||x-c_{i}||^{2}\right](n-k)}$
+[…]
 Batch Nuclear-norm Minimization (BNM) [18] An unsupervised domain adaptation method aims to generate predictions that favor both diversion and confidence; BNM maximizes the nuclear norm of the prediction matrix in a batch and repurposes it as a validation criterion.

### v4

@@
 Extensive experiments validate the effectiveness of this approach, offering actionable insights into SSL with VFMs and paving the way for more scalable and robust semi-supervised learning in the foundation model era.
-footnotetext: *Equal contribution.footnotetext: Our code is available at https://github.com/OSU-MLB/SSL-Foundation-Models.
+footnotetext: Our code is available at https://github.com/OSU-MLB/SSL-Foundation-Models.
 1 Introduction
@@
 First, fine-tuning VFMs with representative SSL algorithms offers limited advantages over using only labeled data for fine-tuning.
-Second, parameter-efficient fine-tuning (PEFT) mai2024lessonslearnedunifyingempirical ; vpetlbench2024 —which updates a small subset of parameters or adds lightweight learnable modules while keeping the VFM largely frozen—consistently yields substantial performance gains, regardless of the learning paradigm.
+Second, parameter-efficient fine-tuning (PEFT) mai2024lessonslearnedunifyingempirical ; vpetlbench2024 ; sun2025strongermixturelowrankexperts —which updates a small subset of parameters or adds lightweight learnable modules while keeping the VFM largely frozen—consistently yields substantial performance gains, regardless of the learning paradigm.
 These observations imply two important directions.
@@
 Vision Foundation Models.
-Vision Transformers (ViT) dosovitskiy2021imageworth16x16words pre-trained on massive amounts of data, has become indispensable to modern AI development awais2025foundation ; mai2025ava ; gu2025bioclip .
+Vision Transformers (ViT) dosovitskiy2021imageworth16x16words pre-trained on massive amounts of data, have become indispensable to modern AI development awais2025foundation ; mai2025ava ; gu2025bioclip .
 These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks.
@@
 Meanwhile, DINOv2 oquab2024dinov2learningrobustvisual pre-trained with self-supervised objectives on extensive sets of well-curated images effectively captures fine-grained localization features.
-Moreover, it is increasingly becoming a common sense that strategically fusing multiple VFMs can unlock synergistic gains mai2024lessonslearnedunifyingempirical ; tong2024eyes , leading to state-of-the-art performance across tasks from visual question answering to object detection.
+Moreover, it is increasingly becoming common sense that strategically fusing multiple VFMs can unlock synergistic gains mai2024lessonslearnedunifyingempirical ; tong2024eyes , leading to state-of-the-art performance across tasks from visual question answering to object detection.
 Leveraging VFMs for Semi-Supervised Learning.
@@
 Our work, in contrast, extends the scope in both the diversity of VFMs and the evaluation benchmarks, establishing the first comprehensive study of SSL in the era of foundation models.
-Table 1: Comparison of supervised linear probing performance between popular SSL datasets (CIFAR, Food101) and our benchmark, where N denotes the number of labeled samples per class.
+Table 1: Comparison of supervised linear probing performance (%) between popular SSL datasets (CIFAR, Food101) and our benchmark, where N denotes the number of labeled samples per class.
 While frozen VFMs already excel in standard SSL datasets, they struggle in our benchmark—comprising diverse tasks, domains, and sizes—underscoring the potential of SSL for unleashing the full potential of VFMs.
@@
 N 44 $100$ $10$ 66 66 22 $80$ 2020 $10$
-CLIP 85.0% 78.3% 80.2% 61.8% 63.7% 69.3% 35.9% 33.1% 51.1%
-DINOv2 91.7% 88.1% 83.1% 66.7% 65.3% 52.4% 41.2% 30.5% 51.3%
+CLIP 85.0 78.3 80.2 61.8 63.7 69.3 35.9 33.1 51.1
+DINOv2 91.7 88.1 83.1 66.7 65.3 52.4 41.2 30.5 51.3
 3 On Evaluation of SSL in the Era of VFMs
@@
 To effectively evaluate SSL methods in the VFM era—focusing exclusively on semi‐supervised image classification—we propose a new benchmark designed to capture the complexities of real‐world applications across different domains and dataset sizes.
-Dataset & Regime We choose six classification datasets from VTAB zhai2020largescalestudyrepresentationlearning , spanning remote sensing, medical imaging, synthetic data, and autonomous driving tasks.
-To test the robustness of SSL methods, we vary the number of labels per class and employ linear probing as our performance metric, selecting shot counts that keep each task sufficiently challenging for frozen VFMs.
-These settings impose substantial difficulty on frozen VFMs (Table 1), underscoring the importance of SSL for unlocking their full potential.
-A summary of the datasets appears in Table 2, with full details provided in Appendix A.
+Dataset & Regime Following the VTAB protocol zhai2020largescalestudyrepresentationlearning , we select six classification datasets covering the three VTAB categories— Natural , Specialized , and Structured .
+Specifically, we choose two datasets from each category: DTD and SUN397 from Natural , RESISC45 and Retinopathy from Specialized , and CLEVR-C and KITTI from Structured .
+These datasets span diverse domains, including texture recognition, scene understanding, remote sensing, medical imaging, synthetic reasoning, and autonomous driving.
+To evaluate the robustness of SSL methods, we vary the number of labeled samples per class and adopt linear probing as the evaluation protocol, with shot counts chosen to keep each task sufficiently challenging for frozen VFMs.
+This configuration imposes substantial difficulty on frozen representations (Table 1), underscoring the necessity of SSL for unlocking their full potential.
+A summary of the datasets appears in Table 2, and detailed descriptions are provided in Appendix A.
 Together, this benchmark offers a more diverse and comprehensive foundation‐model‐era evaluation of semi‐supervised image classification.
@@
 […]
-DTD Recog. Textual 3,008 47 3, 6
+DTD Recog. Textural 3,008 47 3, 6
 SUN397 Recog. Natural 49,601 397 3, 6
@@
 Next, we rank every hyperparameter configuration based on each criterion and choose the one achieving the lowest average rank across all.
-We provide formal definitions and detailed procedures in section˜C.1 and the effectiveness of the proposed method in section˜B.7.
+We provide formal definitions and detailed procedures in section˜C.1 and the effectiveness of the proposed method in section˜6.3.
 By eliminating reliance on held-out labeled validation sets, our method mitigates data leakage and promotes a more practical and fair tuning protocol for SSL.
@@
 If (a) high-confidence predictions are correct and (b) the updated model gradually increases its confidence in the remaining data, self-training can be as effective as fully supervised learning on the entire dataset with true labels.
-In our context, a VFM fine-tuned with PEFT using labeled $\mathcal{L}$ serves as the current model to generate pseudo-labels for $\mathcal{U}$, which are then used to further fine-tune the VFM.
+In our context, a VFM fine-tuned with PEFT using labeled dataset $\mathcal{L}$ serves as the current model to generate pseudo-labels for $\mathcal{U}$, which are then used to further fine-tune the VFM.
 Challenges.
@@
 We present our proposed pipeline in algorithm 1.
+Time Efficiency of Ensembling.
+Among steps (a)(d) in algorithm 1, step (d) dominates the wall-clock time since the pseudo-labeled set is orders of magnitude larger than the labeled one.
+Although our method ensembles $N\times M$ models, the overall time overhead remains marginal: step (a) is fast owing to the small labeled set, and steps (b) and (c) are computationally light.
+For instance, for V-PET, the estimated runtime is only about 1.16× that of other SSL baselines.
 6 Experiment
@@
 We summarize the performance comparison between ST, PET, V-PET and existing SSL methods across all the 12 settings in Table 3.
-Overall, V-PET achieve superior performance compared with others, as shown in the Average column.
+Overall, V-PET achieves superior performance compared with others, as shown in the Average column.
 To further analyze the results, Figure 5 visualize the ranking frequency, where each matrix entry $(i,j)$ indicates how often method $i$ is ranked $j$-th across 12 settings.
@@
 CLIP
-LoRA Labeled Only 57.9% 64.1% 60.7% 66.8% 59.5% 72.1% 35.4% 42.2% 35.2% 50.8% 60.1% 64.3% 55.7%
-Fixmatch 56.8% 68.6% 62.2% 72.8% 47.6% 77.4% 32.5% 39.1% 32.2% 42.6% 59.8% 52.6% 53.7%
-FlexMatch 56.5% 67.0% 66.4% 71.9% 51.2% 78.4% 32.8% 38.7% 36.9% 75.4% 44.6% 55.0% 56.2%
-SoftMatch 59.8% 68.0% 66.3% 70.9% 78.7% 83.7% 42.1% 42.9% 29.8% 71.2% 42.9% 60.5% 59.7%
-FineSSL 63.1% 69.1% 56.3% 61.8% 74.2% 83.5% 35.0% 30.4% 33.6% 38.1% 53.7% 48.2% 53.9%
-ST 58.9% 63.8% 49.2% 52.7% 64.1% 77.4% 35.7% 42.8% 36.5% 51.6% 59.8% 64.3% 54.7%
-PET 63.4% 68.2% 65.6% 70.8% 67.5% 78.6% 36.4% 42.4% 38.3% 54.6% 60.3% 65.7% 59.3%
-V-PET 65.6% 71.7% 67.2% 72.8% 66.6% 77.2% 38.9% 51.1% 36.7% 58.5% 58.0% 62.3% 60.5%
-AdaptFormer Labeled Only 61.5% 65.0% 61.1% 68.2% 61.0% 72.8% 34.4% 39.0% 37.8% 48.6% 57.8% 59.1% 55.6%
-Fixmatch 61.2% 67.3% 64.6% 71.2% 63.1% 80.5% 29.0% 42.4% 33.1% 44.8% 30.0% 57.2% 53.7%
-FlexMatch 60.3% 68.8% 64.3% 70.7% 58.1% 78.1% 31.0% 35.9% 38.5% 52.2% 34.2% 54.2% 53.9%
-SoftMatch 61.3% 69.3% 65.3% 70.2% 65.8% 82.7% 38.2% 42.7% 34.5% 56.1% 33.9% 55.1% 56.3%
-FineSSL 61.9% 68.1% 58.4% 64.9% 73.9% 84.2% 25.3% 29.8% 26.0% 36.9% 42.2% 47.1% 51.6%
-ST 62.0% 67.2% 58.1% 61.3% 68.8% 79.3% 34.4% 39.6% 39.0% 50.3% 58.2% 61.9% 56.7%
-PET 63.8% 68.9% 66.8% 71.7% 69.5% 80.2% 37.1% 41.5% 38.4% 55.8% 60.5% 61.7% 59.7%
-V-PET 65.7% 71.8% 67.8% 73.2% 67.9% 78.4% 38.9% 51.2% 37.0% 59.0% 58.2% 63.3% 61.0%
+LoRA Labeled Only 57.9 64.1 60.7 66.8 59.5 72.1 35.4 42.2 35.2 50.8 60.1 64.3 55.7
+Fixmatch 56.8 68.6 62.2 72.8 47.6 77.4 32.5 39.1 32.2 42.6 59.8 52.6 53.7
+FlexMatch 56.5 67.0 66.4 71.9 51.2 78.4 32.8 38.7 36.9 75.4 44.6 55.0 56.2
+SoftMatch 59.8 68.0 66.3 70.9 78.7 83.7 42.1 42.9 29.8 71.2 42.9 60.5 59.7
+FineSSL 63.1 69.1 56.3 61.8 74.2 83.5 35.0 30.4 33.6 38.1 53.7 48.2 53.9
+ST 58.9 63.8 49.2 52.7 64.1 77.4 35.7 42.8 36.5 51.6 59.8 64.3 54.7
+PET 63.4 68.2 65.6 70.8 67.5 78.6 36.4 42.4 38.3 54.6 60.3 65.7 59.3
+V-PET 65.6 71.7 67.2 72.8 66.6 77.2 38.9 51.1 36.7 58.5 58.0 62.3 60.5
+AdaptFormer Labeled Only 61.5 65.0 61.1 68.2 61.0 72.8 34.4 39.0 37.8 48.6 57.8 59.1 55.6
+Fixmatch 61.2 67.3 64.6 71.2 63.1 80.5 29.0 42.4 33.1 44.8 30.0 57.2 53.7
+FlexMatch 60.3 68.8 64.3 70.7 58.1 78.1 31.0 35.9 38.5 52.2 34.2 54.2 53.9
+SoftMatch 61.3 69.3 65.3 70.2 65.8 82.7 38.2 42.7 34.5 56.1 33.9 55.1 56.3
+FineSSL 61.9 68.1 58.4 64.9 73.9 84.2 25.3 29.8 26.0 36.9 42.2 47.1 51.6
+ST 62.0 67.2 58.1 61.3 68.8 79.3 34.4 39.6 39.0 50.3 58.2 61.9 56.7
+PET 63.8 68.9 66.8 71.7 69.5 80.2 37.1 41.5 38.4 55.8 60.5 61.7 59.7
+V-PET 65.7 71.8 67.8 73.2 67.9 78.4 38.9 51.2 37.0 59.0 58.2 63.3 61.0
 DINOv2
-LoRA Labeled Only 61.0% 68.4% 59.7% 66.8% 48.6% 63.1% 40.7% 48.0% 35.9% 61.7% 46.7% 71.7% 56.0%
-Fixmatch 57.4% 68.4% 56.6% 71.4% 36.6% 66.7% 38.1% 41.2% 35.7% 41.0% 14.8% 42.9% 47.6%
-FlexMatch 52.8% 65.4% 55.7% 67.7% 33.3% 72.6% 36.6% 35.8% 37.1% 66.9% 26.2% 33.6% 48.6%
-SoftMatch 53.6% 64.0% 60.4% 67.3% 46.6% 79.5% 32.1% 44.5% 33.6% 65.1% 27.4% 27.9% 50.2%
-FineSSL 66.9% 72.7% 54.6% 62.1% 69.8% 87.4% 31.0% 24.6% 24.8% 31.7% 31.9% 39.8% 49.8%
-ST 52.3% 56.9% 45.4% 50.5% 54.7% 72.4% 42.6% 49.6% 36.3% 62.7% 33.2% 77.8% 52.9%
-PET 66.7% 72.2% 58.6% 65.3% 55.8% 70.6% 41.0% 54.3% 33.1% 56.9% 42.6% 63.3% 56.7%
-V-PET 67.8% 74.1% 67.1% 73.2% 66.0% 77.0% 39.2% 52.5% 36.8% 59.3% 58.2% 63.7% 61.2%
-AdaptFormer Labeled Only 63.0% 67.6% 58.9% 67.0% 47.0% 57.4% 34.4% 53.1% 27.3% 41.8% 52.3% 49.8% 51.6%
-Fixmatch 59.5% 70.4% 57.9% 69.7% 26.4% 62.3% 38.7% 45.6% 32.7% 38.9% 19.8% 50.4% 47.7%
-FlexMatch 57.1% 65.9% 56.0% 67.5% 51.3% 64.0% 33.4% 40.2% 33.4% 42.2% 34.3% 27.9% 47.8%
-SoftMatch 58.7% 66.2% 60.0% 66.7% 53.1% 78.1% 33.5% 32.2% 32.9% 45.0% 24.5% 30.9% 48.5%
-FineSSL 64.0% 71.2% 56.0% 63.3% 62.3% 69.7% 17.7% 24.3% 25.8% 30.0% 31.8% 45.0% 46.8%
-ST 62.3% 67.0% 59.0% 67.1% 54.1% 66.2% 34.9% 56.0% 27.4% 43.1% 43.9% 48.7% 52.5%
-PET 67.2% 73.0% 63.5% 70.9% 56.2% 72.0% 40.2% 55.5% 33.1% 58.0% 45.2% 60.9% 58.0%
-V-PET 68.2% 74.2% 67.7% 73.4% 66.3% 78.0% 39.4% 52.5% 36.8% 60.0% 59.2% 61.7% 61.5%
-Table 3: Performance comparison of baselines, existing SSL approaches, and our proposed methods on six diverse datasets (12 settings).
+LoRA Labeled Only 61.0 68.4 59.7 66.8 48.6 63.1 40.7 48.0 35.9 61.7 46.7 71.7 56.0
+Fixmatch 57.4 68.4 56.6 71.4 36.6 66.7 38.1 41.2 35.7 41.0 14.8 42.9 47.6
+FlexMatch 52.8 65.4 55.7 67.7 33.3 72.6 36.6 35.8 37.1 66.9 26.2 33.6 48.6
+SoftMatch 53.6 64.0 60.4 67.3 46.6 79.5 32.1 44.5 33.6 65.1 27.4 27.9 50.2
+FineSSL 66.9 72.7 54.6 62.1 69.8 87.4 31.0 24.6 24.8 31.7 31.9 39.8 49.8
+ST 52.3 56.9 45.4 50.5 54.7 72.4 42.6 49.6 36.3 62.7 33.2 77.8 52.9
+PET 66.7 72.2 58.6 65.3 55.8 70.6 41.0 54.3 33.1 56.9 42.6 63.3 56.7
+V-PET 67.8 74.1 67.1 73.2 66.0 77.0 39.2 52.5 36.8 59.3 58.2 63.7 61.2
+AdaptFormer Labeled Only 63.0 67.6 58.9 67.0 47.0 57.4 34.4 53.1 27.3 41.8 52.3 49.8 51.6
+Fixmatch 59.5 70.4 57.9 69.7 26.4 62.3 38.7 45.6 32.7 38.9 19.8 50.4 47.7
+FlexMatch 57.1 65.9 56.0 67.5 51.3 64.0 33.4 40.2 33.4 42.2 34.3 27.9 47.8
+SoftMatch 58.7 66.2 60.0 66.7 53.1 78.1 33.5 32.2 32.9 45.0 24.5 30.9 48.5
+FineSSL 64.0 71.2 56.0 63.3 62.3 69.7 17.7 24.3 25.8 30.0 31.8 45.0 46.8
+ST 62.3 67.0 59.0 67.1 54.1 66.2 34.9 56.0 27.4 43.1 43.9 48.7 52.5
+PET 67.2 73.0 63.5 70.9 56.2 72.0 40.2 55.5 33.1 58.0 45.2 60.9 58.0
+V-PET 68.2 74.2 67.7 73.4 66.3 78.0 39.4 52.5 36.8 60.0 59.2 61.7 61.5
+Table 3: Performance (%) comparison of baselines, existing SSL approaches, and our proposed methods on six diverse datasets (12 settings).
 The best result within each PEFT and VFM is highlighted in bold.
@@
 In particular, we expand our evaluation across two dimensions—VFMs and PEFT methods—to assess the broader applicability of our ensembling strategy.
-Figure 5: Scaling analysis for ensembling more PEFT methods and VFMs. Starting from LoRA fine-tuning, performance improves with diverse pseudo-labels, but gain diminishes with more ensemble sources.
 VFMs Beyond ViT-B (CLIP, DINOv2), we incorporate four more VFMs, for a total of six, including four ViT-B (CLIP, DINOv2, OpenCLIP, ImangeNet21k) cherti2023reproducible ; ridnik2021imagenet21kpretrainingmasses and two ViT-L (CLIP, DINOv2).
@@
 The number in $(i,j)$ indicates the frequency of method $i$ is ranked $j$-th across 12 settings.
-The number in bracket indicates the average rank.
+The number in brackets indicates the average rank, where the higher rank is better.
+6.3 The Effectiveness of Hyperparameter Tuning
+Table 4: Hyperparameter tuning comparison: our method versus baselines (random or chosen by one criterion).
+The absolute errors (%) from oracle test accuracy (smaller the better) are reported averaged across 12 datasets, 5 SSL baselines, and 2 VFMs trained on 2 PEFT.
+The reported values are in the format: mean $\pm$ standard deviation.
+RankMe AMI ARI V-Measure FMI CHI BNM Random Ours
+[…]
+We also investigated whether the proposed hyperparameter tuning strategy accurately gauges SSL methods’ performance.
+To illustrate its advantages, we compare it with two baselines, both measured via the absolute difference from the oracle test performance.
+Specifically, the oracle test performance is defined as the highest accuracy across the entire hyperparameter search space; each method selects a hyperparameter configuration, and we record how far its resulting accuracy is from this optimum.
+By taking the absolute value of this difference, we quantify how closely each tuning method approximates the best possible result.
+Figure 6: Scaling analysis for ensembling more PEFT methods and VFMs.
+Starting from LoRA fine-tuning, performance improves with diverse pseudo-labels, but the gain diminishes with more ensemble sources.
+Mean Label consistently outperforms Mean Logits and Mean Probabilities.
+Our baselines are: (1) Random expectation, which represents the average performance obtained by randomly choosing hyperparameters from the search space, and (2) Single criterion, which selects the best hyperparameter configuration using only one metric.
+Comparing with (1) offers insight into how well our method leverages the search space relative to a purely random choice, while comparing with (2) illustrates the advantage of integrating multiple evaluation criteria.
+As shown in Table 4, we measure the average absolute error between each method’s chosen accuracy and the oracle’s best accuracy (i.e., the maximum accuracy achievable in the search space) across 12 datasets, 5 SSL baselines, 2 VFMs (CLIP and DINOv2), and 2 PEFT methods (LoRA and AdaptFormer).
+Our method outperforms both baselines in most cases, suggesting that combining multiple metrics produces a more reliable and robust assessment of SSL methods.
+We hope future research adopts this approach to strengthen SSL evaluations.
 7 Conclusion
-As the machine learning community increasingly standardizes pre-trained foundation models as backbones across various applications, it becomes crucial to assess whether machine learning algorithms developed prior to the foundation model era remain effective.
-Specifically, given that vision foundation models (VFM) are pre-trained on millions of images, it remains unclear whether they would still benefit from abundant unlabeled data in conjunction with limited labeled data in a semi-supervised learning (SSL) setting.
-To this end, we conduct a comprehensive and systematic empirical study, revealing that VFMs can still benefit from SSL but require a specifically tailored approach.
+We conduct a comprehensive and systematic empirical study for SSL in the era of VFMs, revealing that VFMs can still benefit from SSL but require a specifically tailored approach.
 In particular, we propose a simple yet highly effective self-training algorithm that ensembles pseudo-labels from multiple VFMs and parameter-efficient fine-tuning (PEFT) techniques.
@@
 Results in Table 11 show that, despite ViT-B-IN21k’s lower standalone accuracy, combining its labels with CLIP’s improves overall performance.
-B.7 The Effectiveness of Hyperparameter Tuning
-Table 11: Hyperparameter tuning comparison: our method versus baselines (random or chosen by one criterion).
-The absolute errors from oracle test accuracy (smaller the better) are reported averaged across 12 datasets, 5 SSL baselines, and 2 VFMs trained on 2 PEFT.
-The reported values are in the format: mean $\pm$ standard deviation.
-RankMe AMI ARI V-Measure FMI CHI BNM Random Ours
-[…]
-We begin by investigating whether the proposed hyperparameter tuning strategy accurately gauges SSL methods’ performance.
-To illustrate its advantages, we compare it with two baselines, both measured via the absolute difference from the oracle test performance.
-Specifically, the oracle test performance is defined as the highest accuracy across the entire hyperparameter search space; each method selects a hyperparameter configuration, and we record how far its resulting accuracy is from this optimum.
-By taking the absolute value of this difference, we quantify how closely each tuning method approximates the best possible result.
-Our baselines are: (1) Random expectation, which represents the average performance obtained by randomly choosing hyperparameters from the search space, and (2) Single criterion, which selects the best hyperparameter configuration using only one metric.
-Comparing with (1) offers insight into how well our method leverages the search space relative to a purely random choice, while comparing with (2) illustrates the advantage of integrating multiple evaluation criteria.
-As shown in Table 11, we measure the average absolute error between each method’s chosen accuracy and the oracle’s best accuracy (i.e., the maximum accuracy achievable in the search space) across 12 datasets, 5 SSL baselines, 2 VFMs (CLIP and DINOv2), and 2 PEFT methods (LoRA and AdaptFormer).
-Our method outperforms both baselines in most cases, suggesting that combining multiple metrics produces a more reliable and robust assessment of SSL methods.
-We hope future research adopts this approach to strengthen SSL evaluations.
 Appendix C Hyperparameter Tuning Details
@@
 Adjusted Rand Index (ARI) [30] ARI is a widely used metric in cluster analysis by measuring the similarity between two clustering solutions.
-Given $RI$ is the Rank Index between the true and predicted clustering, ARI is defined by:
+Given $RI$ is the Rand Index between the true and predicted clustering, ARI is defined by:
 $ARI\coloneqq\frac{RI-\text{Expected RI}}{\max(RI)-\text{Expected RI}}$
@@
 Table 12: Complete hyperparameter tuning results for all datasets, models, and SSL baselines. For single validators we only include RankMe.
+Appendix D Limitations
+Our study primarily focuses on the image classification setting, following representative SSL frameworks such as FixMatch.
+While this scope allows for controlled and fair evaluation of SSL principles, it may limit the direct applicability of our findings to other tasks (e.g., detection or segmentation) that often demand task-specific architectures and training paradigms.
+Nevertheless, we believe the core insights from our analysis—such as the behaviors of pseudo-labeling, ensembling, and label-efficiency dynamics—are conceptually transferable.
+We leave a systematic exploration of these directions to future work.